### v2

# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

Abstract
Vision-language foundation models, such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks. Nevertheless, these models may be unreliable under distributional shifts, as their performance is significantly degraded. In this work, we explore how to efficiently leverage class text information to mitigate these distribution drifts encountered by large pre-trained vision-language models (VLMs) during test-time inference. In particular, we propose to generate pseudo-labels for the test-time samples by exploiting generic class text embeddings as fixed centroids of a label assignment problem, which is efficiently solved with Optimal Transport. Furthermore, the proposed adaptation method (CLIP-OT) integrates a multiple template knowledge distillation approach, which replicates multi-view contrastive learning strategies in unsupervised representation learning but without incurring additional computational complexity. Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of CLIP-OT, achieving performance gains of up to $7\%$ over recent state-of-the-art methods, yet being computationally and memory efficient [^1] Code available at https://github.com/ShambhaviCodes/CLIPOT.
1 Introduction
Large pre-trained vision-language models (VLMs), such as CLIP [33] and ALIGN [19], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability. Nevertheless, if the model is exposed to domain drifts between the training and test samples, its performance can be largely degraded [44, 37].
While a straightforward solution to bridge this gap consists in fine-tuning the trained model using domain-specific labeled data [22, 13], this strategy brings several limitations in real-world scenarios, which may hamper their scalability and deployment. First, adapting to new domains requires collecting labeled samples drawn from each distinct distribution. This might be impractical for specific domains and further hinder the real-time adaptation of the trained model for each input test sample. Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [40].
Test-Time Adaptation (TTA) presents a realistic and practical scenario of unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [39, 17, 45]. Nevertheless, despite the recent rise of VLMs, and the popularity of TTA in more traditional deep models, such as CNNs and ViTs, the study of TTA in large pre-trained vision-language models remains less explored. Initial approaches, such as TENT [39], which simply minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [36]. In contrast, more recent strategies resort to pseudo-labels on the inference batch to guide the model adaptation [31]. In particular, the core idea of these later methods is to minimize a classification loss, typically the standard cross-entropy, between the generated pseudo-labels and model predictions, which guides the network updates over multiple iterations. Nonetheless, while this strategy is common in semi-supervised learning, where labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions [17].
In light of the above limitation, we propose a novel approach for CLIP test-time adaptation based on the simple observation that this task can draw inspiration from clustering. Nevertheless, instead of jointly learning the cluster centroids during training, these can be replaced by the fixed class prototypes derived from the text representations. By doing this, the unlabeled samples in a batch will be assigned to the most similar class prototypes, guided by the text embeddings, which serve as a strong class guidance without requiring explicit sample-wise supervision. Furthermore, as our primary goal is to match distributions between different modalities, i.e., class text embeddings and features from the test images, traditional prototype-based clustering methods may lead to suboptimal performance, as they typically assume an unimodal distribution for each cluster. In contrast, we frame the label assignment problem as an Optimal Transport problem, as it can more naturally model multi-modal distributions [23], and is efficiently solved using the Sinkhorn algorithm [9].
Our key contributions can be summarized as follows:
  • We propose to cast the pseudo-labeling strategy in CLIP test-time adaptation as an Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed cluster centroids without requiring further annotations.
  • To solve the label assignment task, we resort to the Sinkhorn algorithm, as it can handle multi-modal distributions and compute label assignments efficiently.
  • We introduce a multi-template knowledge distillation approach, which leverages richer information derived from different text prompts to better guide the adaptation without incurring significant computational overhead.
  • Experiments across 240 scenarios demonstrate the superiority of CLIP-OT over recent state-of-the-art methods, all while avoiding additional computational complexity.
2 Related Work
Test-Time Adaptation.  TTA adapts a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [39, 47, 8, 30]. Existing approaches for TTA in more traditional deep models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [27, 35]; ii) entropy-based approaches [39, 29, 12], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [25], which employ the test-time generated labels for supervising the model. With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt large pre-trained foundation models, particularly CLIP [36, 26, 31, 14], which mainly differs on the different parameter groups updated during adaptation, i.e., prompt tuning [36, 26, 10] or layer-norm [14, 31] strategies. Test-Time Prompt Tuning (TPT) [36] fine-tunes only the input text prompts, which are optimized by minimizing the entropy of the model predictions. To achieve this, TPT generates multiple random augmented views for each test image, incurring additional costs due to performing multiple forward passes through the CLIP vision encoder for each image. Furthermore, TPT is a unimodal approach, which restricts a multi-modal model such as CLIP from fully exploiting its multi-modal capabilities for adaptation. Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to the CLIP text encoder, where both text and visual encoders remain frozen. Nonetheless, keeping the vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively. Most recently, CLIPArTT [14] and WATT [31] propose to fine-tune the normalization layers instead, resorting to pseudo-labels computed from image-to-image and text-to-text similarities.
Clustering for unsupervised representation learning.  Jointly learning how to adapt a set of learnable parameters of a deep network and identify the class assignments can be viewed as clustering and unsupervised representation learning. Thus, our work is closely related to recent literature on deep clustering-based approaches [1, 3, 4, 16, 42, 43, 18, 5]. In [3], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, strategy that is later employed in [4] to pre-train standard supervised deep models. Nevertheless, applying naive $k$-means risks collapsing to only a few imbalanced clusters, making it ineffective. The label assignment task is framed as an instance of the Optimal Transport problem in [1], which stresses that the formulation presented in [3] is not principled, since potential degenerate solutions are avoided via specific implementation choices. Caron et al. [5] propose to enforce consistency between cluster assignments produced for different image views, or augmentations, which avoids expensive pairwise comparisons typically performed in contrastive learning. While we consider a similar formulation to some of these works, several vital differences exist. In particular, all these works, e.g., [4, 1, 3] are used in unsupervised representation learning, and thus the prototypes are learned from the data distribution. In contrast, our approach leverages the text representations to enhance the cluster assignments. Note that text embeddings are accessible for free in VLMs, not incurring additional supervision, as the test image label remains unknown. A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [5], which introduce a computational burden. Indeed, we can leverage individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring additional costs. In fact, these embeddings are computed only once offline. In contrast, image augmentation or “multi-views” strategies must create the additional images for each incoming data batch, which means they perform one forward pass per new augmentation at test time.
3 Background
3.1 Vision-language models
[…]
$p(y=k|\mathbf{x}_{i})=\frac{\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{k}/\tau)}}{\sum_{j=1}^{K}\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{j}/\tau)}},$ (1)
where $\tau$ is a temperature parameter, whose value is learned during pre-training [33], and $\mathbf{z}^{\top}\mathbf{t}$ is the dot product operator.
3.2 Test-Time Adaptation using CLIP
Problem setting.  We address the problem of adapting a pre-trained VLM at test time. In particular, given a model trained on the source domain $\mathcal{D}_{S}$, our goal is to adjust this model online to the new target domain $\mathcal{D}_{T}$, where only unlabeled test data $\{\mathbf{x}_{i}\}^{N}_{i=1}$ is available, and which is received as a stream of batches, from which predictions must be provided.
Vanilla entropy minimization.  Inspired by the semi-supervised learning literature, a straightforward solution to leverage the predicted probability in Eq. 1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [39]:
$\mathcal{L}(p)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
with $N$ denoting the size of the test set, or batch. However, relying on the Shannon entropy [39] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq. (2) might be trivially minimized by assigning all data points to a single, arbitrary label.
4 Proposed solution
The proposed approach aims to overcome the limitations in Eq. 2 objective by finding reliable pseudo-labels in a batch of test-time samples without supervision. To achieve this, we propose CLIP-OT. This novel TTA method aims to better leverage the text embeddings generated by the frozen CLIP text encoder, serving as strong class descriptors to yield label assignments for test samples through Optimal Transport (OT). Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [3], where cluster centroids are iteratively updated based on the prior label assignments. Below, we elaborate on the details of our method, which is illustrated in Figure 1.
Learning objective.  To overcome the limitation of naively minimizing Eq. 2 in the fully unlabeled scenario, we first encode the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
4.1 Online pseudolabels assignments
In the following, we detail how to encode robust posteriors into Eq. 3 by exploiting multi-modal capabilities of CLIP and the test-time batch data distribution. Considering the probability is obtained as in Eq. 1, the objective in Eq. 3 can be expressed as:
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
where $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$ represents the matrix containing the class text prototypes, and the logarithmic term in the right side does not depend on $\mathbf{q}$. If we now express Eq. 4 in its matrix form over all the test images, $\mathbf{Z}$, and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}),$ (5)
where, following [9], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in^{K\times N}_{+}\mid\mathbf{Q}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\},$ (6)
with $\mathbf{1}_{K}$ and $\mathbf{1}_{N}$ denoting the vectors of ones in dimension $K$ and $N$, respectively. The constraints on $\mathcal{Q}$ enforce that on average each prototype is selected at least $\frac{N}{K}$ times, encouraging $\mathbf{Q}$ to be a matrix with uniform marginals in both rows and columns. Note that such constraints “break” the direct dependency between the original prediction $p$ in Eq. 2 and the posterior $q$ in Eq. 4, since the latter consider the whole batch distribution and its marginal properties for the assignment. This results in more robust supervisory signals, later demonstrated empirically (Figure 2).
Figure 1: CLIP-OT. Our method leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq. 9). Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq. 3). Note that, at each test batch, our model runs for $m$ iterations, each leveraging a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
Codes optimization.  As the objective function in Eq. (5) is linear, and the constraints defining $\mathcal{Q}$ are also linear, this is a linear program. Nevertheless, directly optimizing the above learning objective might be time-consuming, particularly as the number of data points and classes increases. To address this issue and allow a faster optimization, we apply the Sinkhorn algorithm [9], which integrates an entropic constraint, enforcing a simple structure on the optimal regularized transport. Hence, the optimization problem becomes:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z})+\varepsilon\mathcal{H}(\mathbf{Q}),$ (7)
where $\mathcal{H}(\mathbf{Q})$ denotes the Shannon entropy function $\mathcal{H}(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ and $\varepsilon$ controls its weight. Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [5], in contrast to other works that employ the full dataset [1]. Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{s}$, where $B_{s}$ denotes the batch size. Now, the soft codes $\mathbf{Q}^{*}$ are the solution of the problem presented in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations as:
$\mathbf{Q}^{*}=\text{Diag}(\mathbf{u}^{(t)})\exp\left(\frac{\mathbf{T}^{\top}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}),$ (8)
where $\mathbf{u}$ and $\mathbf{v}$ are renormalization vectors in K and B respectively, with $t$ denoting the iteration.
Optimizing codes for individual textual views.  Very recent literature [31] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$. Thus, we modify the solution presented in Eq. 8 to be specialized for specific class template [^2] The text templates are the same as in WATT [31] – see Appendix A., where the codes for each template $m$ are:
$\mathbf{Q}^{*}_{m}=\text{Diag}(\mathbf{u}^{(t)}_{m})\exp\left(\frac{\mathbf{T}^{\top}_{m}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}_{m}).$ (9)
The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm, [20], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$. Furthermore, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$ denotes the matrix containing the class text prototypes obtained from the $m$-th template. Note that the computational burden introduced by OT is negligible, representing roughly $1\%$ of the total runtime for a batch, as discussed in Section 5.3.
4.2 Updating CLIP through textual codes
Distilling knowledge from multiple text prototypes.  To optimize the objective in Eq. 3, we employ the soft label assignments, as they yield superior performance compared to their hard counterpart in other problems [5]. Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions per category to the model, the cross-entropy in Eq. 3 is optimized $M$ times, each one using a different template $m$, which can be defined for each image and template as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template as in Eq. 1. Thus, for a given test image $i$ in a batch, after each update produced by the $m$-th text template, the visual encoder produces novel visual embeddings $\mathbf{z}_{i}$, which are then used to generate new predictions $\mathbf{p}_{i}$ for $i$.
Whole learning strategy.  The final proposed algorithm consists in updating the affine parameters in Layer Normalization, such as in [39, 31, 14], and learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq. (7). This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq. (9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given the current label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq. (3) w.r.t. the layer norm parameters, where stochastic gradient descent is applied over the whole batch. This process is repeated $M$ times. Finally, the class predictions on the batch are inferred with the updated model. A view of the proposed method is detailed in Algorithm 1.
Thus, in a nutshell, we generate multiple pseudo-labels per sample, each derived from one text template. Note that pseudo-labeling is a common strategy in other fields, such as self-supervised [49] or semi-supervised [2, 46] learning. Thus, while $\mathbf{Q}$, or pseudo-labels in general, may contain errors, using these to refine the model is a standard practice, which typically yields consistent improvements. We refer the reader to Appendix G for a detailed discussion of the benefits of our approach compared to pseudo-labeling in other learning scenarios.
Interpretation.  Given a batch of $B_{s}$ visual embeddings $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{B}]$, we aim at optimizing a function to map these embeddings into the class prototypes $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$. If we let $\mathbf{Q}$ denote the codes (or mapping), our goal is optimizing $\mathbf{Q}$ to maximize the similarity between the image features and the class text prototypes. While in fully-supervised, or semi-supervised regimes, labeled data is available to generate the class prototypes, in our Test-Time Adaptation scenario this becomes more challenging, as no label information about the test images is provided. Nevertheless, in the context of large vision-language models, the text can be leveraged as a form of supervision to create the prototypes $\mathbf{T}$, which represent a robust representation of each class. Furthermore, the entropic constraint term controls the smoothness of the mapping, promoting equally distributed predictions over all classes in each batch, which avoids degenerate solutions. Lastly, resorting to multiple templates individually, instead of averaging their embeddings, fully leverage the diverse information present in each text prompt.
Algorithm 1 CLIP-OT adaptation for one corruption.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:     for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do    // Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:         $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:         $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:         Compute codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:         // Step 2: Distill knowledge to the model.
9:         $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}\quad(\mathbf{t}_{k}\in\mathbb{R}^{d})$ // average class text embeddings over all M templates.
10:         $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{t}_{k},\forall k$, Eq. (1).
11:         Min. cross-entropy with $\mathbf{P}$ and $\mathbf{Q}^{*}_{m}$ // Eq. (3).
12:         $\theta_{\text{LN}}^{(m-1)}\rightarrow\theta_{\text{LN}}^{(m)}$ // Update layer norm (LN) of $\boldsymbol{\theta}$.
13:     end for // 2 - Inference (for all images in the batch).
14:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:     $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
16:end for
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
CIFAR-10 88.74 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13 91.41 $\pm$0.17 91.05 $\pm$0.06 93.15 $\pm$0.10 +2.10 $\blacktriangle$
CIFAR-10.1 83.25 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27 87.78 $\pm$0.05 86.98 $\pm$0.31 88.37 $\pm$0.10 +1.39 $\blacktriangle$
CIFAR-10C Gaussian Noise 35.27 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36 61.89 $\pm$0.24 63.84 $\pm$0.24 64.85 $\pm$0.14 +1.01 $\blacktriangle$
Shot Noise 39.67 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07 63.52 $\pm$0.08 65.28 $\pm$0.21 67.34 $\pm$0.21 +2.06 $\blacktriangle$
Impulse Noise 42.61 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16 57.13 $\pm$0.02 58.64 $\pm$0.11 62.27 $\pm$0.43 +3.63 $\blacktriangle$
Defocus Blur 69.76 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05 78.86 $\pm$0.09 78.94 $\pm$0.12 82.09 $\pm$0.05 +3.15 $\blacktriangle$
Glass Blur 42.40 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16 62.88 $\pm$0.06 65.12 $\pm$0.07 68.07 $\pm$0.02 +2.95 $\blacktriangle$
Motion Blur 63.97 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19 76.85 $\pm$0.26 77.81 $\pm$0.14 81.30 $\pm$0.18 +3.49 $\blacktriangle$
Zoom Blur 69.83 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20 79.35 $\pm$0.04 79.32 $\pm$0.07 83.13 $\pm$0.24 +3.81 $\blacktriangle$
Snow 71.78 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16 79.44 $\pm$0.09 79.79 $\pm$0.06 83.71 $\pm$0.17 +3.92 $\blacktriangle$
Frost 72.86 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08 80.13 $\pm$0.10 80.54 $\pm$0.12 83.40 $\pm$0.20 +2.86 $\blacktriangle$
Fog 67.04 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14 77.68 $\pm$0.07 78.53 $\pm$0.22 82.56 $\pm$0.08 +4.03 $\blacktriangle$
Brightness 81.87 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16 87.10 $\pm$0.10 87.11 $\pm$0.11 89.90 $\pm$0.06 +2.79 $\blacktriangle$
Contrast 64.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14 80.04 $\pm$0.24 81.20 $\pm$0.22 84.86 $\pm$0.11 +3.66 $\blacktriangle$
Elastic Transform 60.83 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01 71.76 $\pm$0.10 72.66 $\pm$0.15 76.08 $\pm$0.30 +3.42 $\blacktriangle$
Pixelate 50.53 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13 69.28 $\pm$0.09 71.11 $\pm$0.13 76.25 $\pm$0.05 +5.14 $\blacktriangle$
JPEG Compression 55.48 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14 66.49 $\pm$0.14 67.36 $\pm$0.28 70.03 $\pm$0.37 +2.67 $\blacktriangle$
Average 59.22 67.56 56.80 71.17 72.83 73.82 77.06 +3.24 $\blacktriangle$
CIFAR-100 61.68 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04 70.38 $\pm$0.14 70.74 $\pm$0.20 71.68 $\pm$0.09 +0.94 $\blacktriangle$
CIFAR-100C Gaussian Noise 14.80 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14 31.28 $\pm$0.03 32.07 $\pm$0.23 33.43 $\pm$0.13 +1.36 $\blacktriangle$
Shot Noise 16.03 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05 33.44 $\pm$0.11 34.36 $\pm$0.11 35.60 $\pm$0.19 +1.24 $\blacktriangle$
Impulse Noise 13.85 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09 29.40 $\pm$0.11 30.33 $\pm$0.03 30.94 $\pm$0.41 +0.61 $\blacktriangle$
Defocus Blur 36.74 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23 52.32 $\pm$0.28 52.99 $\pm$0.16 53.87 $\pm$0.22 +0.88 $\blacktriangle$
Glass Blur 14.19 3.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03 31.20 $\pm$0.12 32.15 $\pm$0.30 35.26 $\pm$0.22 +3.11 $\blacktriangle$
Motion Blur 36.14 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14 49.72 $\pm$0.15 50.53 $\pm$0.12 52.77 $\pm$0.07 +2.24 $\blacktriangle$
Zoom Blur 40.24 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06 54.72 $\pm$0.04 55.30 $\pm$0.22 56.71 $\pm$0.17 +1.41 $\blacktriangle$
Snow 38.95 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01 51.79 $\pm$0.04 52.77 $\pm$0.15 54.30 $\pm$0.19 +1.53 $\blacktriangle$
Frost 40.56 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12 53.04 $\pm$0.08 53.79 $\pm$0.31 54.92 $\pm$0.07 +1.13 $\blacktriangle$
Fog 38.00 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04 50.78 $\pm$0.24 51.49 $\pm$0.21 53.57 $\pm$0.02 +2.08 $\blacktriangle$
Brightness 48.18 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08 62.65 $\pm$0.25 63.57 $\pm$0.21 64.43 $\pm$0.31 +0.86 $\blacktriangle$
Contrast 29.53 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24 51.34 $\pm$0.10 52.76 $\pm$0.27 55.01 $\pm$0.29 +2.25 $\blacktriangle$
Elastic Transform 26.33 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08 39.97 $\pm$0.06 40.90 $\pm$0.43 43.79 $\pm$0.27 +2.89 $\blacktriangle$
Pixelate 21.98 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14 39.59 $\pm$0.09 40.97 $\pm$0.16 44.51 $\pm$0.32 +3.54 $\blacktriangle$
JPEG Compression 25.91 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32 38.99 $\pm$0.16 39.59 $\pm$0.08 40.83 $\pm$0.14 +1.24 $\blacktriangle$
Average 29.43 35.19 30.46 41.51 44.68 45.57 47.33 +1.76 $\blacktriangle$
[…]
Link to unsupervised representation learning.  The proposed distilling strategy takes inspiration from unsupervised representation learning [7, 28, 5], where comparing random crops of the same image has the potential to yield stronger image representations. In particular, these strategies leverage these novel “views” of the same image as an image augmentation strategy, whose integration improves the semantic quality of the learned image representations. Nevertheless, i) this strategy substantially increases the computational and memory burden [7, 28], ii) specific composition of data augmentation operations is crucial for learning good representations [7], and iii) low-quality crops can alter the model performance [5]. In contrast, exploiting multiple text embeddings to guide the model adaptation process serves as a similar strategy, while also alleviating these shortcomings. More concretely, the text embeddings are low dimensional vectors, which can be computed off-line as they are image agnostic, considerably reducing computation requirements. Furthermore, in the absence of supervision, we argue that diverse text descriptions associated with a given class convey more informative representations than individual crops or “views” of an image to learn more generalizable features. While images provide valuable visual information, their parts may lack the contextual richness necessary for effective representation learning. This distinction suggests that leveraging a variety of textual prompts can indeed serve as a more promising alternative to contrastive learning based on multiple image crops while also being more efficient.
5 Experiments
5.1 Setting
Datasets.  All methods are benchmarked on standard TTA datasets known for their complexity and task diversity. These datasets include corrupted modifications from well-known datasets proposed in [15], such as CIFAR-10C, CIFAR-100C, or Tiny-ImageNet-C. Also, we employ its non-corrupted versions, i.e., CIFAR-10 and CIFAR-100 [21], CIFAR-10.1 [34], and Tiny-ImageNet [41]. Additionaly, we include Domainbed datasets (PACS [24], OfficeHome [38], VisDA-C [32]). More details can be found in Appendix B.
Baselines.  We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [39], TPT [36], CLIPArTT [14], and WATT [31] (both versions WATT-S and WATT-P).
Architectures.  We employed the Vision Transformer architecture as the backbone for our experiments. The choice of ViT is driven by its superior performance on image classification tasks due to its attention mechanisms, which effectively capture long-range dependencies and spatial relationships within images. If the opposite is not stated, experiments are carried out using ViT-B/32, although we also provide further evaluations with larger ViT-backbones, e.g., ViT-B/16.
Implementation details.  Following the existing literature [31, 14], we used a batch size of 128 in our experiments and a learning rate of $10^{-4}$. The only additional hyper-parameters specific to our formulation are the entropic constraint weight, $\epsilon$ in Eq. 7, and the number of repetitions in the Sinkhorn algorithm. These are fixed to the same values across all experiments. $\epsilon$ is set to $\epsilon=0.7$, whereas the latter is fixed to $T=3$, which was enough for convergence, as in previous works [3]. Finally, we ran the adaptation experiments for each corruption/dataset three times under different random seeds and averaged the obtained accuracy accordingly.
5.2 Main results
We start our empirical validation by comparing the proposed approach’s performance to existing state-of-the-art CLIP-based TTA approaches across the different datasets.
Evaluation in the presence of natural or no domain shift.  Table 1 reports the performance of the different methods for CIFAR-10, CIFAR-10.1, and CIFAR-100, which introduce a natural or no domain shift. The proposed method consistently outperforms existing approaches, even when the domain shift is small, highlighting the usefulness of resorting to our approach over CLIP zero-shot predictions.
Impact of the performance under common corruptions.  Furthermore, in Table 1, we report a detailed analysis of the results obtained on CIFAR-10C and CIFAR-100C, which show the superiority of CLIP-OT in adapting CLIP in the presence of common corruptions. In particular, compared to the baseline CLIP, our model brings performance gains of nearly $18\%$ in CIFAR-10C and CIFAR-100C, without additional supervision. These performance gains are similar when compared to other popular TTA methods, e.g., TENT, or TPT, with differences ranging from $10\%$ to $21\%$. While this gap is reduced compared to very recent approaches, i.e., both versions of WATT and CLIPArTT, per-corruption classification accuracy differences are still significant, with CLIP-OT outperforming these two approaches by up to $5\%$ compared to WATT, i.e., $5.14\%$ in pixelate (CIFAR-10C) or $3.11\%$ in Glass Blur (CIFAR-100C), and nearly $11\%$ compared to CLIPArTT, e.g., $10.62\%$ in pixelate and $8.11\%$ in Gaussian Noise. As average, CLIP-OT improves the performance of very recent state-of-the-art approaches by a substantial difference of $3.24\%$ (CIFAR-10C), and $1.76\%$ in the more challenging scenario of CIFAR-100C, which contains $\times$10 number of classes.
What if the number of classes further increases?  In the following, we assess the performance of all methods on Tiny-ImageNet and Tiny-ImageNet-C benchmarks, which bring additional complexity due to the increased amount of categories present. From these results (Table 2), we can observe that the findings reported in previous sections hold, i.e., CLIP-OT significantly outperforming very recent state-of-the-art TTA approaches. More concretely, compared to WATT-S, the average improvement over all the corruptions is above $7.0\%$, with our method showing performance gains superior to $10\%$ in multiple corruptions, e.g., Shot Noise or Pixelate, and above $4.5\%$ in all remaining corruptions. Given the complexity and additional challenges that Tiny-ImageNet-C brings compared to CIFAR versions, as well as the non-negligible improvement provided by our method, we believe that these results showcase the clear superiority of CLIP-OT in more challenging scenarios.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.29 57.72 58.90 59.85 61.35 63.69$\blacktriangle$
Gaussian Noise 7.08 8.01 9.29 14.44 13.02 21.40 $\blacktriangle$
Shot Noise 9.41 10.04 11.70 17.44 15.94 24.90 $\blacktriangle$
Impulse Noise 3.44 4.18 4.85 10.37 6.90 17.34 $\blacktriangle$
Defocus Blur 21.71 24.53 27.56 31.46 29.91 35.39 $\blacktriangle$
Glass Blur 9.12 10.09 11.03 15.84 14.01 21.16 $\blacktriangle$
Motion Blur 34.52 36.94 38.97 41.34 41.26 46.26 $\blacktriangle$
Zoom Blur 27.44 29.48 34.29 35.06 33.96 40.93 $\blacktriangle$
Snow 32.51 32.20 34.45 36.86 37.76 42.32 $\blacktriangle$
Frost 36.33 35.72 37.13 38.20 39.65 44.97 $\blacktriangle$
Fog 25.94 27.46 28.89 33.44 32.13 38.60 $\blacktriangle$
Brightness 40.15 39.79 43.31 46.43 46.93 53.10 $\blacktriangle$
Contrast 1.81 2.24 3.15 6.24 3.53 11.88 $\blacktriangle$
Elastic Transf. 30.40 31.92 33.88 33.89 35.01 40.73 $\blacktriangle$
Pixelate 22.78 24.79 27.70 34.85 31.55 41.84 $\blacktriangle$
JPEG Compr. 29.59 30.93 33.60 37.32 36.46 42.86 $\blacktriangle$
Mean 22.14 23.22 25.32 28.88 27.87 34.91 $\blacktriangle$
Table 2: Performance comparison on Tiny-ImageNet and Tiny-ImageNet-C corruption benchmarks using ViT-B/32 backbone.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C (3D) 84.43 84.86 79.35 85.09 85.42 85.36 90.73 +5.37 $\blacktriangle$
VisDA-C (YT) 84.45 84.68 83.57 84.40 84.57 84.69 85.44 +0.75 $\blacktriangle$
Table 3: Performance comparison across PACS, OfficeHome, and VisDA-C datasets using ViT-B/32 backbone. $\Delta$ represents the difference between CLIP-OT and the State-of-the-art WATT-S [31]. Extended numerical values are presented in Appendix C.2, Table 9.
Performance under simulated and video shifts.  Table 3 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of CLIP-OT compared to other methods. More concretely, the average difference compared to SoTA (e.g., WATT-S) is $3.03\%$, whereas per-dataset differences can go up to $11\%$, e.g., $11.38\%$ compared to TPT in 3D domain.
Texture and style shifts.  We now evaluate CLIP-OT on different distributional drifts, which may bring additional complexity. The results from these experiments (Table 3) show that the trend of superior performance is observed in almost all datasets (i.e., PACS and OfficeHome).
5.3 Ablation studies
Figure 2: Ablation on each component across multiple datasets. Different results come from adding each element of the proposed formulation. In green, performance differences compared to CLIP. The numerical values are provided in Appendix C.
Impact of each component.  In Figure 2, we empirically validate the benefit of each component in CLIP-OT design. Each configuration is detailed in Appendix C.1. First, we can observe that computing the code $Q^{*}_{m}$ from each template $m$ (Eq. (9)) and averaging those to obtain the final predictions consistently improves over zero-shot CLIP. It is noteworthy to highlight that the transportation polytope $\mathcal{Q}$ enforces additional explicit constraints on $\mathbf{Q}$, which ensure a well-defined probabilistic assignment and “break” the direct dependency between $\mathbf{Q}$ and $\mathbf{P}$. This explains the differences between direct CLIP predictions and this version of our approach, which directly leverages $\mathbf{Q}$ as predictions. Furthermore, while this performance is far from state-of-the-art approaches, this strategy is training-free, not requiring additional computational complexity due to gradient descent updates. Next, Average Template employs the pseudo-codes computed in the previous version to minimize the cross-entropy in Eq. (3). Furthermore, the prediction for the $i$-th image, $\mathbf{p}_{i}$, is computed with the average template for the class $k$, i.e., $\mathbf{t}_{k}$. This shows that leveraging the obtained pseudo-codes to refine the model leads to a performance improvement, which is consistent across all datasets. Last, we introduce the novel knowledge distillation strategy proposed in this work, where distinct templates are used to get multiple pseudo-codes and guide the model training, i.e., CLIP-OT. Its substantial performance gains are especially notable in the most challenging scenarios, e.g., CIFAR-100C.
On the impact of multiple templates.  We now assess how the performance evolves as the number of templates increases (Table 4). Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$. For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset. Results from this experiment show that CLIP-OT performance consistently increases with the number of templates. This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
##\## Templates 1 2 3 4 5 6 7 8
CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
Table 4: Accuracy (%) of CLIP-OT with varying numbers of templates (see Appendix A) on natural domain shift benchmarks.
Do observations hold across backbones?  Table 5 reports the performance across different datasets when ViT-B/16 is employed as the visual encoder of CLIP. These results demonstrate that CLIP-OT is model agnostic, as its superiority is consistently maintained across different backbones.
Dataset CLIP TENT TPT CLIPArTT WATT-S CLIP-OT
CIFAR-10 89.25 92.75 89.90 92.61 91.97 94.69
CIFAR-10.1 84.00 88.52 83.75 88.72 88.10 89.25
CIFAR-10C 60.15 68.00 59.75 73.22 76.22 80.11
CIFAR-100 64.76 71.73 67.15 71.34 72.85 74.06
CIFAR-100C 32.01 37.90 33.73 40.08 48.95 51.24
TinyImageNet 58.12 63.59 63.87 64.73 66.05 66.34
TinyImageNet-C 20.92 29.78 26.96 32.90 31.66 37.69
Table 5: Quantitative results with ViT-B/16 as backbone for TTA across different datasets and baselines.
Computational analysis.  Figure 3 depicts the running time required for relevant baselines and the proposed CLIP-OT across several datasets, ordered in terms of increasing complexity, i.e., number of categories. These results expose that different TTA methods, particularly SoTA, substantially differ in total runtimes. In particular, recent CLIPArTT and WATT-S constantly increase the required runtime with the number of classes, driven by their iterative nature on the multiple template evaluations and weight averaging during inference. In contrast, CLIP-OT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once. Moreover, the Sinkhorn algorithm—used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime. This combination of speed and consistency establishes CLIP-OT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT-S for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
Figure 3: Computational efficiency. The runtime of recent CLIP TTA methods is assessed, using a standard input batch of size 128. Runtime differences compared to our CLIP-OT highlighted in red.
6 Conclusions
We have presented a novel approach for TTA of VLMs. In particular, CLIP-OT generates pseudo-labels for test samples using multiple, generic class templates, optimizing a label assignment problem solved with Optimal Transport. This strategy yields state-of-the-art results compared to recent works in several benchmarks, yet being a significantly more efficient solution in terms of computational complexity.
Supplementary Material
Appendix A Templates Used
In the experimental setup, several predefined text templates from CLIP [11] were used to evaluate the proposed model’s adaptability and performance. These are designed to generalize across different contexts and image types, enabling a robust assessment of the model’s capabilities in handling diverse visual and textual representations. In Table 6, we provide the templates used in our experiments. Each template includes a placeholder , which is dynamically replaced with the class name during the generation of text prompts. Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [31].
Template
1: a photo of a {class $k$}
2: itap of a {class $k$}
3: a bad photo of the {class $k$}
4: a origami {class $k$}
5: a photo of the large {class $k$}
6: a {class $k$} in a video game
7: art of the {class $k$}
8: a photo of the small {class $k$}
Table 6: The different templates used during the experiments.
Appendix B Additional details on the datasets
CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems. These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set. Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes. It is a useful benchmark for lightweight models and educational purposes, making it easier to implement and experiment with domain adaptation techniques. Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions, similar to CIFAR-10-C and CIFAR-100-C. This dataset is instrumental for testing the robustness of models against a variety of distortions in a more manageable setting. To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [14]. Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
Appendix C Additional details of the ablation studies
C.1 Configuration: Baselines
In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and CLIP-OT with Average Template. Note that, in short, Training-free OT motivates the use of the pseudo-codes $\mathbf{Q}^{*}$ over the CLIP baseline, whereas CLIP-OT with Average Template includes these pseudo-codes to fine-tune the model. Furthermore, we want to highlight that the later does not fully leverage multiple individual templates, which is introduced in our proposed CLIP-OT model. A more detailed description follows. Given a batch of test images, Training-free OT (Algorithm 2) computes pseudo-codes for each text template, which are later averaged to produce a final prediction. On the other hand, CLIP-OT with Average Template (Algorithm 3) utilizes the averaged pseudo-codes (obtained by Training-free OT) to refine the visual encoder. More concretely, at each batch, the average pseudo-codes $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model. These predictions are obtained by resorting to Eq. (1), where the average class embedding $\mathbf{t}_{k}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$. Differences between the pseudo-codes and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model. Then, once the model is updated, we can do the final inference for the test images.
C.2 Extended numerical values
We further substantiate the findings presented in Figure 2 by providing detailed results in Tables 7 and 8. The tables showcase a comprehensive performance comparison of the impact of the different components of our approach, which empirically motivate our model. These results are reported for different corruption benchmarks on the CIFAR-10, CIFAR-10.1, CIFAR-10C, CIFAR-100, and CIFAR-100C datasets, respectively, using a ViT-B/32 backbone. Also, Table 9 extends the results obtained on the different domains of PACS and OfficeHome, whose average results were reported in the main paper. We can observe that only in one out of the different datasets and domains, CLIP-OT underperforms previous SoTA, with a marginal performance gap, i.e., -0.19%.
Algorithm 2 Training-free OT.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:       end for // Inference (for all images in the batch).
9:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
10:end for
Algorithm 3 CLIP-OT with Average Template.
1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$). // Split $\mathcal{X}$ into $B$ batches of size $B_{s}$. // Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
8:       end for // Step 2: Compute average pseudo-codes and refine the encoder.
9:       $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average pseudo-codes.
10:       $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}$ // average class text embeddings over all templates.
11:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
12:       Min. cross-entropy loss between $\mathbf{P}$ and $\tilde{\mathbf{Q}}$ // Eq. (3).
13:       $\theta_{\text{LN}}^{(t)}\rightarrow\theta_{\text{LN}}^{(t+1)}$ // update layer norm. // 2 - Inference (for all images in the batch).
14:       $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
15:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
16:end for
Dataset CLIP Training-Free OT Avg. Template CLIP-OT
CIFAR-10 88.74 89.44 91.40 93.15
CIFAR-10.1 83.25 84.30 85.50 88.37
Gaussian Noise 35.27 46.68 54.68 64.85
Shot Noise 39.67 49.84 57.13 67.34
Impulse Noise 42.61 47.18 52.72 62.27
Defocus Blur 69.76 73.22 77.84 82.09
Glass Blur 42.40 50.09 56.88 68.07
Motion Blur 63.97 70.15 74.81 81.30
Zoom Blur 69.83 74.19 77.99 83.13
Snow 71.78 73.93 78.41 83.71
Frost 72.86 75.66 79.31 83.40
Fog 67.04 69.28 75.54 82.56
Brightness 81.87 82.90 86.87 89.90
Contrast 64.37 67.30 75.30 84.86
Elastic Transform 60.83 64.06 69.54 76.08
Pixelate 50.53 56.65 62.91 76.25
JPEG Compression 55.48 59.75 64.31 70.03
Mean 59.22 64.05 69.62 77.06
Table 7: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets as shown in Figure 2.
Corruption CLIP Training-Free OT Avg. Template CLIP-OT
CIFAR-100 61.68 61.55 68.26 71.68
Gaussian Noise 14.80 21.26 23.29 33.43
Shot Noise 16.03 22.98 25.15 35.60
Impulse Noise 13.85 22.27 22.14 30.94
Defocus Blur 36.74 41.40 46.84 53.87
Glass Blur 14.19 22.47 23.60 35.26
Motion Blur 36.14 40.71 44.41 52.77
Zoom Blur 40.24 45.14 49.97 56.71
Snow 38.95 44.14 46.73 54.30
Frost 40.56 45.03 47.85 54.92
Fog 38.00 40.00 45.45 53.57
Brightness 48.18 51.26 57.22 64.43
Contrast 29.53 33.69 41.16 55.01
Elastic Transform 26.33 33.02 34.85 43.79
Pixelate 21.98 27.72 30.13 44.51
JPEG Compression 25.91 31.69 33.80 40.83
Mean 29.43 34.85 38.17 47.33
Table 8: Numerical Analysis for CIFAR-100 and CIFAR-100-C datasets as shown in Figure 2.
Dataset Domain CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
PACS Art 96.34 96.65 $\pm$0.05 95.52 $\pm$0.20 96.57 $\pm$0.09 96.31 $\pm$0.00 96.39 $\pm$0.00 96.54 $\pm$0.05 +0.15 $\blacktriangle$
Cartoon 96.08 96.22 $\pm$0.05 94.77 $\pm$0.20 96.00 $\pm$0.02 96.52 $\pm$0.02 96.62 $\pm$0.02 98.34 $\pm$0.13 +1.72 $\blacktriangle$
Photo 99.34 99.40 $\pm$0.00 99.42 $\pm$0.06 99.28 $\pm$0.00 99.48 $\pm$0.03 99.52 $\pm$0.00 99.86 $\pm$0.03 +0.34 $\blacktriangle$
Sketch 82.85 82.96 $\pm$0.12 83.22 $\pm$0.14 83.93 $\pm$0.14 86.92 $\pm$0.04 86.65 $\pm$0.12 89.01 $\pm$0.29 +2.36 $\blacktriangle$
Mean 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
OfficeHome Art 73.75 74.03 $\pm$0.27 75.76 $\pm$0.27 73.84 $\pm$0.20 75.65 $\pm$0.27 75.76 $\pm$0.39 79.39 ±$\pm 0.31$ +3.63 $\blacktriangle$
Clipart 63.33 63.42 $\pm$0.04 63.08 $\pm$0.31 63.54 $\pm$0.06 66.23 $\pm$0.13 65.77 $\pm$0.11 66.21 $\pm$0.17 +0.44 $\blacktriangle$
Product 85.32 85.51 $\pm$0.08 84.07 $\pm$0.28 85.23 $\pm$0.16 85.41 $\pm$0.09 85.41 $\pm$0.01 86.81 $\pm$0.03 +1.40 $\blacktriangle$
[…]
Mean 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
VisDA-C 3D (trainset) 84.43 84.86 $\pm$0.01 79.35 $\pm$0.04 85.09 $\pm$0.01 85.42 $\pm$0.03 85.36 $\pm$0.01 90.73 $\pm$0.65 +5.37 $\blacktriangle$
YT (valset) 84.45 84.68 $\pm$0.01 83.57 $\pm$0.04 84.40 $\pm$0.01 84.57 $\pm$0.00 84.69 $\pm$0.01 85.44 $\pm$0.65 +0.75 $\blacktriangle$
Mean 84.44 84.77 81.46 84.75 85.00 85.03 88.08 +3.05 $\blacktriangle$
[…]
Appendix D Comparison to OT-VP
The very recent OT-VP [48] has presented a solution integrating optimal transport for test-time adaptation. Nevertheless, it presents several fundamental differences with our work. First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized. Thus, optimal transport is used for a different task. Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality. And third, the adaptation scenario strongly differs from our setting. In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome). While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 10. In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps. To expand the extent of our empirical validation, we compare our approach to OT-VP [48] whose results are reported in Table 10. To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
Method PACS OfficeHome Avg.
First Setting : Single Source
[…]
Second Setting : Multi Source
[…]
Our setting: No specific source
CLIP-OT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
Table 10: Performance comparison to [48] across PACS and OfficeHome datasets. Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
Appendix E Differences wrt PLOT [6]
Note that despite the apparent similarities, our work presents fundamental differences w.r.t. Prompt Learning with Optimal Transport (PLOT) [6], which prevents from adding the later in the empirical evaluation. a) Target task. PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available. Thus PLOT cannot be compared in the TTA setting. b) Adaptation. Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category). c) Use of OT. PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities. In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.
Appendix F Additional Details on Epsilon ($\epsilon$)
In the main paper, we stated that $\epsilon$, the entropic constraint weight in Eq. 7 was set to $\epsilon=0.7$ based on preliminary experiments. However, here in the appendix, we explore the sensitivity of the model to different values of $\epsilon$ to assess its impact on accuracy under various corruptions and datasets. As shown in Table 11, we experimented with $\epsilon$ values of 0.3, 0.5, 0.7 and 0.9 across CIFAR-10C and CIFAR-100C datasets. Further, we show the values for CIFAR-10 and CIFAR-100 datasets along with averaged performance over corruptions in the Figure 4.
Corruption CIFAR-10 CIFAR-100
0.3 0.5 0.7 0.9 0.3 0.5 0.7 0.9
Original 91.15 93.01 93.15 92.81 62.26 70.95 71.62 71.01
Gaussian Noise 56.16 63.65 64.43 65.26 13.74 32.03 33.2 32.44
Shot Noise 57.53 66.29 66.27 67.4 15.62 34.69 35.33 35.46
Impulse Noise 54.27 61.22 61.84 62.42 19.44 29.94 30.99 30.36
Defocus Blur 76.27 81.33 82.1 82.49 39.16 52.77 54.16 53.32
Glass Blur 57.63 66.75 67.55 67.96 17.06 34.46 35.24 34.89
Motion Blur 71.99 80.11 81.31 80.9 36.59 51.33 52.67 51.62
Zoom Blur 76.24 82.54 83.11 83.46 42.58 55.53 56.93 56.13
Snow 78.59 83.39 82.99 83.56 40.91 53.54 54.06 54.12
Frost 78.92 82.67 82.93 83.41 40.98 54.1 54.93 54.37
Fog 75.75 81.64 81.9 82.47 40.76 52.69 53.6 53.12
Brightness 85.97 89.59 89.72 89.71 48.31 63.27 64.87 63.57
Contrast 70.38 83.55 84.53 84.6 32.09 53.61 55.39 54.41
Elastic Transform 69.17 75.42 75.67 76.09 27.78 42.69 43.78 43.01
Pixelate 61.31 74.57 76.51 76.08 22.79 42.64 44.96 43.18
JPEG Compression 66.02 69.42 70.46 70.03 30.16 40.16 41.02 40.67
Mean 70.46 76.14 76.75 77.06 31.20 46.23 47.41 46.71
Table 11: Mean Accuracy for CIFAR-10 and CIFAR-100 under Different Corruptions and Epsilon Values
We can observe that higher $\epsilon$ values (0.7 and 0.9) improved stability and performance due to stronger regularization introduced by higher entropic constraints, which prevents degenerate solutions during Sinkhorn normalization. While smaller $\epsilon$ values sometimes performed better under mild corruptions, they exhibited significant instability during Sinkhorn normalization. In several cases, the $\mathbf{Q}$ matrix became NaN, especially under high noise corruptions, rendering the adaptation process unusable which can be attributed to insufficient regularization, causing numerical instabilities during iterative updates. The chosen value ($\epsilon=0.7$) strikes a balance between stability and performance, yielding robust results across both clean and corrupted datasets.
Figure 4: Ablation on epsilon values across multiple datasets.
Appendix G Detailed benefits of our multi-template distillation approach.
The proposed multi-template distillation strategy resembles to the unsupervised approach employed to leverage pseudo-labels in semi-supervised learning, e.g., [2]. In this learning scenario, unlabeled samples are augmented, whose predictions are averaged to obtain a pseudo-label as $\mathbf{q}_{u}=\frac{1}{M}\sum_{m}f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$, where $\bar{\mathbf{x}}_{u,m}$ is the $m$-th augmented version of a given unlabeled image $\mathbf{x}_{u}$. Once the average pseudo-label is obtained, we can minimize the pseudo cross-entropy as $-\sum_{k}\mathbf{q}_{u}\log f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$ [^3] For simplicity, we consider only one unlabeled image.. In this setting, the predictions from multiple augmented images are guided by the average pseudo-label. Once the loss is computed over all the predictions from the augmented images, the gradient is back-propagated and the network parameters updated. Limitations of this method: First, this approach forces all augmentations to align with a single static target $\mathbf{q}$, potentially causing the model to converge to a suboptimal solution because it cannot make incremental adjustments based on each individual augmentation. Second, it treats all augmentations as being equally aligned with a single target $\mathbf{q}$, which could limit the model’s ability to handle diverse augmentations, especially if some augmentations are very different from others. And last, this strategy applies a global update that may force the model too quickly into a solution where all augmentations have the same prediction target $\mathbf{q}$, potentially making it harder for the model to learn robust features. To overcome these limitations in the TTA scenario, where no single supervision is available, the proposed method takes an opposite direction to these well-established practices. More concretely, our approach minimizes the following pseudo cross-entropy loss at each step $m$: $-\mathbf{q}_{m}\log\mathbf{p}$, where $\mathbf{q}_{m}$ is the pseudo-label at step $m$ (derived from the $m$-th text prompt, which is equivalent to one augmentation), and $\mathbf{p}$ is the mean prediction considering all the $M$ text prompts (similar to averaging over $M$ augmentations). At step $m$, the gradient is backpropagated, the network parameters updated. Then, we can recompute the pseudo-labels at step $m+1$, as well as the new mean prediction $\mathbf{p}$ (which will be now different due to the parameters update). Advantages: First, by updating the model after each step and recomputing $\mathbf{p}$ as the mean of the model’s outputs, our approach encourages a more fine-grained, iterative refinement of predictions, potentially helping the model to converge more smoothly. Indeed, the model can adjust its predictions for each template (i.e., augmentation) before being evaluated again in the next step. Furthermore, this step-by-step refinement process allows the model to better adapt to the diversity in the augmentations and learn a more generalizable representation.
Appendix H Discussion on the Two Hyperparameters of WATT and Motivation for CLIP-OT
WATT [31] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process). While these hyperparameters play an essential role in improving WATT’s adaptation capabilities, they introduce significant computational and scalability challenges, particularly as the number of classes grows. The iterative nature of $L$ requires multiple forward and backward passes for every template, leading to substantial runtime overhead, especially for datasets with many classes like Tiny-ImageNet. Additionally, higher $L$ values, while effective for complex corruptions, risk overfitting to noisy pseudo-labels, while lower $L$ values may lead to under-adaptation. Similarly, $M$ stabilizes adaptation through repeated weight averaging, but its effectiveness diminishes beyond a certain point, with larger values providing negligible performance improvements while exponentially increasing runtime. This dependence on repeated updates for each class and template becomes especially prohibitive for datasets with a large number of classes, as demonstrated in Figure 3, where runtime increases from CIFAR-10 (10 classes) to Tiny-ImageNet (200 classes).
These limitations motivate our proposed approach, CLIP-OT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$. CLIP-OT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging. This not only drastically reduces computational overhead but also ensures scalable, robust adaptation across diverse corruption levels and datasets, offering a practical alternative to WATT’s hyperparameter-dependent framework.
Appendix I Further results on other backbones
In this section, we present additional experimental results leveraging different backbone architectures to evaluate the robustness and generalizability of our proposed method. Specifically, we analyze the performance of CLIP models with ViT-B/16 and ViT-L/14 backbones across various datasets, including CIFAR, Tiny-ImageNet and their corruptions. These experiments validate our approach across diverse settings and highlight its adaptability to challenging domain shifts.
Table 12 details the performance of CLIP-OT with the ViT-B/16 backbone on CIFAR-10, CIFAR-10.1, and CIFAR-10-C datasets. Our proposed method demonstrates consistent improvements across both clean and corrupted data, outperforming baseline methods such as CLIP, TENT, and TPT. Similarly for ViT-L/14 backbone on these datasets, we obtain similar gains as illustrated in Table 13. Table 14 extends the evaluation to CIFAR-100 and CIFAR-100-C, demonstrating the effectiveness of our method on more complex datasets with larger label spaces. Using CLIP with the ViT-B/16 backbone, our approach achieves the highest average accuracy, with substantial improvements under corruptions such as Glass Blur ($+5.30\%$) and Contrast ($+6.23\%$). We also evaluate these datasets with backbone ViT/L-14 in Table 15.
For Tiny-ImageNet and its corrupted version, Tiny-ImageNet-C, Table 16 and Table 17 provides a detailed breakdown. On datasets with higher complexity and larger label spaces, such as Tiny-ImageNet and Tiny-ImageNet-C (Table 16), CLIP-OT exhibits substantial gains, consistent with the main paper’s observations. Specifically, our method improves over WATT-S by over 6% on average across Tiny-ImageNet-C corruptions, showcasing its scalability to more challenging scenarios.
I.1 Tiny-Imagenet Experiments
We report the mean and standard deviation over Tiny-Imagenet experiments (Table 2, main paper) in Table 18.
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05 92.31 $\pm$0.10 91.97 $\pm$0.03 94.69 $\pm$0.18
CIFAR-10.1 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33 87.90 $\pm$0.11 88.10 $\pm$0.08 89.25 $\pm$0.08
Gaussian Noise 37.75 31.04 $\pm$0.38 35.35 $\pm$0.15 60.89 $\pm$0.26 63.10 $\pm$0.12 65.57 $\pm$0.22 68.50 $\pm$0.19
Shot Noise 41.10 40.54 $\pm$0.41 41.03 $\pm$0.19 65.19 $\pm$0.21 66.31 $\pm$0.10 68.67 $\pm$0.37 71.24 $\pm$0.22
Impulse Noise 51.71 58.03 $\pm$0.16 54.86 $\pm$0.07 67.55 $\pm$0.09 69.62 $\pm$0.12 70.39 $\pm$0.11 76.97 $\pm$0.17
Defocus Blur 70.07 77.57 $\pm$0.03 70.29 $\pm$0.02 78.92 $\pm$0.12 79.64 $\pm$0.08 79.90 $\pm$0.07 83.33 $\pm$0.09
Glass Blur 42.24 47.16 $\pm$0.05 37.86 $\pm$0.17 57.18 $\pm$0.20 58.98 $\pm$0.12 61.62 $\pm$0.21 68.33 $\pm$0.20
Motion Blur 65.81 76.16 $\pm$0.05 67.43 $\pm$0.11 76.59 $\pm$0.06 78.32 $\pm$0.16 79.02 $\pm$0.07 83.13 $\pm$0.10
Zoom Blur 72.50 79.64 $\pm$0.12 72.91 $\pm$0.02 79.62 $\pm$0.11 80.67 $\pm$0.07 81.10 $\pm$0.08 85.35 $\pm$0.11
Snow 73.23 81.68 $\pm$0.03 72.98 $\pm$0.32 81.13 $\pm$0.29 81.99 $\pm$0.10 82.54 $\pm$0.18 86.65 $\pm$0.12
Frost 76.52 83.22 $\pm$0.05 75.87 $\pm$0.16 81.24 $\pm$0.08 83.41 $\pm$0.16 83.46 $\pm$0.15 86.37 $\pm$0.09
Fog 68.35 80.78 $\pm$0.15 69.13 $\pm$0.27 78.47 $\pm$0.19 81.36 $\pm$0.12 81.88 $\pm$0.12 85.60 $\pm$0.24
Brightness 83.36 89.85 $\pm$0.11 83.67 $\pm$0.14 88.66 $\pm$0.15 89.06 $\pm$0.05 89.10 $\pm$0.14 92.64 $\pm$0.11
Contrast 61.90 79.24 $\pm$0.19 62.16 $\pm$0.06 75.15 $\pm$0.07 81.57 $\pm$0.23 83.79 $\pm$0.12 86.92 $\pm$0.18
Elastic Transform 53.16 62.54 $\pm$0.08 51.26 $\pm$0.23 69.49 $\pm$0.08 69.14 $\pm$0.09 70.93 $\pm$0.20 75.18 $\pm$0.14
Pixelate 48.48 67.08 $\pm$0.24 44.65 $\pm$0.21 71.80 $\pm$0.16 73.38 $\pm$0.29 75.67 $\pm$0.32 78.73 $\pm$0.30
JPEG Compression 56.05 65.42 $\pm$0.05 56.73 $\pm$0.07 66.42 $\pm$0.25 69.02 $\pm$0.10 69.65 $\pm$0.23 72.65 $\pm$0.18
Mean 60.15 68.00 59.75 73.22 75.04 76.22 80.11
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-10 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03 95.91 $\pm$0.10 95.71 $\pm$0.03 97.61 $\pm$0.03
CIFAR-10.1 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02 92.97 $\pm$0.13 92.10 $\pm$0.33 93.45 $\pm$0.30
Gaussian Noise 64.64 68.87 $\pm$0.20 64.44 $\pm$0.11 70.04 $\pm$0.31 72.81 $\pm$0.09 72.73 $\pm$0.03 80.47 $\pm$0.26
Shot Noise 67.82 71.95 $\pm$0.06 66.81 $\pm$0.19 71.44 $\pm$0.16 74.45 $\pm$0.16 74.60 $\pm$0.03 82.64 $\pm$0.29
Impulse Noise 78.21 80.22 $\pm$0.19 76.46 $\pm$0.17 79.42 $\pm$0.15 81.36 $\pm$0.09 80.95 $\pm$0.15 87.80 $\pm$0.15
Defocus Blur 80.73 83.10 $\pm$0.03 79.01 $\pm$0.23 81.75 $\pm$0.19 83.20 $\pm$0.10 83.15 $\pm$0.18 90.33 $\pm$0.12
Glass Blur 50.29 57.12 $\pm$0.07 49.64 $\pm$0.23 58.13 $\pm$0.23 61.51 $\pm$0.07 62.35 $\pm$0.15 77.72 $\pm$0.31
Motion Blur 80.75 82.69 $\pm$0.11 78.85 $\pm$0.04 80.76 $\pm$0.12 82.60 $\pm$0.13 82.61 $\pm$0.12 88.93 $\pm$0.17
Zoom Blur 82.75 84.91 $\pm$0.08 82.32 $\pm$0.13 83.39 $\pm$0.05 85.76 $\pm$0.06 85.44 $\pm$0.13 91.69 $\pm$0.13
Snow 83.01 85.99 $\pm$0.11 82.69 $\pm$0.10 84.48 $\pm$0.07 84.91 $\pm$0.13 85.61 $\pm$0.15 91.56 $\pm$0.11
Frost 84.90 87.15 $\pm$0.12 84.63 $\pm$0.08 85.21 $\pm$0.06 87.17 $\pm$0.13 86.88 $\pm$0.04 91.43 $\pm$0.12
Fog 78.44 81.30 $\pm$0.07 77.56 $\pm$0.17 79.27 $\pm$0.07 81.80 $\pm$0.11 81.79 $\pm$0.09 90.82 $\pm$0.11
Brightness 91.67 93.07 $\pm$0.04 90.94 $\pm$0.04 91.87 $\pm$0.09 92.78 $\pm$0.05 92.59 $\pm$0.16 95.86 $\pm$0.00
Contrast 84.20 87.93 $\pm$0.04 82.88 $\pm$0.09 86.19 $\pm$0.06 87.54 $\pm$0.12 87.38 $\pm$0.02 94.53 $\pm$0.08
Elastic Transform 65.45 69.96 $\pm$0.12 64.81 $\pm$0.14 67.43 $\pm$0.24 71.19 $\pm$0.07 71.25 $\pm$0.09 82.16 $\pm$0.23
Pixelate 75.10 77.89 $\pm$0.05 72.92 $\pm$0.12 77.11 $\pm$0.10 77.88 $\pm$0.13 77.67 $\pm$0.16 87.13 $\pm$0.16
JPEG Compression 72.58 75.49 $\pm$0.07 71.18 $\pm$0.19 74.46 $\pm$0.11 75.88 $\pm$0.16 75.84 $\pm$0.18 82.60 $\pm$0.18
Mean 76.04 79.18 75.01 78.06 80.05 80.06 86.35
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07 72.98 $\pm$0.07 72.85 $\pm$0.15 74.07 $\pm$0.20
Gaussian Noise 15.88 12.28 $\pm$0.20 15.43 $\pm$0.03 19.01 $\pm$0.24 34.23 $\pm$0.03 35.95 $\pm$0.27 38.80 $\pm$0.51
Shot Noise 17.49 15.07 $\pm$0.21 16.88 $\pm$0.07 20.27 $\pm$0.21 36.68 $\pm$0.10 37.96 $\pm$0.15 40.89 $\pm$0.10
Impulse Noise 21.43 13.13 $\pm$0.16 22.12 $\pm$0.15 17.66 $\pm$0.10 43.17 $\pm$0.35 44.62 $\pm$0.20 46.39 $\pm$0.17
Defocus Blur 40.10 50.35 $\pm$0.03 41.08 $\pm$0.22 49.86 $\pm$0.13 53.13 $\pm$0.12 53.80 $\pm$0.12 55.50 $\pm$0.24
Glass Blur 13.48 04.84 $\pm$0.14 18.43 $\pm$0.15 18.34 $\pm$0.31 32.53 $\pm$0.03 33.39 $\pm$0.11 37.89 $\pm$0.68
Motion Blur 39.82 49.85 $\pm$0.37 40.85 $\pm$0.26 50.00 $\pm$0.09 51.63 $\pm$0.06 52.72 $\pm$0.30 55.10 $\pm$0.28
Zoom Blur 45.45 54.76 $\pm$0.04 46.77 $\pm$0.06 54.13 $\pm$0.08 56.81 $\pm$0.11 57.51 $\pm$0.09 58.77 $\pm$0.33
Snow 42.77 52.38 $\pm$0.18 47.24 $\pm$0.18 52.80 $\pm$0.27 56.04 $\pm$0.06 56.73 $\pm$0.27 57.84 $\pm$0.13
Frost 45.39 51.66 $\pm$0.04 48.61 $\pm$0.14 49.56 $\pm$0.08 56.00 $\pm$0.11 56.48 $\pm$0.34 58.09 $\pm$0.04
Fog 38.98 50.74 $\pm$0.14 39.92 $\pm$0.16 49.92 $\pm$0.11 52.88 $\pm$0.33 53.83 $\pm$0.19 55.68 $\pm$0.38
Brightness 52.55 64.26 $\pm$0.09 55.83 $\pm$0.10 63.76 $\pm$0.13 65.58 $\pm$0.07 66.67 $\pm$0.19 67.76 $\pm$0.21
Contrast 33.32 48.69 $\pm$0.08 33.13 $\pm$0.16 47.86 $\pm$0.02 52.90 $\pm$0.06 55.06 $\pm$0.15 58.29 $\pm$0.51
Elastic Transform 24.39 33.56 $\pm$0.28 27.36 $\pm$0.10 32.93 $\pm$0.23 39.82 $\pm$0.21 40.37 $\pm$0.26 44.32 $\pm$0.18
Pixelate 21.89 36.20 $\pm$0.28 21.26 $\pm$0.10 39.49 $\pm$0.21 45.10 $\pm$0.06 47.02 $\pm$0.04 49.46 $\pm$0.17
JPEG Compression 27.21 30.80 $\pm$0.05 30.97 $\pm$0.10 35.56 $\pm$0.23 41.43 $\pm$0.18 42.13 $\pm$0.21 43.87 $\pm$0.23
Mean 32.01 37.90 33.73 40.08 47.86 48.95 51.24
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
CIFAR-100 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08 79.33 $\pm$0.05 78.85 $\pm$0.19 81.38 $\pm$0.15
Gaussian Noise 30.55 36.93 $\pm$0.03 36.10 $\pm$0.11 41.46 $\pm$0.15 43.99 $\pm$0.13 44.13 $\pm$0.11 52.26 $\pm$0.18
Shot Noise 34.58 40.96 $\pm$0.16 38.23 $\pm$0.13 44.27 $\pm$0.09 46.28 $\pm$0.22 46.63 $\pm$0.17 54.59 $\pm$0.26
Impulse Noise 44.89 49.09 $\pm$0.14 49.69 $\pm$0.21 51.44 $\pm$0.23 56.15 $\pm$0.04 56.26 $\pm$0.22 63.02 $\pm$0.12
Defocus Blur 48.88 55.23 $\pm$0.07 50.43 $\pm$0.19 56.55 $\pm$0.22 57.46 $\pm$0.01 57.66 $\pm$0.42 65.20 $\pm$0.17
Glass Blur 23.46 27.02 $\pm$0.23 24.35 $\pm$0.22 30.47 $\pm$0.14 32.54 $\pm$0.12 33.54 $\pm$0.16 48.19 $\pm$0.17
Motion Blur 50.83 56.03 $\pm$0.20 51.94 $\pm$0.04 56.98 $\pm$0.18 58.22 $\pm$0.10 57.81 $\pm$0.05 65.09 $\pm$0.17
Zoom Blur 56.02 61.19 $\pm$0.10 56.96 $\pm$0.16 62.56 $\pm$0.04 62.94 $\pm$0.02 62.74 $\pm$0.06 68.92 $\pm$0.15
Snow 49.03 55.60 $\pm$0.09 54.89 $\pm$0.11 58.81 $\pm$0.11 60.68 $\pm$0.06 61.04 $\pm$0.27 66.81 $\pm$0.15
Frost 53.27 58.21 $\pm$0.15 58.15 $\pm$0.33 60.38 $\pm$0.23 62.34 $\pm$0.14 62.76 $\pm$0.22 67.71 $\pm$0.29
Fog 48.51 53.37 $\pm$0.25 49.26 $\pm$0.13 54.38 $\pm$0.04 54.71 $\pm$0.31 54.70 $\pm$0.13 65.54 $\pm$0.23
Brightness 60.53 67.34 $\pm$0.17 66.60 $\pm$0.10 69.63 $\pm$0.14 71.52 $\pm$0.07 71.60 $\pm$0.09 76.25 $\pm$0.15
Contrast 50.24 59.91 $\pm$0.13 53.64 $\pm$0.24 63.39 $\pm$0.13 62.77 $\pm$0.22 63.95 $\pm$0.04 72.34 $\pm$0.27
Elastic Transform 35.07 38.49 $\pm$0.12 35.72 $\pm$0.09 39.57 $\pm$0.39 41.28 $\pm$0.25 41.27 $\pm$0.15 52.72 $\pm$0.24
Pixelate 43.86 48.37 $\pm$0.17 44.32 $\pm$0.10 50.45 $\pm$0.16 51.15 $\pm$0.15 51.22 $\pm$0.13 59.55 $\pm$0.06
JPEG Compression 39.11 44.42 $\pm$0.09 43.44 $\pm$0.11 47.45 $\pm$0.14 49.40 $\pm$0.17 49.78 $\pm$0.18 54.98 $\pm$0.20
Mean 44.59 50.14 47.58 52.52 54.10 54.34 69.48
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 58.12 63.59 63.87 $\pm$0.09 64.73 $\pm$0.11 66.05 $\pm$0.03 66.34 $\pm$0.07
Gaussian Noise 4.80 13.44 $\pm$0.21 10.53 $\pm$0.25 19.25 $\pm$0.18 16.78 $\pm$0.09 24.88 $\pm$0.28
Shot Noise 6.44 17.46 $\pm$0.20 12.90 $\pm$0.45 21.80 $\pm$0.11 19.27 $\pm$0.13 27.99 $\pm$0.32
Impulse Noise 4.36 9.09 $\pm$0.14 6.57 $\pm$0.20 13.46 $\pm$0.01 10.42 $\pm$0.07 20.69 $\pm$0.18
Defocus Blur 22.23 32.33 $\pm$0.09 29.16 $\pm$0.32 33.49 $\pm$0.29 $3$$\pm$0.03 37.27 $\pm$0.16
Glass Blur 7.50 12.47 $\pm$0.14 10.88 $\pm$0.18 18.33 $\pm$0.09 16.26 $\pm$0.08 22.72 $\pm$0.14
Motion Blur 33.46 44.17 $\pm$0.06 39.91 $\pm$0.43 45.17 $\pm$0.11 44.21 $\pm$0.10 48.75 $\pm$0.37
Zoom Blur 29.86 39.24 $\pm$0.26 36.62 $\pm$0.11 40.53 $\pm$0.19 40.26 $\pm$0.21 45.16 $\pm$0.08
Snow 31.64 38.91 $\pm$0.29 38.66 $\pm$0.19 42.71 $\pm$0.11 42.31 $\pm$0.18 45.74 $\pm$0.15
Frost 33.26 42.04 $\pm$0.15 40.95 $\pm$0.20 44.28 $\pm$0.14 44.83 $\pm$0.13 48.09 $\pm$0.05
Fog 24.26 31.06 $\pm$0.18 29.22 $\pm$0.24 34.78 $\pm$0.11 34.08 $\pm$0.13 41.21 $\pm$0.19
Brightness 40.48 49.26 $\pm$0.12 47.78 $\pm$0.44 51.61 $\pm$0.39 52.16 $\pm$0.08 55.75 $\pm$0.24
Contrast 1.72 3.06 $\pm$0.09 3.58 $\pm$0.07 7.24 $\pm$0.09 5.43 $\pm$0.08 13.06 $\pm$0.18
Elastic Transform 23.85 34.57 $\pm$0.05 32.11 $\pm$0.10 35.16 $\pm$0.05 35.11 $\pm$0.05 41.69 $\pm$0.12
Pixelate 22.35 37.73 $\pm$0.18 29.02 $\pm$0.22 41.11 $\pm$0.18 37.47 $\pm$0.21 45.83 $\pm$0.17
JPEG Compression 27.65 41.92 $\pm$0.45 36.51 $\pm$0.12 44.58 $\pm$0.04 42.93 $\pm$0.08 46.55 $\pm$0.26
Mean 20.92 29.78 26.96 32.90 31.66 37.69
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-ImageNet 71.34 74.09 $\pm$0.12 74.51 $\pm$0.20 74.57 $\pm$0.15 75.04 $\pm$0.21 75.01 $\pm$0.09
Gaussian Noise 17.26 26.60 $\pm$0.16 23.22 $\pm$0.12 29.27 $\pm$0.03 27.98 $\pm$0.14 38.14 $\pm$0.09
Shot Noise 21.57 28.93 $\pm$0.27 27.39 $\pm$0.21 33.01 $\pm$0.18 31.47 $\pm$0.04 41.02 $\pm$0.13
Impulse Noise 17.32 $2$$\pm$0.29 19.79 $\pm$0.18 27.93 $\pm$0.09 25.38 $\pm$0.15 35.48 $\pm$0.10
Defocus Blur 35.63 38.66 $\pm$0.19 39.44 $\pm$0.27 39.99 $\pm$0.15 40.55 $\pm$0.04 46.03 $\pm$0.17
Glass Blur 13.62 10.98 $\pm$0.13 15.72 $\pm$0.18 17.09 $\pm$0.15 17.44 $\pm$0.16 27.93 $\pm$0.18
Motion Blur 50.11 52.53 $\pm$0.01 53.91 $\pm$0.32 53.63 $\pm$0.28 54.46 $\pm$0.16 57.78 $\pm$0.06
Zoom Blur 43.65 46.02 $\pm$0.15 48.82 $\pm$0.11 47.39 $\pm$0.04 48.28 $\pm$0.09 52.67 $\pm$0.09
Snow 48.04 49.13 $\pm$0.33 51.50 $\pm$0.27 51.97 $\pm$0.14 52.60 $\pm$0.04 56.85 $\pm$0.13
Frost 49.36 50.46 $\pm$0.09 53.89 $\pm$0.22 51.49 $\pm$0.11 53.77 $\pm$0.16 57.56 $\pm$0.35
Fog 37.58 38.63 $\pm$0.15 40.43 $\pm$0.11 42.35 $\pm$0.04 42.57 $\pm$0.09 50.49 $\pm$0.09
Brightness 55.85 59.25 $\pm$0.04 61.07 $\pm$0.09 60.69 $\pm$0.12 62.01 $\pm$0.04 65.28 $\pm$0.31
Contrast 5.78 8.66 $\pm$0.14 7.48 $\pm$0.18 11.25 $\pm$0.08 9.21 $\pm$0.15 21.57 $\pm$0.13
Elastic Transform 34.15 36.81 $\pm$0.19 40.65 $\pm$0.27 38.44 $\pm$0.22 39.85 $\pm$0.08 47.86 $\pm$0.17
Pixelate 45.90 50.77 $\pm$0.18 52.28 $\pm$0.31 52.22 $\pm$0.06 51.83 $\pm$0.04 58.23 $\pm$0.29
JPEG Compression 50.13 51.88 $\pm$0.13 52.38 $\pm$0.24 52.46 $\pm$0.19 54.19 $\pm$0.16 56.82 $\pm$0.24
Mean 34.98 40.28 41.07 42.98 43.28 50.36
[…]
Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
Tiny-Imagenet 58.29 57.72 $\pm$0.14 58.90 $\pm$0.15 59.85 $\pm$0.01 61.35 $\pm$0.17 63.69 $\pm$0.15
Gaussian Noise 7.08 8.01 $\pm$0.05 9.29 $\pm$0.03 14.44 $\pm$0.25 13.02 $\pm$0.07 21.40 $\pm$0.37
Shot Noise 9.41 10.04 $\pm$0.07 11.70 $\pm$0.07 17.44 $\pm$0.12 15.94 $\pm$0.07 24.90 $\pm$0.17
Impulse Noise 3.44 4.18 $\pm$0.02 4.85 $\pm$0.01 10.37 $\pm$0.16 6.90 $\pm$0.01 17.34 $\pm$0.14
Defocus Blur 21.71 24.53 $\pm$0.05 27.56 $\pm$0.10 31.46 $\pm$0.26 29.91 $\pm$0.13 35.39 $\pm$0.16
Glass Blur 9.12 10.09 $\pm$0.06 11.03 $\pm$0.10 15.84 $\pm$0.09 14.01 $\pm$0.10 21.16 $\pm$0.19
Motion Blur 34.52 36.94 $\pm$0.07 38.97 $\pm$0.05 41.34 $\pm$0.20 41.26 $\pm$0.04 46.26 $\pm$0.37
Zoom Blur 27.44 29.48 $\pm$0.06 34.29 $\pm$0.07 35.06 $\pm$0.15 33.96 $\pm$0.07 40.93 $\pm$0.05
Snow 32.51 32.20 $\pm$0.03 34.45 $\pm$0.10 36.79 $\pm$0.08 37.76 $\pm$0.09 42.32 $\pm$0.14
Frost 36.33 35.72 $\pm$0.02 37.13 $\pm$0.05 38.37 $\pm$0.13 39.65 $\pm$0.03 44.97 $\pm$0.17
Fog 25.94 27.46 $\pm$0.04 28.89 $\pm$0.08 33.51 $\pm$0.19 32.13 $\pm$0.08 38.60 $\pm$0.32
Brightness 40.15 39.79 $\pm$0.08 43.31 $\pm$0.10 46.52 $\pm$0.06 46.93 $\pm$0.13 53.10 $\pm$0.05
Contrast 1.81 2.24 $\pm$0.04 3.15 $\pm$0.04 6.07 $\pm$0.12 3.53 $\pm$0.02 11.88 $\pm$0.23
Elastic Transform 30.40 31.92 $\pm$0.05 33.88 $\pm$0.14 33.74 $\pm$0.11 35.01 $\pm$0.13 40.73 $\pm$0.16
Pixelate 22.78 24.79 $\pm$0.03 27.70 $\pm$0.12 34.84 $\pm$0.09 31.55 $\pm$0.08 41.84 $\pm$0.09
JPEG Compression 29.59 30.93 $\pm$0.13 33.60 $\pm$0.15 37.29 $\pm$0.14 36.46 $\pm$0.11 42.86 $\pm$0.04
Mean 22.14 23.22 25.32 28.87 27.87 34.91
[…]

### v3

@@
-Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation
+Semantic Anchor Transport: Robust Test-Time Adaptation for Vision-Language Models
 Abstract
-Vision-language foundation models, such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks.
+Large pre-trained vision-language models (VLMs), such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks.
 Nevertheless, these models may be unreliable under distributional shifts, as their performance is significantly degraded.
-In this work, we explore how to efficiently leverage class text information to mitigate these distribution drifts encountered by large pre-trained vision-language models (VLMs) during test-time inference.
-In particular, we propose to generate pseudo-labels for the test-time samples by exploiting generic class text embeddings as fixed centroids of a label assignment problem, which is efficiently solved with Optimal Transport.
-Furthermore, the proposed adaptation method (CLIP-OT) integrates a multiple template knowledge distillation approach, which replicates multi-view contrastive learning strategies in unsupervised representation learning but without incurring additional computational complexity.
-Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of CLIP-OT, achieving performance gains of up to $7\%$ over recent state-of-the-art methods, yet being computationally and memory efficient [^1] Code available at https://github.com/ShambhaviCodes/CLIPOT.
+In this work, we investigate how to efficiently utilize class text information to mitigate distribution drifts encountered by VLMs during inference.
+In particular, we propose generating pseudo-labels for the noisy test-time samples by aligning visual embeddings with reliable, text-based semantic anchors.
+Specifically, to maintain the regular structure of the dataset properly, we formulate the problem as a batch-wise label assignment, which is efficiently solved using Optimal Transport.
+Our method, Semantic Anchor Transport (SAT), utilizes such pseudo-labels as supervisory signals for test-time adaptation, yielding a principled cross-modal alignment solution.
+Moreover, SAT further leverages heterogeneous textual clues, with a multi-template distillation approach that replicates multi-view contrastive learning strategies in unsupervised representation learning without incurring additional computational complexity.
+Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of SAT, achieving consistent performance gains over recent state-of-the-art methods, yet being computationally efficient.
 1 Introduction
-Large pre-trained vision-language models (VLMs), such as CLIP [33] and ALIGN [19], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability.
-Nevertheless, if the model is exposed to domain drifts between the training and test samples, its performance can be largely degraded [44, 37].
-While a straightforward solution to bridge this gap consists in fine-tuning the trained model using domain-specific labeled data [22, 13], this strategy brings several limitations in real-world scenarios, which may hamper their scalability and deployment.
+Large pre-trained vision-language models (VLMs), such as CLIP [radford2021learning] and ALIGN [jia2021scaling], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability.
+Nevertheless, if the model is exposed to domain drifts at test time, its performance can be largely degraded [yu2023task, silva2024closer].
+While a straightforward solution to bridge this gap involves fine-tuning the trained model using domain-specific labeled data [lai2023padclip, goyal2023finetune], this strategy presents several limitations in real-world scenarios, which may hinder its scalability and deployment.
 First, adapting to new domains requires collecting labeled samples drawn from each distinct distribution.
 This might be impractical for specific domains and further hinder the real-time adaptation of the trained model for each input test sample.
-Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [40].
-Test-Time Adaptation (TTA) presents a realistic and practical scenario of unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [39, 17, 45].
+Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [WiSE].
+Test-Time Adaptation (TTA) presents a realistic and practical scenario for unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [wangtent, iwasawa2021test, yuan2023robust].
 Nevertheless, despite the recent rise of VLMs, and the popularity of TTA in more traditional deep models, such as CNNs and ViTs, the study of TTA in large pre-trained vision-language models remains less explored.
-Initial approaches, such as TENT [39], which simply minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [36].
-In contrast, more recent strategies resort to pseudo-labels on the inference batch to guide the model adaptation [31].
+Standard approaches, e.g., TENT [wangtent], which minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [shu2022test].
+On the other hand, more recent strategies utilize pseudo-labels within the inference batch to guide model adaptation [osowiechi2024watt, hakim2024clipartt, maharana2024texttt].
 In particular, the core idea of these later methods is to minimize a classification loss, typically the standard cross-entropy, between the generated pseudo-labels and model predictions, which guides the network updates over multiple iterations.
-Nonetheless, while this strategy is common in semi-supervised learning, where labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions [17].
-In light of the above limitation, we propose a novel approach for CLIP test-time adaptation based on the simple observation that this task can draw inspiration from clustering.
-Nevertheless, instead of jointly learning the cluster centroids during training, these can be replaced by the fixed class prototypes derived from the text representations.
-By doing this, the unlabeled samples in a batch will be assigned to the most similar class prototypes, guided by the text embeddings, which serve as a strong class guidance without requiring explicit sample-wise supervision.
-Furthermore, as our primary goal is to match distributions between different modalities, i.e., class text embeddings and features from the test images, traditional prototype-based clustering methods may lead to suboptimal performance, as they typically assume an unimodal distribution for each cluster.
-In contrast, we frame the label assignment problem as an Optimal Transport problem, as it can more naturally model multi-modal distributions [23], and is efficiently solved using the Sinkhorn algorithm [9].
-Our key contributions can be summarized as follows:
-  • We propose to cast the pseudo-labeling strategy in CLIP test-time adaptation as an Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed cluster centroids without requiring further annotations.
+Nonetheless, while this strategy is common in semi-supervised learning, where additional labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions over multiple updates [iwasawa2021test], a problem we refer to as error accumulation.
+Figure 1: Error Accumulation.
+We track examples from all corruptions from CIFAR10C dataset with initial zero-shot CLIP predictions (at $t=0$) that are misclassified.
+Baselines [wangtent, maharana2024texttt, osowiechi2024watt] catastrophically reinforce this error: similarity to the ‘Wrong Class’ (dashed lines) increases while similarity to the ‘True Class’ (solid lines) decreases.
+In contrast, our method, SAT (orange), is the only one that provides a corrective signal, actively reducing similarity to the wrong class and breaking the cycle of error accumulation.
+This problem is particularly prevalent in the literature on test-time adaptation in VLMs.
+To illustrate this phenomenon, we conducted a simple experiment, the findings of which are depicted in Figure 1.
+Considering all corruptions in CIFAR10C, we identify samples misclassified by zero-shot CLIP and track their behavior across consecutive TTA steps.
+Specifically, we compute the average cosine similarity between the adapted visual embeddings and the corresponding class text embeddings of both the wrong (i.e., predicted) and true (i.e., expected) category.
+Ideally, a robust TTA method should reduce the similarity to the wrong class prototype, gradually correcting the initial misalignment.
+Nevertheless, as shown in Figure 1, existing methods tend to reinforce these mistakes, thereby increasing the similarity between the visual and text prototypes of the incorrect class, even when the prediction is wrong.
+In light of the above limitation, we aim to provide a more robust TTA framework that can correct incorrect initial supervisory signals.
+To do so, we reframe TTA as a cross-modal alignment problem, where the goal is to align the embeddings of test images to their corresponding fixed semantic anchors provided by the language modality.
+This alignment can be viewed through the lens of clustering, where the unlabeled samples are grouped around class text prototypes.
+Hence, we can take into account the structure of the joint test batch to better correct sample-specific inconsistencies.
+Nevertheless, instead of jointly learning the cluster centroids during training, we use the fixed class prototypes derived from the text representations, which provide stable semantic guidance without requiring explicit supervision.
+A key challenge is the distribution mismatch between text prototypes and test image embeddings, since they originate from different modalities.
+Traditional clustering methods, which typically assume unimodal distributions, often struggle to bridge this gap, resulting in suboptimal performance.
+In contrast, we formulate label assignment as an Optimal Transport (OT) problem, yielding a global, cost-aware correspondence between visual embeddings and text prototypes that naturally model multi-modal distributions [multimodal_dist_ot].
+This perspective is more robust to outliers or noisy predictions, directly mitigating the error accumulation that undermines existing self-training in TTA.
+The main contributions of this paper can be summarized as:
+  • We propose Semantic Anchor Transport (SAT).
+This novel framework casts the pseudo-labeling strategy in CLIP test-time adaptation as a batch-wise Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed robust cluster centroids without requiring further annotations.
   • To solve the label assignment task, we resort to the Sinkhorn algorithm, as it can handle multi-modal distributions and compute label assignments efficiently.
-  • We introduce a multi-template knowledge distillation approach, which leverages richer information derived from different text prompts to better guide the adaptation without incurring significant computational overhead.
-  • Experiments across 240 scenarios demonstrate the superiority of CLIP-OT over recent state-of-the-art methods, all while avoiding additional computational complexity.
+  • We introduce a multi-template knowledge distillation approach that leverages richer information derived from different text prompts to better guide adaptation without incurring significant computational overhead.
+  • Comprehensive experiments across multiple visual corruptions and domain shifts benchmarks demonstrate the superiority of SAT over recent state-of-the-art methods.
 2 Related Work
-Test-Time Adaptation.
-TTA adapts a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [39, 47, 8, 30].
-Existing approaches for TTA in more traditional deep models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [27, 35]; ii) entropy-based approaches [39, 29, 12], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [25], which employ the test-time generated labels for supervising the model.
-With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt large pre-trained foundation models, particularly CLIP [36, 26, 31, 14], which mainly differs on the different parameter groups updated during adaptation, i.e., prompt tuning [36, 26, 10] or layer-norm [14, 31] strategies.
-Test-Time Prompt Tuning (TPT) [36] fine-tunes only the input text prompts, which are optimized by minimizing the entropy of the model predictions.
-To achieve this, TPT generates multiple random augmented views for each test image, incurring additional costs due to performing multiple forward passes through the CLIP vision encoder for each image.
-Furthermore, TPT is a unimodal approach, which restricts a multi-modal model such as CLIP from fully exploiting its multi-modal capabilities for adaptation.
-Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to the CLIP text encoder, where both text and visual encoders remain frozen.
-Nonetheless, keeping the vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively.
-Most recently, CLIPArTT [14] and WATT [31] propose to fine-tune the normalization layers instead, resorting to pseudo-labels computed from image-to-image and text-to-text similarities.
+Test-Time Adaptation (TTA) aims at adapting a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [wangtent, zhang2022memo, choi2022improving, niutowards].
+Existing approaches in traditional unimodal models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [mirza2022norm, schneider2020improving]; ii) entropy-based approaches [wangtent, niu2022efficient, goyal2022test], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [liang2020we], which employ the test-time generated labels for supervising the model.
+With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt pre-trained foundation models, notably CLIP [shu2022test, ma2024swapprompt, osowiechi2024watt, hakim2024clipartt], which mainly differ on the different parameter groups updated during adaptation, i.e., prompt tuning [shu2022test, ma2024swapprompt, dobler2024lost] or layer-norm [hakim2024clipartt, osowiechi2024watt] strategies.
+For example, Test-Time Prompt Tuning (TPT) [shu2022test] optimizes input text prompts by minimizing the model’s prediction entropy.
+In contrast, Vision-Text-Space Ensemble (VTE) [dobler2024lost] uses an ensemble of prompts as input to the text encoder.
+Nevertheless, keeping both the text and vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively.
+Recent VLM adaptation methods, such as WATT [osowiechi2024watt] and BATCLIP [maharana2024texttt], which employ parameter-efficient updates of the encoders, have developed sophisticated bimodal heuristics to improve pseudo-label quality.
+WATT [osowiechi2024watt] generates them by combining both image-to-image and text-to-text similarity matrices.
+Taking this a step further, BATCLIP [maharana2024texttt] first generates pseudo-labels via a standard argmax prediction, which are refined via additional losses.
+Nevertheless, both approaches rely on locally derived pseudo-labels, failing to explicitly optimize the sample-to-class assignment group cost, making them more sensitive to outliers or ambiguous samples.
 Clustering for unsupervised representation learning.
-Jointly learning how to adapt a set of learnable parameters of a deep network and identify the class assignments can be viewed as clustering and unsupervised representation learning.
-Thus, our work is closely related to recent literature on deep clustering-based approaches [1, 3, 4, 16, 42, 43, 18, 5].
-In [3], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, strategy that is later employed in [4] to pre-train standard supervised deep models.
+Jointly adapting the parameters of a deep network while inferring the class assignments can be viewed as clustering and unsupervised representation learning.
+Thus, our work is closely related to recent literature on deep clustering-based approaches [asanoself, caron2018deep, caron2019unsupervised, huang2019unsupervised, yan2020clusterfit, yang2016joint, jabi2019deep, caron2020unsupervised].
+In [caron2018deep], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, a strategy later employed in [caron2019unsupervised] to pre-train standard supervised deep models.
 Nevertheless, applying naive $k$-means risks collapsing to only a few imbalanced clusters, making it ineffective.
-The label assignment task is framed as an instance of the Optimal Transport problem in [1], which stresses that the formulation presented in [3] is not principled, since potential degenerate solutions are avoided via specific implementation choices.
-Caron et al.
-[5] propose to enforce consistency between cluster assignments produced for different image views, or augmentations, which avoids expensive pairwise comparisons typically performed in contrastive learning.
-While we consider a similar formulation to some of these works, several vital differences exist.
-In particular, all these works, e.g., [4, 1, 3] are used in unsupervised representation learning, and thus the prototypes are learned from the data distribution.
-In contrast, our approach leverages the text representations to enhance the cluster assignments.
+A more principled approach consists in framing the label assignment task as an instance of the Optimal Transport problem [asanoself], whose global, batch-aware optimization and balancing constraints naturally prevent degenerate solutions.
+Furthermore, Caron et al.
+[caron2020unsupervised] enforce consistency between cluster assignments across different image views or augmentations, avoiding expensive pairwise comparisons typically performed in contrastive learning.
+While our formulation shares similarities, key differences exist.
+In particular, all these methods, i.e., [caron2019unsupervised, asanoself, caron2018deep] operate in unsupervised representation learning, where prototypes are learned from the data distribution.
+In contrast, we leverage the text representations as fixed semantic anchors.
 Note that text embeddings are accessible for free in VLMs, not incurring additional supervision, as the test image label remains unknown.
-A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [5], which introduce a computational burden.
-Indeed, we can leverage individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring additional costs.
-In fact, these embeddings are computed only once offline.
-In contrast, image augmentation or “multi-views” strategies must create the additional images for each incoming data batch, which means they perform one forward pass per new augmentation at test time.
+A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [caron2020unsupervised], which introduce a computational burden.
+Instead, we treat individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring extra cost.
+In fact, these embeddings are computed only once offline, whereas image augmentation or “multi-views” strategies must create the additional images for each incoming batch, performing one forward pass per new augmentation at test time.
 3 Background
 3.1 Vision-language models
-[…]
+CLIP [radford2021learning] is a foundation vision-language model, trained via contrastive learning to produce visual representations from images $\mathbf{x}$ paired with their associated text descriptions $T$.
+To do so, CLIP consists of an image encoder $\boldsymbol{\theta}$ and a text encoder $\boldsymbol{\phi}$.
+This generates the corresponding vision $\mathbf{z}\in\real^{d}$ and class text $\mathbf{t}_{k}\in\real^{d}$ embeddings (column vectors), which are typically projected into an $\ell_{2}$-normalized shared embedding space.
+At inference, this learning paradigm enables zero-shot prediction.
+More concretely, for a given set of $K$ classes, and an ensemble of $M$ different templates, we can generate the set of available prompts as $\mathcal{T}=\{\{T_{mk}\}_{m=1}^{M}\}_{k=1}^{K}$, whose embedding for template $m$ and class $k$ is obtained as $\mathbf{t}_{mk}=\boldsymbol{\phi}(\textsc{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}``A photo of a [class${}_{k}$]''}})$.
+Then, a popular strategy [radford2021learning, clipad, WiSE] consists in obtaining a class zero-shot prototype, which is computed as $\mathbf{t}_{k}=\frac{1}{M}\sum^{M}_{m=1}\boldsymbol{\phi}(T_{mk})$.
+Then, for a given test image $\mathbf{x}_{i}$, its zero-shot prediction, $\mathbf{p}_{i}=(p_{ik})_{1\leq k\leq K}$, can be obtained as:
 $p(y=k|\mathbf{x}_{i})=\frac{\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{k}/\tau)}}{\sum_{j=1}^{K}\exp{(\mathbf{z}_{i}^{\top}\mathbf{t}_{j}/\tau)}},$ (1)
-where $\tau$ is a temperature parameter, whose value is learned during pre-training [33], and $\mathbf{z}^{\top}\mathbf{t}$ is the dot product operator.
+where $\tau$ is a temperature parameter learned during pre-training [radford2021learning], and $\mathbf{z}^{\top}\mathbf{t}$ indicates dot product.
 3.2 Test-Time Adaptation using CLIP
@@
 Inspired by the semi-supervised learning literature, a straightforward solution to leverage the predicted probability in Eq.
-1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [39]:
-$\mathcal{L}(p)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
-with $N$ denoting the size of the test set, or batch.
-However, relying on the Shannon entropy [39] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq.
+1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [wangtent]:
+$\mathcal{L}(p)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
+with $B_{T}$ denoting the size of the test batch.
+However, relying on the Shannon entropy [wangtent] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq.
 (2) might be trivially minimized by assigning all data points to a single, arbitrary label.
-4 Proposed solution
+4 SAT: TTA as Cross-Modal Alignment
 The proposed approach aims to overcome the limitations in Eq.
 2 objective by finding reliable pseudo-labels in a batch of test-time samples without supervision.
-To achieve this, we propose CLIP-OT.
-This novel TTA method aims to better leverage the text embeddings generated by the frozen CLIP text encoder, serving as strong class descriptors to yield label assignments for test samples through Optimal Transport (OT).
-Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [3], where cluster centroids are iteratively updated based on the prior label assignments.
+To achieve this, we propose SAT.
+We formulate TTA as a cross-modal alignment problem, where the core task is to generate a robust supervisory signal for each unlabeled test batch.
+This novel TTA method leverages the text embeddings generated by the frozen CLIP text encoder more effectively, serving as strong class descriptors to yield label assignments for test samples, which we propose propagating through Optimal Transport.
+Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [caron2018deep], where cluster centroids are iteratively updated based on the prior label assignments.
+Furthermore, by solving a globally optimal matching problem, SAT is more robust to noisy predictions than existing TTA methods [osowiechi2024watt, maharana2024texttt].
 Below, we elaborate on the details of our method, which is illustrated in Figure 2.
-Learning objective.
+Figure 2: SAT leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq.
+9).
+Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq.
+3).
+Note that, at each test batch, our model runs for $m$ iterations, each using a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
 To overcome the limitation of naively minimizing Eq.
-2 in the fully unlabeled scenario, we first encode the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
-$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
-4.1 Online pseudolabels assignments
+2 in the fully unlabeled scenario, TTA literature typically encodes the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
+$\mathcal{L}(p,q)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
+Thus, the adaptation is driven by the pseudo cross-entropy in Eq.
+(3), where a significant challenge lies in constructing a high-quality target distribution, or pseudo-label $q_{ik}=q(y=k|\mathbf{x}_{i})$, without access to ground truth labels.
+Indeed, as illustrated in Figure 1, prior TTA methods tend to amplify prediction errors, leading to a state of error accumulation.
+Our central thesis is that we can break this loop by exploiting the natural alignment between visual features and their text-based semantic anchors in the batch.
+4.1 Batch-Aware Cross-modal Alignment
 In the following, we detail how to encode robust posteriors into Eq.
@@
 3 can be expressed as:
-$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
+$\mathcal{L}(p,q)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
 where $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$ represents the matrix containing the class text prototypes, and the logarithmic term in the right side does not depend on $\mathbf{q}$.
 If we now express Eq.
-4 in its matrix form over all the test images, $\mathbf{Z}$, and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
-$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}),$ (5)
-where, following [9], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
-$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in^{K\times N}_{+}\mid\mathbf{Q}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\},$ (6)
+4 in its matrix form over all the test images in the batch (with $\mathbf{Z}\in\mathbb{R}^{d\times B_{T}}$ the corresponding embeddings), and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
+$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}).$ (5)
+which seeks the optimal alignment between the batch’s visual features $\mathbf{Z}\in\mathbb{R}^{d\times B_{T}}$ and a set of text-based “semantic anchors” $\mathbf{T}\in\mathbb{R}^{d\times K}$.
+Following [cuturi2013sinkhorn], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
+$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in\real^{K\times B_{T}}_{+}\mid\mathbf{Q}\mathbf{1}_{B_{T}}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{B_{T}}\mathbf{1}_{B_{T}}\},$ (6)
 with $\mathbf{1}_{K}$ and $\mathbf{1}_{N}$ denoting the vectors of ones in dimension $K$ and $N$, respectively.
 The constraints on $\mathcal{Q}$ enforce that on average each prototype is selected at least $\frac{N}{K}$ times, encouraging $\mathbf{Q}$ to be a matrix with uniform marginals in both rows and columns.
-Note that such constraints “break” the direct dependency between the original prediction $p$ in Eq.
+Note that such constraints ‘break’ the direct dependency between the original prediction $p$ in Eq.
 2 and the posterior $q$ in Eq.
-4, since the latter consider the whole batch distribution and its marginal properties for the assignment.
+4, since the latter considers the whole batch distribution and its marginal properties for the assignment.
 This results in more robust supervisory signals, later demonstrated empirically (Figure 3).
-Figure 1: CLIP-OT.
-Our method leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq.
-9).
-Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq.
-3).
-Note that, at each test batch, our model runs for $m$ iterations, each leveraging a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
-Codes optimization.
+Finding optimal Q.
 As the objective function in Eq.
@@
 Nevertheless, directly optimizing the above learning objective might be time-consuming, particularly as the number of data points and classes increases.
-To address this issue and allow a faster optimization, we apply the Sinkhorn algorithm [9], which integrates an entropic constraint, enforcing a simple structure on the optimal regularized transport.
+To address this issue and facilitate faster optimization, we apply the Sinkhorn algorithm [cuturi2013sinkhorn], which incorporates an entropic constraint that enforces a simple structure on the optimal regularized transport.
 Hence, the optimization problem becomes:
@@
 where $\mathcal{H}(\mathbf{Q})$ denotes the Shannon entropy function $\mathcal{H}(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ and $\varepsilon$ controls its weight.
-Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [5], in contrast to other works that employ the full dataset [1].
-Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{s}$, where $B_{s}$ denotes the batch size.
-Now, the soft codes $\mathbf{Q}^{*}$ are the solution of the problem presented in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations as:
+Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [caron2020unsupervised], in contrast to other works that employ the full dataset [asanoself].
+Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{T}$, where $B_{T}$ denotes the batch size.
+Now, the soft assignment matrix $\mathbf{Q}^{*}$ is the solution of the problem in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations:
 $\mathbf{Q}^{*}=\text{Diag}(\mathbf{u}^{(t)})\exp\left(\frac{\mathbf{T}^{\top}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}),$ (8)
-where $\mathbf{u}$ and $\mathbf{v}$ are renormalization vectors in K and B respectively, with $t$ denoting the iteration.
-Optimizing codes for individual textual views.
-Very recent literature [31] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$.
-Thus, we modify the solution presented in Eq.
-8 to be specialized for specific class template [^2] The text templates are the same as in WATT [31] – see Appendix A., where the codes for each template $m$ are:
+with $\mathbf{u}$ and $\mathbf{v}$ representing renormalization vectors in K and $\real{}^{B_{T}}$ respectively, and $t$ the iteration.
+Benefits of the proposed strategy.
+The proposed Optimal Transport (OT) based formulation provides: i) Global consistency, since OT is batch-aware.
+The assignment for any single image is dependent on the features of all other images in the batch.
+The solver finds the best overall configuration, making it inherently robust to individual outliers.
+ii) Intrinsic regularization given by the polytope constraints (Eq.
+6), which enforce that, on average, all classes are represented in the pseudo-labels.
+This structurally prevents the model from collapsing to a single class, yielding robustness.
+4.2 Semantic alignment with multiple templates
+Recent literature [osowiechi2024watt] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$.
+Thus, to create a richer and more robust supervisory signal, we leverage the ensemble of $M$ diverse text templates typically used in TTA, each providing a unique “semantic view”, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$, where $\mathbf{T}_{m}$ denotes the matrix containing the class text prototypes obtained from the $m$-th template.
+Compared to existing TTA approaches, we introduce a more sophisticated knowledge distillation strategy that disentangles the goals of assignment and generalization.
+1
+For Assignment, we need specificity.
+We compute a separate, “clean” transport plan $\mathbf{Q}^{*}_{m}$ for each specific semantic view $\mathbf{T}_{m}$.
+This provides an unambiguous matching based on a single, clear context (e.g., aligning to “a sketch of a class” vs.
+“a photo of a class”).
+Concretely, we modify the solution presented in Eq.
+8 to be specialized for specific class template [^1] The text templates are the same as in WATT [osowiechi2024watt] – see Appendix A., where the codes for each template $m$ are:
 $\mathbf{Q}^{*}_{m}=\text{Diag}(\mathbf{u}^{(t)}_{m})\exp\left(\frac{\mathbf{T}^{\top}_{m}\mathbf{Z}}{\lambda}\right)\text{Diag}(\mathbf{v}^{(t)}_{m})$. (9)
-The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm, [20], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$.
-Furthermore, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$ denotes the matrix containing the class text prototypes obtained from the $m$-th template.
-Note that the computational burden introduced by OT is negligible, representing roughly $1\%$ of the total runtime for a batch, as discussed in Section 5.3.
-4.2 Updating CLIP through textual codes
-Distilling knowledge from multiple text prototypes.
-To optimize the objective in Eq.
-3, we employ the soft label assignments, as they yield superior performance compared to their hard counterpart in other problems [5].
-Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions per category to the model, the cross-entropy in Eq.
-3 is optimized $M$ times, each one using a different template $m$, which can be defined for each image and template as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template as in Eq.
-1.
+The renormalization vectors in our setting are computed trough a small number of matrix multiplications using the iterative Sinkhorn-Knopp algorithm [knight2008sinkhorn], where in each iteration $\mathbf{u}^{(t)}_{m}=\mathbf{u}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{v}^{t-1}_{m})$ and $\mathbf{v}^{(t)}_{m}=\mathbf{v}_{m}/((\exp(\mathbf{T}_{m}^{\top}\mathbf{Z}/\lambda)\mathbf{u}^{t}_{m})$, with $\mathbf{v}^{0}_{m}=\mathbf{1}$.
+2
+For Generalization, we need consistency.
+To ensure generalization, model predictions must remain consistent across semantic views, preventing the visual encoder from overfitting to any single template, or “view”.
+In the proposed adaptation strategy, the model’s prediction $p(y=k|\mathbf{x}_{i})$ in Eq.
+1 is always computed using a single classification head, defined by the averaged text class prototypes $\mathbf{T}$.
+This design enforces stability across templates, ensuring that the pseudo-supervision $q(y=k|\mathbf{x}_{i})$ is always applied against a consistent prediction function, preventing the encoder from overfitting to template‑specific fluctuations and promoting robust generalization.
+Coupling both together.
+First, to optimize the objective in Eq.
+3, we employ the soft label assignments from (9), as they yield superior performance compared to their hard counterpart in other problems [caron2020unsupervised].
+Now, to distill the knowledge of the more diverse, and richer representations derived from multiple text predictions, the cross-entropy in Eq.
+3 is optimized by iterating through each of the $M$ views, which can be formally defined for each image $i$ and template $m$ as $\ell(\mathbf{p}_{i},\mathbf{q}^{*}_{im})=-\mathbf{q}^{*}_{im}\log\mathbf{p}_{i}$, where $\mathbf{p}_{i}$ is obtained with the average text template.
 Thus, for a given test image $i$ in a batch, after each update produced by the $m$-th text template, the visual encoder produces novel visual embeddings $\mathbf{z}_{i}$, which are then used to generate new predictions $\mathbf{p}_{i}$ for $i$.
-Whole learning strategy.
-The final proposed algorithm consists in updating the affine parameters in Layer Normalization, such as in [39, 31, 14], and learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq.
-(7).
-This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq.
-(9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given the current label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq.
+4.3 Whole learning strategy
+The complete Semantic Anchor Transport (SAT) algorithm (Alg.
+1) consists in learning a label assignment matrix $\mathbf{Q}_{m}$ per text template by solving the optimization problem presented in Eq.
+(7), and then updating the affine parameters of the visual encoder’s Layer Normalization layers, such as [wangtent, osowiechi2024watt, hakim2024clipartt].
+This is done iteratively by alternating two steps across the batches and text templates: i) with the learnable parameters fixed, we compute the visual features $\mathbf{Z}\in\real^{d\times B_{s}}$ of the test batch samples, and find $\mathbf{Q}_{m}^{*}$ through Eq.
+(9) by iterating the updates on $\mathbf{u}_{m}$ and $\mathbf{v}_{m}$; and ii) given label assignments $\mathbf{Q}_{m}^{*}$, and the softmax predictions for the test samples obtained with each average class prototype $\mathbf{t}_{k}$, the set of learnable parameters is updated by minimizing Eq.
 (3) w.r.t.
@@
 Finally, the class predictions on the batch are inferred with the updated model.
-A view of the proposed method is detailed in Algorithm 1.
-Thus, in a nutshell, we generate multiple pseudo-labels per sample, each derived from one text template.
-Note that pseudo-labeling is a common strategy in other fields, such as self-supervised [49] or semi-supervised [2, 46] learning.
-Thus, while $\mathbf{Q}$, or pseudo-labels in general, may contain errors, using these to refine the model is a standard practice, which typically yields consistent improvements.
-We refer the reader to Appendix G for a detailed discussion of the benefits of our approach compared to pseudo-labeling in other learning scenarios.
-Interpretation.
-Given a batch of $B_{s}$ visual embeddings $\mathbf{Z}=[\mathbf{z}_{1},...,\mathbf{z}_{B}]$, we aim at optimizing a function to map these embeddings into the class prototypes $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$.
-If we let $\mathbf{Q}$ denote the codes (or mapping), our goal is optimizing $\mathbf{Q}$ to maximize the similarity between the image features and the class text prototypes.
-While in fully-supervised, or semi-supervised regimes, labeled data is available to generate the class prototypes, in our Test-Time Adaptation scenario this becomes more challenging, as no label information about the test images is provided.
-Nevertheless, in the context of large vision-language models, the text can be leveraged as a form of supervision to create the prototypes $\mathbf{T}$, which represent a robust representation of each class.
-Furthermore, the entropic constraint term controls the smoothness of the mapping, promoting equally distributed predictions over all classes in each batch, which avoids degenerate solutions.
-Lastly, resorting to multiple templates individually, instead of averaging their embeddings, fully leverage the diverse information present in each text prompt.
-Algorithm 1 CLIP-OT adaptation for one corruption.
-1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$).
-// Split $\mathcal{X}$ into $B$ batches of size $B_{s}$.
-// Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
-2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
-3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
-4:     for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do    // Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
-5:         $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
-6:         $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
-7:         Compute codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
-8:         // Step 2: Distill knowledge to the model.
-9:         $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}\quad(\mathbf{t}_{k}\in\mathbb{R}^{d})$ // average class text embeddings over all M templates.
-10:         $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{t}_{k},\forall k$, Eq. (1).
-11:         Min. cross-entropy with $\mathbf{P}$ and $\mathbf{Q}^{*}_{m}$ // Eq. (3).
-12:         $\theta_{\text{LN}}^{(m-1)}\rightarrow\theta_{\text{LN}}^{(m)}$ // Update layer norm (LN) of $\boldsymbol{\theta}$.
-13:     end for // 2 - Inference (for all images in the batch).
-14:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
-15:     $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
-16:end for
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
-CIFAR-10 88.74 91.69 $\pm$0.10 88.06 $\pm$0.06 90.04 $\pm$0.13 91.41 $\pm$0.17 91.05 $\pm$0.06 93.15 $\pm$0.10 +2.10 $\blacktriangle$
-CIFAR-10.1 83.25 87.60 $\pm$0.45 81.80 $\pm$0.27 86.35 $\pm$0.27 87.78 $\pm$0.05 86.98 $\pm$0.31 88.37 $\pm$0.10 +1.39 $\blacktriangle$
-CIFAR-10C Gaussian Noise 35.27 41.27 $\pm$0.27 33.90 $\pm$0.08 59.90 $\pm$0.36 61.89 $\pm$0.24 63.84 $\pm$0.24 64.85 $\pm$0.14 +1.01 $\blacktriangle$
-Shot Noise 39.67 47.20 $\pm$0.23 38.20 $\pm$0.02 62.77 $\pm$0.07 63.52 $\pm$0.08 65.28 $\pm$0.21 67.34 $\pm$0.21 +2.06 $\blacktriangle$
-Impulse Noise 42.61 48.58 $\pm$0.31 37.66 $\pm$0.20 56.02 $\pm$0.16 57.13 $\pm$0.02 58.64 $\pm$0.11 62.27 $\pm$0.43 +3.63 $\blacktriangle$
-Defocus Blur 69.76 77.12 $\pm$0.16 67.83 $\pm$0.28 76.74 $\pm$0.05 78.86 $\pm$0.09 78.94 $\pm$0.12 82.09 $\pm$0.05 +3.15 $\blacktriangle$
-Glass Blur 42.40 52.65 $\pm$0.30 38.81 $\pm$0.12 61.77 $\pm$0.16 62.88 $\pm$0.06 65.12 $\pm$0.07 68.07 $\pm$0.02 +2.95 $\blacktriangle$
-Motion Blur 63.97 71.25 $\pm$0.09 63.39 $\pm$0.13 76.01 $\pm$0.19 76.85 $\pm$0.26 77.81 $\pm$0.14 81.30 $\pm$0.18 +3.49 $\blacktriangle$
-Zoom Blur 69.83 76.20 $\pm$0.19 68.95 $\pm$0.16 77.40 $\pm$0.20 79.35 $\pm$0.04 79.32 $\pm$0.07 83.13 $\pm$0.24 +3.81 $\blacktriangle$
-Snow 71.78 78.29 $\pm$0.20 70.16 $\pm$0.10 77.29 $\pm$0.16 79.44 $\pm$0.09 79.79 $\pm$0.06 83.71 $\pm$0.17 +3.92 $\blacktriangle$
-Frost 72.86 79.84 $\pm$0.09 72.39 $\pm$0.22 79.20 $\pm$0.08 80.13 $\pm$0.10 80.54 $\pm$0.12 83.40 $\pm$0.20 +2.86 $\blacktriangle$
-Fog 67.04 77.39 $\pm$0.01 64.31 $\pm$0.28 75.74 $\pm$0.14 77.68 $\pm$0.07 78.53 $\pm$0.22 82.56 $\pm$0.08 +4.03 $\blacktriangle$
-Brightness 81.87 87.78 $\pm$0.03 81.30 $\pm$0.18 86.59 $\pm$0.16 87.10 $\pm$0.10 87.11 $\pm$0.11 89.90 $\pm$0.06 +2.79 $\blacktriangle$
-Contrast 64.37 79.47 $\pm$0.11 62.26 $\pm$0.31 77.82 $\pm$0.14 80.04 $\pm$0.24 81.20 $\pm$0.22 84.86 $\pm$0.11 +3.66 $\blacktriangle$
-Elastic Transform 60.83 70.00 $\pm$0.25 56.43 $\pm$0.27 70.20 $\pm$0.01 71.76 $\pm$0.10 72.66 $\pm$0.15 76.08 $\pm$0.30 +3.42 $\blacktriangle$
-Pixelate 50.53 63.74 $\pm$0.18 42.80 $\pm$0.40 66.52 $\pm$0.13 69.28 $\pm$0.09 71.11 $\pm$0.13 76.25 $\pm$0.05 +5.14 $\blacktriangle$
-JPEG Compression 55.48 62.64 $\pm$0.14 53.67 $\pm$0.25 63.51 $\pm$0.14 66.49 $\pm$0.14 67.36 $\pm$0.28 70.03 $\pm$0.37 +2.67 $\blacktriangle$
-Average 59.22 67.56 56.80 71.17 72.83 73.82 77.06 +3.24 $\blacktriangle$
-CIFAR-100 61.68 69.74 $\pm$0.16 63.78 $\pm$0.28 69.79 $\pm$0.04 70.38 $\pm$0.14 70.74 $\pm$0.20 71.68 $\pm$0.09 +0.94 $\blacktriangle$
-CIFAR-100C Gaussian Noise 14.80 14.38 $\pm$0.14 14.03 $\pm$0.10 25.32 $\pm$0.14 31.28 $\pm$0.03 32.07 $\pm$0.23 33.43 $\pm$0.13 +1.36 $\blacktriangle$
-Shot Noise 16.03 17.34 $\pm$0.27 15.25 $\pm$0.17 27.90 $\pm$0.05 33.44 $\pm$0.11 34.36 $\pm$0.11 35.60 $\pm$0.19 +1.24 $\blacktriangle$
-Impulse Noise 13.85 10.03 $\pm$0.13 13.01 $\pm$0.13 25.62 $\pm$0.09 29.40 $\pm$0.11 30.33 $\pm$0.03 30.94 $\pm$0.41 +0.61 $\blacktriangle$
-Defocus Blur 36.74 49.05 $\pm$0.07 37.60 $\pm$0.17 49.88 $\pm$0.23 52.32 $\pm$0.28 52.99 $\pm$0.16 53.87 $\pm$0.22 +0.88 $\blacktriangle$
-Glass Blur 14.19 3.71 $\pm$0.07 16.41 $\pm$0.02 27.89 $\pm$0.03 31.20 $\pm$0.12 32.15 $\pm$0.30 35.26 $\pm$0.22 +3.11 $\blacktriangle$
-Motion Blur 36.14 46.62 $\pm$0.27 37.52 $\pm$0.23 47.93 $\pm$0.14 49.72 $\pm$0.15 50.53 $\pm$0.12 52.77 $\pm$0.07 +2.24 $\blacktriangle$
-Zoom Blur 40.24 51.84 $\pm$0.15 42.99 $\pm$0.11 52.70 $\pm$0.06 54.72 $\pm$0.04 55.30 $\pm$0.22 56.71 $\pm$0.17 +1.41 $\blacktriangle$
-Snow 38.95 46.71 $\pm$0.21 42.35 $\pm$0.13 49.72 $\pm$0.01 51.79 $\pm$0.04 52.77 $\pm$0.15 54.30 $\pm$0.19 +1.53 $\blacktriangle$
-Frost 40.56 44.90 $\pm$0.27 43.31 $\pm$0.14 49.63 $\pm$0.12 53.04 $\pm$0.08 53.79 $\pm$0.31 54.92 $\pm$0.07 +1.13 $\blacktriangle$
-Fog 38.00 47.31 $\pm$0.04 38.81 $\pm$0.17 48.77 $\pm$0.04 50.78 $\pm$0.24 51.49 $\pm$0.21 53.57 $\pm$0.02 +2.08 $\blacktriangle$
-Brightness 48.18 60.58 $\pm$0.18 50.23 $\pm$0.11 61.27 $\pm$0.08 62.65 $\pm$0.25 63.57 $\pm$0.21 64.43 $\pm$0.31 +0.86 $\blacktriangle$
-Contrast 29.53 45.90 $\pm$0.11 28.09 $\pm$0.09 48.55 $\pm$0.24 51.34 $\pm$0.10 52.76 $\pm$0.27 55.01 $\pm$0.29 +2.25 $\blacktriangle$
-Elastic Transform 26.33 33.09 $\pm$0.08 28.12 $\pm$0.15 37.45 $\pm$0.08 39.97 $\pm$0.06 40.90 $\pm$0.43 43.79 $\pm$0.27 +2.89 $\blacktriangle$
-Pixelate 21.98 26.47 $\pm$0.09 20.43 $\pm$0.14 33.88 $\pm$0.14 39.59 $\pm$0.09 40.97 $\pm$0.16 44.51 $\pm$0.32 +3.54 $\blacktriangle$
-JPEG Compression 25.91 29.89 $\pm$0.07 28.82 $\pm$0.09 36.07 $\pm$0.32 38.99 $\pm$0.16 39.59 $\pm$0.08 40.83 $\pm$0.14 +1.24 $\blacktriangle$
-Average 29.43 35.19 30.46 41.51 44.68 45.57 47.33 +1.76 $\blacktriangle$
+Algorithm 1 Semantic Anchor Transport for one test batch.
+1:Input: Test batch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$, set of $M$ semantic anchor matrices $\{\mathbf{T}_{m}\}_{m=1}^{M}$, visual encoder $\boldsymbol{\theta}$.
+2:// $\mathbf{T}_{m}$ (and thus $\mathbf{T}$) are pre-computed offline from a systematically generated set of text templates.
 […]
-Link to unsupervised representation learning.
-The proposed distilling strategy takes inspiration from unsupervised representation learning [7, 28, 5], where comparing random crops of the same image has the potential to yield stronger image representations.
-In particular, these strategies leverage these novel “views” of the same image as an image augmentation strategy, whose integration improves the semantic quality of the learned image representations.
-Nevertheless, i) this strategy substantially increases the computational and memory burden [7, 28], ii) specific composition of data augmentation operations is crucial for learning good representations [7], and iii) low-quality crops can alter the model performance [5].
-In contrast, exploiting multiple text embeddings to guide the model adaptation process serves as a similar strategy, while also alleviating these shortcomings.
-More concretely, the text embeddings are low dimensional vectors, which can be computed off-line as they are image agnostic, considerably reducing computation requirements.
-Furthermore, in the absence of supervision, we argue that diverse text descriptions associated with a given class convey more informative representations than individual crops or “views” of an image to learn more generalizable features.
-While images provide valuable visual information, their parts may lack the contextual richness necessary for effective representation learning.
-This distinction suggests that leveraging a variety of textual prompts can indeed serve as a more promising alternative to contrastive learning based on multiple image crops while also being more efficient.
+4:// — Adaptation Phase —
+5:for each template $m$ in a random permutation of $\{1,\dots,M\}$ do
+6:  // Step 1: Align - Compute soft assignments that maximize cross-modal alignment for the m-th “view”
+7:  $\mathbf{Z}\leftarrow[\boldsymbol{\theta}(\mathbf{x}_{1}),\dots,\boldsymbol{\theta}(\mathbf{x}_{B_{T}})]$
+8:  $\mathbf{Q}^{*}_{m}\leftarrow\text{SolveOT}(\mathbf{Z},\mathbf{T}_{m})\quad$ Eq. 9
+9:  // Step 2: Adapt - Distill knowledge into a general representation
+10:  $\mathbf{P}\leftarrow\text{Predict}(\mathbf{Z},\mathbf{T})\quad$ Eq. 1
+11:  $\mathcal{L}\leftarrow\text{CrossEntropy}(\mathbf{P},\mathbf{Q}^{*}_{m})$
+12:  Update LayerNorm parameters of $\boldsymbol{\theta}$ using $\nabla_{\boldsymbol{\theta}}\mathcal{L}$.
+13:end for
+14:// — Inference Phase —
+[…]
+[…]
+17:Return $\mathbf{P}_{\text{final}}$
 5 Experiments
@@
 Datasets.
-All methods are benchmarked on standard TTA datasets known for their complexity and task diversity.
-These datasets include corrupted modifications from well-known datasets proposed in [15], such as CIFAR-10C, CIFAR-100C, or Tiny-ImageNet-C.
-Also, we employ its non-corrupted versions, i.e., CIFAR-10 and CIFAR-100 [21], CIFAR-10.1 [34], and Tiny-ImageNet [41].
-Additionaly, we include Domainbed datasets (PACS [24], OfficeHome [38], VisDA-C [32]).
-More details can be found in Appendix B.
-Baselines.
-We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [39], TPT [36], CLIPArTT [14], and WATT [31] (both versions WATT-S and WATT-P).
-Architectures.
-We employed the Vision Transformer architecture as the backbone for our experiments.
-The choice of ViT is driven by its superior performance on image classification tasks due to its attention mechanisms, which effectively capture long-range dependencies and spatial relationships within images.
-If the opposite is not stated, experiments are carried out using ViT-B/32, although we also provide further evaluations with larger ViT-backbones, e.g., ViT-B/16.
-Implementation details.
-Following the existing literature [31, 14], we used a batch size of 128 in our experiments and a learning rate of $10^{-4}$.
-The only additional hyper-parameters specific to our formulation are the entropic constraint weight, $\epsilon$ in Eq.
-7, and the number of repetitions in the Sinkhorn algorithm.
-These are fixed to the same values across all experiments.
-$\epsilon$ is set to $\epsilon=0.7$, whereas the latter is fixed to $T=3$, which was enough for convergence, as in previous works [3].
-Finally, we ran the adaptation experiments for each corruption/dataset three times under different random seeds and averaged the obtained accuracy accordingly.
+We evaluate on standard TTA benchmarks [maharana2024texttt, osowiechi2024watt, hakim2024clipartt], covering two primary types of domain shift: i) Visual corruptions: CIFAR-10C, CIFAR-100C, ImageNet-C, and Tiny-ImageNet-C [datasets_corruptions]; and ii) Domain generalization: PACS [pacs], VLCS [vlcs], OfficeHome [officehome], and VisDA-C [visda].
+We also use the original clean benchmarks (e.g., CIFAR-10 [cifar], CIFAR-10.1 [cifar101], and Tiny-ImageNet [TinyImagenet]) for ablation.
+Full dataset details are presented in Appendix A.
+Baselines: We resort to relevant CLIP-based TTA parameter-efficient adaptation approaches, i.e., modify a set of learnable parameters, including TENT [wangtent], SAR [niutowards], VTE [dobler2024lost], TPT [shu2022test], CLIPArTT [hakim2024clipartt], WATT [osowiechi2024watt], and BATCLIP [maharana2024texttt], where [hakim2024clipartt, osowiechi2024watt, maharana2024texttt] represent the state-of-the-art.
+Architectures: Following prior work, we use Vision Transformer (ViT) backbones from CLIP [radford2021learning].
+Our main experiments use ViT-B/32 as in [hakim2024clipartt, osowiechi2024watt], and we demonstrate generalization with additional evaluations on larger ViT-B/16 and L/14 backbones, particularly CLIP’s and SigLIP’s [zhai2023sigmoidlosslanguageimage].
+Implementation details: Following prior work [hakim2024clipartt, osowiechi2024watt], we use a batch size of 128 for all experiments (see Appendix B.4 for a sensitivity analysis on batch size).
+All reported results are the average of three runs using different random seeds.
+We follow the officially recommended hyperparameters for all baseline methods.
+Specifically, TENT [wangtent] and CLIPArTT [hakim2024clipartt] are run for 10 adaptation steps per batch.
+For WATT [osowiechi2024watt], we use their suggested $L=2$, $M=5$, and 8 templates, totaling 80 iterations per batch.
+For BATCLIP [maharana2024texttt], we adapt both the text and visual encoders.
+For all other baselines, we use the learning rates specified in their respective papers for each dataset.
+We set the learning rate for our method to $10^{-4}$ for all datasets, except for ImageNet-C, where we use $5\times 10^{-5}$ as in [maharana2024texttt].
+Our formulation introduces only two new hyperparameters: the entropic constraint weight, $\epsilon$ from Eq.
+7, and the number of Sinkhorn iterations, $T$.
+These are fixed across all datasets and experiments.
+We set $\epsilon=0.7$ (see Appendix B.3) and $T=3$.
+We found $T=3$ to be sufficient for convergence, as in prior work [caron2018deep].
 5.2 Main results
-We start our empirical validation by comparing the proposed approach’s performance to existing state-of-the-art CLIP-based TTA approaches across the different datasets.
-Evaluation in the presence of natural or no domain shift.
-Table 1 reports the performance of the different methods for CIFAR-10, CIFAR-10.1, and CIFAR-100, which introduce a natural or no domain shift.
-The proposed method consistently outperforms existing approaches, even when the domain shift is small, highlighting the usefulness of resorting to our approach over CLIP zero-shot predictions.
-Impact of the performance under common corruptions.
-Furthermore, in Table 1, we report a detailed analysis of the results obtained on CIFAR-10C and CIFAR-100C, which show the superiority of CLIP-OT in adapting CLIP in the presence of common corruptions.
-In particular, compared to the baseline CLIP, our model brings performance gains of nearly $18\%$ in CIFAR-10C and CIFAR-100C, without additional supervision.
-These performance gains are similar when compared to other popular TTA methods, e.g., TENT, or TPT, with differences ranging from $10\%$ to $21\%$.
-While this gap is reduced compared to very recent approaches, i.e., both versions of WATT and CLIPArTT, per-corruption classification accuracy differences are still significant, with CLIP-OT outperforming these two approaches by up to $5\%$ compared to WATT, i.e., $5.14\%$ in pixelate (CIFAR-10C) or $3.11\%$ in Glass Blur (CIFAR-100C), and nearly $11\%$ compared to CLIPArTT, e.g., $10.62\%$ in pixelate and $8.11\%$ in Gaussian Noise.
-As average, CLIP-OT improves the performance of very recent state-of-the-art approaches by a substantial difference of $3.24\%$ (CIFAR-10C), and $1.76\%$ in the more challenging scenario of CIFAR-100C, which contains $\times$10 number of classes.
-What if the number of classes further increases?
-In the following, we assess the performance of all methods on Tiny-ImageNet and Tiny-ImageNet-C benchmarks, which bring additional complexity due to the increased amount of categories present.
-From these results (Table 2), we can observe that the findings reported in previous sections hold, i.e., CLIP-OT significantly outperforming very recent state-of-the-art TTA approaches.
-More concretely, compared to WATT-S, the average improvement over all the corruptions is above $7.0\%$, with our method showing performance gains superior to $10\%$ in multiple corruptions, e.g., Shot Noise or Pixelate, and above $4.5\%$ in all remaining corruptions.
-Given the complexity and additional challenges that Tiny-ImageNet-C brings compared to CIFAR versions, as well as the non-negligible improvement provided by our method, we believe that these results showcase the clear superiority of CLIP-OT in more challenging scenarios.
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
-Tiny-ImageNet 58.29 57.72 58.90 59.85 61.35 63.69$\blacktriangle$
-Gaussian Noise 7.08 8.01 9.29 14.44 13.02 21.40 $\blacktriangle$
-Shot Noise 9.41 10.04 11.70 17.44 15.94 24.90 $\blacktriangle$
-Impulse Noise 3.44 4.18 4.85 10.37 6.90 17.34 $\blacktriangle$
-Defocus Blur 21.71 24.53 27.56 31.46 29.91 35.39 $\blacktriangle$
-Glass Blur 9.12 10.09 11.03 15.84 14.01 21.16 $\blacktriangle$
-Motion Blur 34.52 36.94 38.97 41.34 41.26 46.26 $\blacktriangle$
-Zoom Blur 27.44 29.48 34.29 35.06 33.96 40.93 $\blacktriangle$
-Snow 32.51 32.20 34.45 36.86 37.76 42.32 $\blacktriangle$
-Frost 36.33 35.72 37.13 38.20 39.65 44.97 $\blacktriangle$
-Fog 25.94 27.46 28.89 33.44 32.13 38.60 $\blacktriangle$
-Brightness 40.15 39.79 43.31 46.43 46.93 53.10 $\blacktriangle$
-Contrast 1.81 2.24 3.15 6.24 3.53 11.88 $\blacktriangle$
-Elastic Transf. 30.40 31.92 33.88 33.89 35.01 40.73 $\blacktriangle$
-Pixelate 22.78 24.79 27.70 34.85 31.55 41.84 $\blacktriangle$
-JPEG Compr. 29.59 30.93 33.60 37.32 36.46 42.86 $\blacktriangle$
-Mean 22.14 23.22 25.32 28.88 27.87 34.91 $\blacktriangle$
-Table 2: Performance comparison on Tiny-ImageNet and Tiny-ImageNet-C corruption benchmarks using ViT-B/32 backbone.
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
-PACS 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
-OfficeHome 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
-VisDA-C (3D) 84.43 84.86 79.35 85.09 85.42 85.36 90.73 +5.37 $\blacktriangle$
-VisDA-C (YT) 84.45 84.68 83.57 84.40 84.57 84.69 85.44 +0.75 $\blacktriangle$
-Table 3: Performance comparison across PACS, OfficeHome, and VisDA-C datasets using ViT-B/32 backbone.
-$\Delta$ represents the difference between CLIP-OT and the State-of-the-art WATT-S [31].
-Extended numerical values are presented in Appendix C.2, Table 9.
-Performance under simulated and video shifts.
-Table 3 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of CLIP-OT compared to other methods.
-More concretely, the average difference compared to SoTA (e.g., WATT-S) is $3.03\%$, whereas per-dataset differences can go up to $11\%$, e.g., $11.38\%$ compared to TPT in 3D domain.
-Texture and style shifts.
-We now evaluate CLIP-OT on different distributional drifts, which may bring additional complexity.
-The results from these experiments (Table 3) show that the trend of superior performance is observed in almost all datasets (i.e., PACS and OfficeHome).
-5.3 Ablation studies
-Figure 2: Ablation on each component across multiple datasets.
-Different results come from adding each element of the proposed formulation.
-In green, performance differences compared to CLIP.
-The numerical values are provided in Appendix C.
+Performance under common visual corruptions.
+We present a unified evaluation across four standard benchmarks in Table 1, which shows the superiority of SAT in adapting CLIP in the presence of common corruptions against a comprehensive suite of recent TTA methods.
+Compared to vanilla CLIP, our model brings performance gains of $17.8\%$ (CIFAR-10C) and $17.9\%$ (CIFAR-100C) without requiring additional supervision.
+These performance gains are similar when compared to other popular TTA methods, e.g., TENT or TPT, with differences ranging from $10\%$ to $21\%$.
+While this gap is reduced compared to recent approaches, such as WATT, CLIPArTT, and BATCLIP, the differences remain significant.
+SAT outperforms the second-best competitor, i.e., WATT, by up to $3.2\%$ in CIFAR-10C (e.g., $5.1\%$ in pixelate) and $1.8\%$ in the more challenging scenario of CIFAR-100C (which contains $\times$10 classes).
+Additionally, it achieves significant gains over recent baselines in specific corruptions, e.g., $19.8\%$ compared to BATCLIP in pixelate (CIFAR-10C) or $16.3\%$ in Glass Blur (CIFAR-100C).
+Also, compared to CLIPArTT, SAT shows gains of nearly $10.6\%$ on pixelate (CIFAR-100C) and $8.1\%$ on Gaussian Noise (CIFAR-100C).
+Table 1: Performance in visual corruptions benchmarks. Results using CLIP ViT-B/32. Best method in bold, second best underlined.
+Method
+Gaussian
+Shot
+Impulse
+Defocus
+Glass
+Motion
+Zoom
+Snow
+Frost
+Fog
+Bright
+Contrast
+Elastic
+Pixelate
+JPEG
+Mean $\uparrow$
+CIFAR-10C
+CLIP [ICLR’21] 35.27 39.67 42.61 69.76 42.40 63.97 69.83 71.78 72.86 67.04 81.87 64.37 60.83 50.53 55.48 59.22
+TENT [ICLR’21] 41.27 47.20 48.58 77.12 52.65 71.25 76.20 78.29 79.84 77.39 87.78 79.47 70.00 63.74 62.64 67.56
+SAR [ICLR’22] 47.58 50.39 47.19 71.65 49.34 70.27 72.63 71.66 72.82 69.48 82.34 70.54 60.98 48.07 58.48 62.89
+VTE [ECCVw’24] 44.40 47.70 42.90 64.90 45.00 66.70 67.00 67.40 64.50 65.30 74.90 53.60 61.20 42.60 50.80 57.26
+TPT [NeurIPS’22] 33.90 38.20 37.66 67.83 38.81 63.39 68.95 70.16 72.39 64.31 81.30 62.26 56.43 42.80 53.67 56.80
+WATT [NeurIPS’24] 63.84 65.28 58.64 78.94 65.12 77.81 79.32 79.79 80.54 78.53 87.11 81.20 72.66 71.11 67.36 73.82
+CLIPArTT [WACV’25] 59.90 62.77 56.02 76.74 61.77 76.01 77.40 77.29 79.20 75.74 86.59 77.82 70.20 66.52 63.51 71.17
+BATCLIP [ICCV’25] 50.89 56.01 54.35 76.17 56.11 74.71 76.01 77.74 79.33 75.87 86.46 78.65 68.76 56.41 61.79 68.62
+SAT (Ours) 64.85 67.34 62.27 82.09 68.07 81.30 83.13 83.71 83.40 82.56 89.90 84.86 76.08 76.25 70.03 77.06
+CIFAR-100C
+CLIP [ICLR’21] 14.80 16.03 13.85 36.74 14.19 36.14 40.24 38.95 40.56 38.00 48.18 29.53 26.33 21.98 25.91 29.43
+TENT [ICLR’21] 14.38 17.34 10.03 49.05 3.71 46.62 51.84 46.71 44.90 47.31 60.58 45.90 33.09 26.47 29.89 35.19
+SAR [ICLR’22] 22.82 25.10 18.68 44.51 21.78 43.04 47.04 46.75 47.34 44.62 57.00 42.17 31.51 25.09 30.83 36.55
+VTE [ECCVw’24] 10.00 10.30 13.30 36.10 20.40 37.90 39.80 42.20 40.80 36.60 45.50 29.20 30.80 17.00 20.70 28.71
+TPT [NeurIPS’22] 14.03 15.25 13.01 37.60 16.41 37.52 42.99 42.35 43.31 38.81 50.23 28.09 28.12 20.43 28.82 30.46
+WATT [NeurIPS’24] 32.07 34.36 30.33 52.99 32.15 50.53 55.30 52.77 53.79 51.49 63.57 52.76 40.90 40.97 39.59 45.57
+CLIPArTT [WACV’25] 25.32 27.90 25.62 49.88 27.89 47.93 52.70 49.72 49.63 48.77 61.27 48.55 37.45 33.88 36.07 41.51
+BATCLIP [ICCV’25] 17.25 19.76 18.98 42.20 19.00 40.81 46.59 41.34 40.14 41.56 53.85 34.07 31.38 25.51 28.96 33.43
+SAT (Ours) 33.43 35.60 30.94 53.87 35.26 52.77 56.71 54.30 54.92 53.57 64.43 55.01 43.79 44.51 40.83 47.33
+Tiny-ImageNet-C
+CLIP [ICLR’21] 7.08 9.41 3.44 21.71 9.12 34.52 27.44 32.51 36.33 25.94 40.15 1.81 30.40 22.78 29.59 22.15
+TENT [ICLR’21] 8.01 10.04 4.18 24.53 10.09 36.94 29.48 32.20 35.72 27.46 39.79 2.24 31.92 24.79 30.93 23.22
+SAR [ICLR’22] 9.09 10.94 3.65 5.50 1.68 14.02 12.08 20.72 24.62 8.37 32.35 0.71 15.32 12.39 25.35 13.12
+VTE [ECCVw’24] 18.63 20.34 4.71 9.62 2.21 30.37 21.68 38.84 40.27 17.41 41.22 0.63 31.64 25.33 37.79 22.71
+TPT [NeurIPS’22] 9.29 11.70 4.85 27.56 11.03 38.97 34.29 34.45 37.13 28.89 43.31 3.15 33.88 27.70 33.60 25.32
+WATT [NeurIPS’24] 13.02 15.94 6.90 29.91 14.01 41.26 33.96 37.76 39.65 32.13 46.93 3.53 35.01 31.55 36.46 27.87
+CLIPArTT [WACV’25] 14.44 17.44 10.37 31.46 15.84 41.34 35.06 36.86 38.20 33.44 46.43 6.24 33.89 34.85 37.32 28.88
+BATCLIP [ICCV’25] 11.96 15.48 10.05 31.89 14.76 43.31 39.07 39.02 39.05 31.91 49.06 5.65 32.79 36.63 39.12 29.32
+SAT (Ours) 21.40 24.90 17.34 35.39 21.16 46.26 40.93 42.32 44.97 38.60 53.10 11.88 40.73 41.84 42.86 34.91
+ImageNet-C
+CLIP [ICLR’21] 11.30 11.58 12.28 20.88 8.92 19.78 17.62 19.92 23.48 25.90 47.34 15.48 17.02 28.00 27.60 20.47
+TENT [ICLR’21] 8.00 7.20 9.20 23.04 10.84 22.86 19.04 21.24 23.86 26.54 48.54 18.32 17.90 30.32 29.66 21.10
+SAR [ICLR’22] 13.07 15.69 13.92 22.74 14.53 23.41 19.49 22.65 24.89 29.47 48.39 18.88 19.61 31.68 29.07 23.17
+VTE [ECCVw’24] 7.12 10.24 9.18 27.31 10.27 26.42 27.36 24.28 26.15 31.22 49.37 13.09 14.18 32.44 31.33 22.66
+TPT [NeurIPS’22] 8.94 7.22 7.55 20.47 9.13 21.78 23.92 24.61 21.54 24.98 40.37 15.22 13.18 30.74 24.63 20.01
+WATT [NeurIPS’24] 7.76 7.06 8.94 24.16 12.46 25.00 21.52 21.58 24.16 26.62 49.74 21.14 19.90 32.70 32.16 22.33
+CLIPArTT [WACV’25] 14.74 15.10 15.30 10.82 9.02 13.82 12.30 16.96 22.52 19.90 41.78 0.26 12.84 22.80 31.94 18.15
+BATCLIP [ICCV’25] 14.84 15.10 15.52 24.42 17.18 25.64 23.08 25.06 25.58 31.08 49.66 18.44 22.20 33.42 33.02 24.95
+SAT (Ours) 18.98 24.54 25.04 25.46 21.98 26.34 24.46 27.78 28.74 29.32 41.98 17.90 25.70 26.32 25.68 26.01
+What if the number of classes increases?
+The gap is even more pronounced on the more difficult Tiny-ImageNet-C testbed, where our $34.9\%$ mean accuracy is $5.6\%$ higher than that of the strongest baseline, BATCLIP.
+Finally, on the highly challenging 1000-class ImageNet-C benchmark, where many methods show limited gains, SAT again proves its robustness.
+Specifically, achieves the highest mean accuracy, $26.0\%$, a promising $5.5\%$ gain over the CLIP baseline, outperforming all other TTA baselines.
+Table 2: Performance comparison under domain shifts. Results using CLIP ViT-B/32. Best method in bold, second best underlined.
+Method PACS O-Home VLCS VisDA-3D VisDA-YT Mean
+(texture / style) (simulated / video)
+CLIP 93.65 77.53 80.16 84.43 84.45 84.04
+TENT 93.81 77.68 80.27 84.86 84.68 84.26
+TPT 93.23 77.20 74.57 79.35 83.57 81.58
+CLIPArTT 93.95 77.56 80.06 85.09 84.40 84.21
+WATT 94.80 78.83 81.14 85.36 84.69 84.96
+BATCLIP 94.52 78.90 80.78 81.97 83.60 83.95
+SAT (Ours) 95.94 80.15 78.33 90.73 85.44 86.12
+Performance under additional shifts:  Table 2 (VisDA-C) reports the results of simulated (3D) and video (YT) shifts, which showcase the superiority of SAT.
+More concretely, the differences compared to the second-competitor SoTA method (WATT) in these two datasets are $5.4\%$ and $0.8\%$.
+Regarding other popular baselines, per-dataset differences can go up to $11\%$, e.g., $11.4\%$ compared to TPT in the 3D domain.
+It is worth noting also the significant performance drop of the very recent BATCLIP on this benchmark, which underscores the instability of methods that rely too heavily on their own initial predictions in the face of large domain gaps.
+Furthermore, Table 2 results on texture and style shifts show that the trend of superior performance is observed in other datasets (i.e., PACS and OfficeHome), whereas performance is degraded in only one out of the five datasets, i.e., VLCS.
+All considered, SAT emerges as a powerful and reliable TTA approach, where average improvement gains compared to recent methods are $2.2\%$ (BATCLIP), $1.2\%$ (WATT), and $1.9\%$ (CLIPArTT), showcasing its versatility.
+5.3 In-depth studies
 Impact of each component.
-In Figure 2, we empirically validate the benefit of each component in CLIP-OT design.
-Each configuration is detailed in Appendix C.1.
-First, we can observe that computing the code $Q^{*}_{m}$ from each template $m$ (Eq.
-(9)) and averaging those to obtain the final predictions consistently improves over zero-shot CLIP.
-It is noteworthy to highlight that the transportation polytope $\mathcal{Q}$ enforces additional explicit constraints on $\mathbf{Q}$, which ensure a well-defined probabilistic assignment and “break” the direct dependency between $\mathbf{Q}$ and $\mathbf{P}$.
-This explains the differences between direct CLIP predictions and this version of our approach, which directly leverages $\mathbf{Q}$ as predictions.
-Furthermore, while this performance is far from state-of-the-art approaches, this strategy is training-free, not requiring additional computational complexity due to gradient descent updates.
-Next, Average Template employs the pseudo-codes computed in the previous version to minimize the cross-entropy in Eq.
-(3).
-Furthermore, the prediction for the $i$-th image, $\mathbf{p}_{i}$, is computed with the average template for the class $k$, i.e., $\mathbf{t}_{k}$.
-This shows that leveraging the obtained pseudo-codes to refine the model leads to a performance improvement, which is consistent across all datasets.
-Last, we introduce the novel knowledge distillation strategy proposed in this work, where distinct templates are used to get multiple pseudo-codes and guide the model training, i.e., CLIP-OT.
-Its substantial performance gains are especially notable in the most challenging scenarios, e.g., CIFAR-100C.
-On the impact of multiple templates.
-We now assess how the performance evolves as the number of templates increases (Table 4).
-Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$.
-For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset.
-Results from this experiment show that CLIP-OT performance consistently increases with the number of templates.
-This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
-##\## Templates 1 2 3 4 5 6 7 8
-CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
-CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
-Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
-Table 4: Accuracy (%) of CLIP-OT with varying numbers of templates (see Appendix A) on natural domain shift benchmarks.
+In Fig.
+3, we empirically validate the benefits of each component in SAT (each configuration is detailed in Appendix B.1).
+The ‘Training-Free OT’ (using OT for assignments but not updating the model) already improves over zero-shot CLIP, demonstrating the power of our batch-aware assignment.
+‘Average Template’ (using an averaged text prototype) provides a further boost.
+Finally, our full method, which utilizes ‘Multi-Templates’ knowledge distillation, yields the largest gains, demonstrating that all components are necessary.
+The benefit of leveraging multiple textual semantic views is also supported in Appendix B.5 experiments, where we showcase that performance consistently scales with the number of templates.
+Figure 3: Ablation on each component. Results come from adding each element of SAT. In green, performance differences compared to CLIP (numerical values in Appendix B.2).
 Do observations hold across backbones?
-Table 5 reports the performance across different datasets when ViT-B/16 is employed as the visual encoder of CLIP.
-These results demonstrate that CLIP-OT is model agnostic, as its superiority is consistently maintained across different backbones.
-Dataset CLIP TENT TPT CLIPArTT WATT-S CLIP-OT
-CIFAR-10 89.25 92.75 89.90 92.61 91.97 94.69
-CIFAR-10.1 84.00 88.52 83.75 88.72 88.10 89.25
-CIFAR-10C 60.15 68.00 59.75 73.22 76.22 80.11
-CIFAR-100 64.76 71.73 67.15 71.34 72.85 74.06
-CIFAR-100C 32.01 37.90 33.73 40.08 48.95 51.24
-TinyImageNet 58.12 63.59 63.87 64.73 66.05 66.34
-TinyImageNet-C 20.92 29.78 26.96 32.90 31.66 37.69
-Table 5: Quantitative results with ViT-B/16 as backbone for TTA across different datasets and baselines.
+Tab.
+3 reports the performance across different datasets when larger CLIP pre-trained models are employed.
+These results demonstrate the model-agnostic nature of SAT, as its superiority remains consistent across different backbones.
+Table 3: TTA results using larger CLIP ViT backbones.
+a) ViT-B/16 CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
+CLIP 60.15 32.01 20.92 20.89
+TENT 68.00 37.90 29.78 22.79
+TPT 59.75 33.73 26.96 22.35
+CLIPArTT 73.22 40.08 32.90 23.47
+WATT 76.22 48.95 31.66 24.38
+BATCLIP 73.52 38.85 29.30 27.16
+SAT 80.11 ($\uparrow$ 19.96) 51.24 ($\uparrow$ 19.23) 37.69 ($\uparrow$ 16.77) 28.35 ($\uparrow$ 7.46)
+b) ViT-L/14 CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
+CLIP 76.04 44.59 34.98 32.05
+TENT 79.18 50.14 40.28 33.09
+TPT 75.01 47.58 41.07 30.67
+CLIPArTT 78.06 52.52 42.98 34.13
+WATT 80.06 54.34 43.28 36.30
+BATCLIP 83.79 48.84 35.72 37.07
+SAT (Ours) 86.35 ($\uparrow$ 10.31) 62.21 ($\uparrow$ 17.62) 50.36 ($\uparrow$ 15.38) 38.82 ($\uparrow$ 6.77)
+Generalization across VLMs.
+SAT’s principles are also model-agnostic beyond the CLIP pre-training framework.
+As shown in Table 4, SAT consistently delivers state-of-the-art performance when applied to a modern SigLIP backbone.
+The gains are even more pronounced on this architecture, with a $15.0\%$ improvement on CIFAR-100C and $5.4\%$ on ImageNet-C over zero-shot SigLIP, proving the general applicability of our cross-modal alignment paradigm.
+Table 4: Performance using SigLIP [zhai2023sigmoidlosslanguageimage].
+Method CIFAR-10C CIFAR-100C Tiny-IN-C ImageNet-C
+SigLIP 59.04 34.76 22.00 26.49
+BATCLIP 67.45 39.26 24.72 30.40
+SAT (Ours) 76.20 ($\uparrow$ 17.16) 49.72 ($\uparrow$ 14.96) 26.81 ($\uparrow$ 4.81) 31.92 ($\uparrow$ 5.43)
 Computational analysis.
-Figure 3 depicts the running time required for relevant baselines and the proposed CLIP-OT across several datasets, ordered in terms of increasing complexity, i.e., number of categories.
+Fig.
+4 depicts the running time required for relevant baselines and the proposed SAT across several datasets, ordered in terms of increasing complexity, i.e., number of categories.
 These results expose that different TTA methods, particularly SoTA, substantially differ in total runtimes.
-In particular, recent CLIPArTT and WATT-S constantly increase the required runtime with the number of classes, driven by their iterative nature on the multiple template evaluations and weight averaging during inference.
-In contrast, CLIP-OT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once.
-Moreover, the Sinkhorn algorithm—used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime.
-This combination of speed and consistency establishes CLIP-OT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT-S for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
-Figure 3: Computational efficiency. The runtime of recent CLIP TTA methods is assessed, using a standard input batch of size 128. Runtime differences compared to our CLIP-OT highlighted in red.
+In particular, recent CLIPArTT and WATT constantly increase the required runtime with the number of classes, driven by their iterative nature, which involves multiple template evaluations and weight averaging during inference.
+BATCLIP is also highly efficient, yet its runtime still shows a slight increase as the number of classes grows.
+In contrast, SAT avoids this overhead by distilling this information during the adaptation stage and computing multiple text embeddings off-line only once.
+It is worth noting that the Sinkhorn algorithm used for generating pseudo-codes is highly efficient, as its iterative updates converge faster on the low-dimensional similarity matrices and benefit from parallelizable computations, accounting for only nearly $1\%$ of the total runtime.
+This combination of speed and consistency establishes SAT as a scalable and effective method, providing runtimes from one to two orders of magnitude lower than CLIPArTT and WATT for the most complex dataset (i.e., Tiny-ImageNet), while significantly outperforming them from a discriminative standpoint.
+Figure 4: Inference Runtime Comparison. Runtime in seconds for ViT-B/32 TTA methods on an NVIDIA RTX A6000 using a test-time batch of size=128 images.
 6 Conclusions
-We have presented a novel approach for TTA of VLMs.
-In particular, CLIP-OT generates pseudo-labels for test samples using multiple, generic class templates, optimizing a label assignment problem solved with Optimal Transport.
-This strategy yields state-of-the-art results compared to recent works in several benchmarks, yet being a significantly more efficient solution in terms of computational complexity.
+We have presented Semantic Anchor Transport (SAT), a novel approach for the test-time adaptation of vision-language models.
+SAT reformulates TTA as a principled cross-modal alignment problem.
+It generates robust, batch-aware pseudo-labels by aligning visual embeddings to fixed text-based semantic anchors using Optimal Transport.
+This global assignment strategy fundamentally mitigates the error accumulation demonstrated by other TTA methods.
+Furthermore, SAT employs a sophisticated multi-template distillation strategy to harness diverse textual clues, enhancing robustness without incurring significant computational overhead.
+Extensive experiments demonstrate that SAT achieves state-of-the-art performance on TTA across multiple visual domain shift benchmarks and multi-modal backbones.
+[…]
 Supplementary Material
-Appendix A Templates Used
-In the experimental setup, several predefined text templates from CLIP [11] were used to evaluate the proposed model’s adaptability and performance.
+A Additional datasets details.
+Datasets.
+CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems.
+These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set.
+Similarly, ImageNet-C provides these corruptions for the 1000-class ImageNet validation set.
+These datasets are critical for evaluating how well models can generalize to real-world variations and noise.
+Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes.
+Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions.
+This dataset is instrumental for testing robustness against a variety of distortions.
+To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [hakim2024clipartt].
+Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
+CLIP’s text templates.
+In the experimental setup, several predefined text templates from CLIP [clipad] were used to evaluate the proposed model’s adaptability and performance.
 These are designed to generalize across different contexts and image types, enabling a robust assessment of the model’s capabilities in handling diverse visual and textual representations.
@@
 Each template includes a placeholder , which is dynamically replaced with the class name during the generation of text prompts.
-Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [31].
+Note that, for the sake of fairness with prior literature, the templates are the same as the ones employed in WATT [osowiechi2024watt].
+Table 5: The different templates used during the experiments.
 Template
-1: a photo of a {class $k$}
-2: itap of a {class $k$}
-3: a bad photo of the {class $k$}
-4: a origami {class $k$}
-5: a photo of the large {class $k$}
-6: a {class $k$} in a video game
-7: art of the {class $k$}
-8: a photo of the small {class $k$}
-Table 6: The different templates used during the experiments.
-Appendix B Additional details on the datasets
-CIFAR-10.1 introduces a natural shift from CIFAR-10, whereas CIFAR-10-C and CIFAR-100-C are augmented with 15 different corruptions across 5 severity levels (each containing 10,000 images), leading to 75 corruption scenarios commonly employed in domain shift problems.
-These datasets are critical for evaluating how well models can generalize to real-world variations and noise that are not present in the training set.
-Tiny-ImageNet is a downsized version of the original ImageNet dataset, providing a more accessible challenge with 200 classes.
-It is a useful benchmark for lightweight models and educational purposes, making it easier to implement and experiment with domain adaptation techniques.
-Tiny-ImageNet-C further extends Tiny-ImageNet by incorporating various common corruptions, similar to CIFAR-10-C and CIFAR-100-C.
-This dataset is instrumental for testing the robustness of models against a variety of distortions in a more manageable setting.
-To assess the performance on class imbalanced datasets, we employ several datasets from Domainbed (PACS, OfficeHome, VisDA-C) that are often utilized for benchmarking domain adaptation algorithms, and have been recently adopted in TTA scenarios [14].
-Each dataset presents unique domain shift challenges and diverse visual categories, allowing for a comprehensive evaluation of a model’s ability to generalize across different environments.
-Appendix C Additional details of the ablation studies
-C.1 Configuration: Baselines
-In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and CLIP-OT with Average Template.
-Note that, in short, Training-free OT motivates the use of the pseudo-codes $\mathbf{Q}^{*}$ over the CLIP baseline, whereas CLIP-OT with Average Template includes these pseudo-codes to fine-tune the model.
-Furthermore, we want to highlight that the later does not fully leverage multiple individual templates, which is introduced in our proposed CLIP-OT model.
+1: a photo of a {class kk}
+2: itap of a {class kk}
+3: a bad photo of the {class kk}
+4: a origami {class kk}
+5: a photo of the large {class kk}
+6: a {class kk} in a video game
+7: art of the {class kk}
+8: a photo of the small {class kk}
+B Additional experimental details.
+B.1 Configuration: Baselines
+In this section, we detail the configurations used to assess the impact of each main component of our approach, which we refer to Training-free OT and SAT with Average Template.
+Note that, in short, Training-free OT motivates the use of the soft assignments $\mathbf{Q}^{*}$ over the CLIP baseline, whereas SAT with Average Template includes these assignments to fine-tune the model.
+Furthermore, we want to highlight that the latter does not fully leverage multiple individual templates, which is introduced in our proposed method.
 A more detailed description follows.
-Given a batch of test images, Training-free OT (Algorithm 2) computes pseudo-codes for each text template, which are later averaged to produce a final prediction.
-On the other hand, CLIP-OT with Average Template (Algorithm 3) utilizes the averaged pseudo-codes (obtained by Training-free OT) to refine the visual encoder.
-More concretely, at each batch, the average pseudo-codes $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model.
+Given a batch of test images, Training-free OT (Algorithm 2) computes soft assignments for each text template, which are later averaged to produce a final prediction.
+On the other hand, SAT with Average Template (Algorithm 3) utilizes the averaged assignments (obtained by Training-free OT) to refine the visual encoder.
+More concretely, at each batch, the average assignments $\tilde{\mathbf{Q}}$ supervise the predictions of the test images produced by the model.
 These predictions are obtained by resorting to Eq.
-(1), where the average class embedding $\mathbf{t}_{k}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$.
-Differences between the pseudo-codes and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model.
+(1), where the average class embedding $\mathbf{T}$ is used, together with the visual embeddings of the test images, $\mathbf{Z}$.
+Differences between the assignments and the predictions are minimized via a cross-entropy loss, whose gradients are used to update the layer norm parameters of the model.
 Then, once the model is updated, we can do the final inference for the test images.
-C.2 Extended numerical values
-We further substantiate the findings presented in Figure 2 by providing detailed results in Tables 7 and 8.
+Algorithm 2 Training-free OT.
+1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$).
+// Split $\mathcal{X}$ into $B$ batches of size $B_{T}$.
+// Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
+[…]
+3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$ do
+4:   for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Align - Compute soft assignments $\mathbf{Q}^{*}_{m}$.
+5:     $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
+6:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
+7:     Compute soft assignments $\mathbf{Q}^{*}_{m}$ // Eq. (9).
+8:   end for// Inference (for all images in the batch).
+9:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
+10:end for
+Algorithm 3 SAT with Average Template.
+1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$).
+// Split $\mathcal{X}$ into $B$ batches of size $B_{T}$.
+// Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
+[…]
+3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{T}}$ do // 1 - Adaptation
+4:   for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Align - Compute soft assignments $\mathbf{Q}^{*}_{m}$.
+5:     $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
+6:     $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
+7:     Compute soft assignments $\mathbf{Q}^{*}_{m}$ // Eq. (9).
+8:   end for// Step 2: Compute average assignments and refine the encoder.
+9:   $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average assignments.
+10:   $\mathbf{T}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{T}_{m}$ // average class text embeddings.
+11:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict with $\mathbf{T}$, Eq. (1).
+12:   Min. cross-entropy loss between $\mathbf{P}$ and $\tilde{\mathbf{Q}}$ // Eq. (3).
+13:   $\theta_{\text{LN}}^{(t)}\rightarrow\theta_{\text{LN}}^{(t+1)}$ // update layer norm. // 2 - Inference (for all images in the batch).
+14:   $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{T}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
+15:   $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{T}}]$ // predict with $\mathbf{T}$, Eq. (1).
+16:end for
+B.2 Extended numerical values
+We further substantiate the findings presented in Figure 3 by providing detailed results in Tables 6 and 7.
 The tables showcase a comprehensive performance comparison of the impact of the different components of our approach, which empirically motivate our model.
 These results are reported for different corruption benchmarks on the CIFAR-10, CIFAR-10.1, CIFAR-10C, CIFAR-100, and CIFAR-100C datasets, respectively, using a ViT-B/32 backbone.
-Also, Table 9 extends the results obtained on the different domains of PACS and OfficeHome, whose average results were reported in the main paper.
-We can observe that only in one out of the different datasets and domains, CLIP-OT underperforms previous SoTA, with a marginal performance gap, i.e., -0.19%.
-Algorithm 2 Training-free OT.
-1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$).
-// Split $\mathcal{X}$ into $B$ batches of size $B_{s}$.
-// Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
-2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
-3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do
-4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
-5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
-6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
-7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
-8:       end for // Inference (for all images in the batch).
-9:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict: $\mathbf{p}_{i}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{q}_{im}$.
-10:end for
-Algorithm 3 CLIP-OT with Average Template.
-1:input: test dataset $\mathcal{X}=\{\mathbf{x}_{n}\}$ (images), text templates $\mathcal{T}$ (class descriptions), visual and text encoders ($\boldsymbol{\theta}$, $\phi$).
-// Split $\mathcal{X}$ into $B$ batches of size $B_{s}$.
-// Compute $K\times M$ text prototypes ($\boldsymbol{\phi}(\mathcal{T}),\forall m,k)$.
-2:$\mathbf{T}\in\mathbb{R}^{d\times K\times M}=\left[\mathbf{t}_{km}\right]_{k=1,\dots,K,m=1,\dots,M}$
-3:for sampled minibatch $\{\mathbf{x}_{i}\}_{i=1}^{B_{s}}$ do // 1 - Adaptation
-4:       for each template $m$ in a random permutation of $\{1,2,\dots,M\}$ do// Step 1: Generation of pseudo-codes $Q^{*}_{m}$.
-5:             $\mathbf{T}_{m}\in\mathbb{R}^{d\times K}$ // class text embeddings for $m$.
-6:             $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features ($\boldsymbol{\theta}_{\text{LN}}^{(m-1)}$).
-7:             Compute pseudo-codes $\mathbf{Q}^{*}_{m}$ // Eq. (9).
-8:       end for // Step 2: Compute average pseudo-codes and refine the encoder.
-9:       $\tilde{\mathbf{Q}}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{Q}^{*}_{m}$ // average pseudo-codes.
-10:       $\mathbf{t}_{k}=\frac{1}{M}\sum_{m=1}^{M}\mathbf{t}_{km}$ // average class text embeddings over all templates.
-11:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
-12:       Min. cross-entropy loss between $\mathbf{P}$ and $\tilde{\mathbf{Q}}$ // Eq. (3).
-13:       $\theta_{\text{LN}}^{(t)}\rightarrow\theta_{\text{LN}}^{(t+1)}$ // update layer norm. // 2 - Inference (for all images in the batch).
-14:       $\mathbf{Z}=[\mathbf{z}_{1},\dots,\mathbf{z}_{B_{s}}]$ // visual features (with $\boldsymbol{\theta}_{\text{LN}}^{(m)}$).
-15:       $\mathbf{P}=[\mathbf{p}_{1},\dots,\mathbf{p}_{B_{s}}]$ // predict with $\mathbf{t}_{k},\forall k$, Eq. (1).
-16:end for
-Dataset CLIP Training-Free OT Avg. Template CLIP-OT
+Table 6: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C as shown in Figure 3.
+Dataset CLIP Training-Free OT Avg. Template SAT
 CIFAR-10 88.74 89.44 91.40 93.15
@@
 Mean 59.22 64.05 69.62 77.06
-Table 7: Numerical Analysis for CIFAR-10, CIFAR-10.1 and CIFAR-10-C datasets as shown in Figure 2.
-Corruption CLIP Training-Free OT Avg. Template CLIP-OT
+Table 7: Numerical Analysis for CIFAR-100 and CIFAR-100-C datasets as shown in Figure 3.
+Corruption CLIP Training-Free OT Avg. Template SAT
 CIFAR-100 61.68 61.55 68.26 71.68
@@
 Mean 29.43 34.85 38.17 47.33
-Table 8: Numerical Analysis for CIFAR-100 and CIFAR-100-C datasets as shown in Figure 2.
-Dataset Domain CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours $\Delta$
-PACS Art 96.34 96.65 $\pm$0.05 95.52 $\pm$0.20 96.57 $\pm$0.09 96.31 $\pm$0.00 96.39 $\pm$0.00 96.54 $\pm$0.05 +0.15 $\blacktriangle$
-Cartoon 96.08 96.22 $\pm$0.05 94.77 $\pm$0.20 96.00 $\pm$0.02 96.52 $\pm$0.02 96.62 $\pm$0.02 98.34 $\pm$0.13 +1.72 $\blacktriangle$
-Photo 99.34 99.40 $\pm$0.00 99.42 $\pm$0.06 99.28 $\pm$0.00 99.48 $\pm$0.03 99.52 $\pm$0.00 99.86 $\pm$0.03 +0.34 $\blacktriangle$
-Sketch 82.85 82.96 $\pm$0.12 83.22 $\pm$0.14 83.93 $\pm$0.14 86.92 $\pm$0.04 86.65 $\pm$0.12 89.01 $\pm$0.29 +2.36 $\blacktriangle$
-Mean 93.65 93.81 93.23 93.95 94.81 94.80 95.94 +1.14 $\blacktriangle$
-OfficeHome Art 73.75 74.03 $\pm$0.27 75.76 $\pm$0.27 73.84 $\pm$0.20 75.65 $\pm$0.27 75.76 $\pm$0.39 79.39 ±$\pm 0.31$ +3.63 $\blacktriangle$
-Clipart 63.33 63.42 $\pm$0.04 63.08 $\pm$0.31 63.54 $\pm$0.06 66.23 $\pm$0.13 65.77 $\pm$0.11 66.21 $\pm$0.17 +0.44 $\blacktriangle$
-Product 85.32 85.51 $\pm$0.08 84.07 $\pm$0.28 85.23 $\pm$0.16 85.41 $\pm$0.09 85.41 $\pm$0.01 86.81 $\pm$0.03 +1.40 $\blacktriangle$
-[…]
-Mean 77.53 77.68 77.20 77.56 78.88 78.83 80.15 +1.32 $\blacktriangle$
-VisDA-C 3D (trainset) 84.43 84.86 $\pm$0.01 79.35 $\pm$0.04 85.09 $\pm$0.01 85.42 $\pm$0.03 85.36 $\pm$0.01 90.73 $\pm$0.65 +5.37 $\blacktriangle$
-YT (valset) 84.45 84.68 $\pm$0.01 83.57 $\pm$0.04 84.40 $\pm$0.01 84.57 $\pm$0.00 84.69 $\pm$0.01 85.44 $\pm$0.65 +0.75 $\blacktriangle$
-Mean 84.44 84.77 81.46 84.75 85.00 85.03 88.08 +3.05 $\blacktriangle$
-[…]
-Appendix D Comparison to OT-VP
-The very recent OT-VP [48] has presented a solution integrating optimal transport for test-time adaptation.
-Nevertheless, it presents several fundamental differences with our work.
-First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized.
-Thus, optimal transport is used for a different task.
-Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality.
-And third, the adaptation scenario strongly differs from our setting.
-In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome).
-While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 10.
-In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps.
-To expand the extent of our empirical validation, we compare our approach to OT-VP [48] whose results are reported in Table 10.
-To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
-Method PACS OfficeHome Avg.
-First Setting : Single Source
-[…]
-[…]
-Second Setting : Multi Source
-[…]
-[…]
-Our setting: No specific source
-CLIP-OT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
-Table 10: Performance comparison to [48] across PACS and OfficeHome datasets.
-Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
-Appendix E Differences wrt PLOT [6]
-Note that despite the apparent similarities, our work presents fundamental differences w.r.t.
-Prompt Learning with Optimal Transport (PLOT) [6], which prevents from adding the later in the empirical evaluation.
-a) Target task.
-PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available.
-Thus PLOT cannot be compared in the TTA setting.
-b) Adaptation.
-Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category).
-c) Use of OT.
-PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities.
-In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.
-Appendix F Additional Details on Epsilon ($\epsilon$)
+B.3 Additional Details on Epsilon ($\epsilon$)
 In the main paper, we stated that $\epsilon$, the entropic constraint weight in Eq.
@@
 Further, we show the values for CIFAR-10 and CIFAR-100 datasets along with averaged performance over corruptions in the Figure 5.
+Figure 5: Ablation on epsilon values across multiple datasets.
+Table 8: Accuracy on CIFAR-10 and CIFAR-100 under different corruptions and epsilon values.
 Corruption CIFAR-10 CIFAR-100
@@
 Mean 70.46 76.14 76.75 77.06 31.20 46.23 47.41 46.71
-Table 11: Mean Accuracy for CIFAR-10 and CIFAR-100 under Different Corruptions and Epsilon Values
 We can observe that higher $\epsilon$ values (0.7 and 0.9) improved stability and performance due to stronger regularization introduced by higher entropic constraints, which prevents degenerate solutions during Sinkhorn normalization.
@@
 The chosen value ($\epsilon=0.7$) strikes a balance between stability and performance, yielding robust results across both clean and corrupted datasets.
-Figure 4: Ablation on epsilon values across multiple datasets.
-Appendix G Detailed benefits of our multi-template distillation approach.
-The proposed multi-template distillation strategy resembles to the unsupervised approach employed to leverage pseudo-labels in semi-supervised learning, e.g., [2].
-In this learning scenario, unlabeled samples are augmented, whose predictions are averaged to obtain a pseudo-label as $\mathbf{q}_{u}=\frac{1}{M}\sum_{m}f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$, where $\bar{\mathbf{x}}_{u,m}$ is the $m$-th augmented version of a given unlabeled image $\mathbf{x}_{u}$.
-Once the average pseudo-label is obtained, we can minimize the pseudo cross-entropy as $-\sum_{k}\mathbf{q}_{u}\log f_{\theta}(y|\bar{\mathbf{x}}_{u,m})$ [^3] For simplicity, we consider only one unlabeled image..
-In this setting, the predictions from multiple augmented images are guided by the average pseudo-label.
-Once the loss is computed over all the predictions from the augmented images, the gradient is back-propagated and the network parameters updated.
-Limitations of this method: First, this approach forces all augmentations to align with a single static target $\mathbf{q}$, potentially causing the model to converge to a suboptimal solution because it cannot make incremental adjustments based on each individual augmentation.
-Second, it treats all augmentations as being equally aligned with a single target $\mathbf{q}$, which could limit the model’s ability to handle diverse augmentations, especially if some augmentations are very different from others.
-And last, this strategy applies a global update that may force the model too quickly into a solution where all augmentations have the same prediction target $\mathbf{q}$, potentially making it harder for the model to learn robust features.
-To overcome these limitations in the TTA scenario, where no single supervision is available, the proposed method takes an opposite direction to these well-established practices.
-More concretely, our approach minimizes the following pseudo cross-entropy loss at each step $m$: $-\mathbf{q}_{m}\log\mathbf{p}$, where $\mathbf{q}_{m}$ is the pseudo-label at step $m$ (derived from the $m$-th text prompt, which is equivalent to one augmentation), and $\mathbf{p}$ is the mean prediction considering all the $M$ text prompts (similar to averaging over $M$ augmentations).
-At step $m$, the gradient is backpropagated, the network parameters updated.
-Then, we can recompute the pseudo-labels at step $m+1$, as well as the new mean prediction $\mathbf{p}$ (which will be now different due to the parameters update).
-Advantages: First, by updating the model after each step and recomputing $\mathbf{p}$ as the mean of the model’s outputs, our approach encourages a more fine-grained, iterative refinement of predictions, potentially helping the model to converge more smoothly.
-Indeed, the model can adjust its predictions for each template (i.e., augmentation) before being evaluated again in the next step.
-Furthermore, this step-by-step refinement process allows the model to better adapt to the diversity in the augmentations and learn a more generalizable representation.
-Appendix H Discussion on the Two Hyperparameters of WATT and Motivation for CLIP-OT
-WATT [31] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process).
+B.4 Sensitivity to Batch Size
+Our method, SAT, is fundamentally a batch-aware approach.
+We evaluate SAT’s performance on CIFAR-10C while varying the batch size from 16 to 128, comparing it against other methods.
+As shown in Table 9, SAT’s performance is robust and scales well, achieving strong results even with a batch size of 32.
+Notably, under distribution shift on CIFAR-10C, SATexhibits substantially lower performance degradation when reducing the batch size (only –2.5%, from 77.35 to 74.40) compared to much larger drops observed for WATT (–7.5%) and BATCLIP (–6%), highlighting SAT’s superior robustness to batch size constraints in corrupted settings.
+Performance for all methods generally improves with a larger batch, confirming that a sufficiently large batch is important for the stability of TTA.
+Table 9: Accuracy (%) on CIFAR-10 and CIFAR-10C vs. Batch Size. Results are on ViT-B/32.
+Dataset Method 16 32 64 128
+CIFAR-10 WATT 89.14 89.51 90.16 91.05
+BATCLIP 84.31 85.22 88.34 90.35
+SAT (Ours) 87.06 92.34 93.05 93.22
+CIFAR-10C WATT 65.66 68.34 71.21 73.00
+BATCLIP 63.08 61.89 65.51 69.61
+SAT (Ours) 74.40 77.01 77.35 77.06
+B.5 On the impact of multiple templates.
+We now assess how the performance evolves as the number of templates increases (Table 10).
+Concretely, we trained our model multiple times, varying the number of templates $N\in\{1,\dots,8\}$.
+For each $N$, the average accuracy over all template combinations was computed, leading to 256 experiments per dataset.
+Results from this experiment show that SAT performance consistently increases with the number of templates.
+This showcases how leveraging multiple templates individually can extract richer information to adapt CLIP at test time.
+Table 10: Accuracy (%) of SAT with varying numbers of templates (see Table 5) on natural domain shift benchmarks.
+#\# Templates 1 2 3 4 5 6 7 8
+CIFAR-10 91.59 92.16 92.46 92.69 92.88 93.04 93.14 93.14
+CIFAR-100 68.36 69.46 70.08 70.56 70.90 71.15 71.38 71.34
+Tiny-ImageNet 60.62 61.73 62.32 62.69 62.95 63.16 63.20 63.11
+C Detailed Experimental Results
+To verify that our adaptation method does not simply overfit to corruptions at the cost of performance on the original data distribution, we evaluate SAT on clean, non-corrupted benchmarks.
+Table 11 presents the performance of SAT and competing baselines on the standard CIFAR-10, CIFAR-10.1, CIFAR-100, and Tiny-ImageNet test sets.
+C.1 Performance on Clean Datasets
+Table 11 details the performance of all ViT-B/32 methods on the original, uncorrupted test sets.
+These results correspond to the clean dataset performance reported in the ablations (Figure 3) and form the basis for the corruption-free evaluation.
+The results show that SAT does not just preserve performance but consistently improves accuracy over the CLIP baseline and other TTA methods, even without a significant domain shift.
+On CIFAR-10, SAT achieves $93.15\%$ accuracy, surpassing WATT at $91.05\%$.
+This trend is even more pronounced in more complex datasets.
+On CIFAR-100, SAT achieves $71.68\%$, a full 10-point gain over the $61.68\%$ from CLIP and is notably higher than all other baselines.
+Similarly, on Tiny-ImageNet, SAT $63.69\%$ provides a clear improvement over both the baseline ($58.29\%$) and WATT ($61.35\%$).
+Table 11: Performance on Clean Datasets (ViT-B/32). Best method in bold.
+Dataset CLIP TENT TPT CLIPArTT WATT BATCLIP SAT
+CIFAR-10 88.74 91.69 88.06 90.04 91.05 88.89 93.15
+CIFAR-10.1 83.25 87.60 81.80 86.35 86.98 83.70 88.37
+CIFAR-100 61.68 69.74 63.78 69.79 70.74 63.94 71.68
+Tiny-IN 58.29 57.72 58.90 59.85 61.35 58.93 63.69
+C.2 Per-Corruption Results
+To supplement the mean accuracy results presented in the main paper (Table 1), this section provides a granular, per-corruption performance breakdown.
+This analysis demonstrates that SAT’s strong average performance stems from a consistent robustness across a wide variety of distribution shifts.
+We extend this detailed analysis to the larger-scale backbones, ViT-B/16 (Table 12) and ViT-L/14 (Table 13).
+On ViT-B/16 (Table 12): SAT achieves the highest accuracy on all 15 corruption types for both CIFAR-10C and CIFAR-100C.
+For example, on CIFAR-10C Pixelate corruption, SAT ($78.73\%$) significantly outperforms the baseline, WATT ($75.67\%$).
+The gains are similarly strong on ImageNet-C, where SAT (mean $28.35\%$) consistently outperforms all other methods.
+On ViT-L/14 (13): The performance gap widens further, demonstrating SAT’s superior scalability.
+On CIFAR-100C, SAT (mean $69.48\%$) again dominates every single corruption type.
+The improvement is validated on challenging corruptions like Glass Blur, where SAT ($48.19\%$) more than doubles the performance of the zero-shot baseline ($23.46\%$) and is far ahead of the next-best method, WATT ($33.54\%$).
+On the difficult Tiny-ImageNet-C, SAT’s mean accuracy of $50.36\%$ is a major leap from the $34.98\%$ baseline and all other TTA methods.
+This comprehensive, per-corruption breakdown confirms that SAT provides a fundamentally more robust adaptation mechanism that is effective against diverse types of data shifts, and that these benefits are amplified when applied to larger, more capable vision-language models.
+Table 12: Unified corruption robustness results (ViT-B/16, severity = 5).
+Top-1 accuracy (%) on CIFAR-10C, CIFAR-100C, Tiny-ImageNet-C, and ImageNet-C.
+Best method in bold (with gray cell), second best underlined.
+Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixelate JPEG Mean $\uparrow$
+CIFAR-10C
+CLIP 37.75 41.10 51.71 70.07 42.24 65.81 72.50 73.23 76.52 68.35 83.36 61.90 53.16 48.48 56.05 60.15
+TENT 31.04 40.54 58.03 77.57 47.16 76.16 79.64 81.68 83.22 80.78 89.85 79.24 62.54 67.08 65.42 68.00
+TPT 35.35 41.03 54.86 70.29 37.86 67.43 72.91 72.98 75.87 69.13 83.67 62.16 51.26 44.65 56.73 59.75
+WATT 65.57 68.67 70.39 79.90 61.62 79.02 81.10 82.54 83.46 81.88 89.10 83.79 70.93 75.67 69.65 76.22
+CLIPArTT 60.89 65.19 67.55 78.92 57.18 76.59 79.62 81.13 81.24 78.47 88.66 75.15 69.49 71.80 66.42 73.22
+BATCLIP 61.81 65.49 64.32 80.03 53.82 79.67 81.42 82.63 83.60 80.84 87.20 81.17 69.44 63.17 68.21 73.52
+Ours 68.50 71.24 76.97 83.33 68.33 83.13 85.35 86.65 86.37 85.60 92.64 86.92 75.18 78.73 72.65 80.11
+CIFAR-100C
+CLIP 15.88 17.49 21.43 40.10 13.48 39.82 45.45 42.77 45.39 38.98 52.55 33.32 24.39 21.89 27.21 32.01
+TENT 12.28 15.07 13.13 50.35 4.84 49.85 54.76 52.38 51.66 50.74 64.26 48.69 33.56 36.20 30.80 37.90
+TPT 15.43 16.88 22.12 41.08 18.43 40.85 46.77 47.24 48.61 39.92 55.83 33.13 27.36 21.26 30.97 33.73
+WATT 35.95 37.96 44.62 53.80 33.39 52.72 57.51 56.73 56.48 53.83 66.67 55.06 40.37 47.02 42.13 48.95
+CLIPArTT 19.01 20.27 17.66 49.86 18.34 50.00 54.13 52.80 49.56 49.92 63.76 47.86 32.93 39.49 35.56 40.08
+BATCLIP 21.79 28.07 30.81 45.46 21.16 45.04 49.82 49.52 47.81 47.32 59.12 43.72 30.89 30.17 32.06 38.85
+Ours 38.80 40.89 46.39 55.50 37.89 55.10 58.77 57.84 58.09 55.68 67.76 58.29 44.32 49.46 43.87 51.24
+Tiny-ImageNet-C
+CLIP 4.80 6.44 4.36 22.23 7.50 33.46 29.86 31.64 33.26 24.26 40.48 1.72 23.85 22.35 27.65 20.92
+TENT 13.44 17.46 9.09 32.33 12.47 44.17 39.24 38.91 42.04 31.06 49.26 3.06 34.57 37.73 41.92 29.78
+TPT 10.53 12.90 6.57 29.16 10.88 39.91 36.62 38.66 40.95 29.22 47.78 3.58 32.11 29.02 36.51 26.96
+WATT 16.78 19.27 10.42 33.33 16.26 44.21 40.26 42.31 44.83 34.08 52.16 5.43 35.11 37.47 42.93 31.66
+CLIPArTT 19.25 21.80 13.46 33.49 18.33 45.17 40.53 42.71 44.28 34.78 51.61 7.24 35.16 41.11 44.58 32.90
+BATCLIP 11.92 15.36 10.14 31.77 14.80 43.38 39.13 38.92 38.98 31.88 49.11 5.54 32.91 36.72 38.98 29.30
+Ours 24.88 27.99 20.69 37.27 22.72 48.75 45.16 45.74 48.09 41.21 55.75 13.06 41.69 45.83 46.55 37.69
+ImageNet-C
+CLIP 10.70 11.28 10.66 19.60 13.40 18.06 17.54 26.78 27.08 32.86 49.02 14.98 11.94 30.00 29.38 20.89
+TENT 7.98 8.68 8.82 21.64 17.08 21.38 19.78 28.26 28.40 33.86 49.54 18.10 13.18 32.56 32.54 22.79
+TPT 7.15 6.23 7.03 20.11 11.24 21.36 23.19 26.04 31.17 32.01 50.12 21.05 8.13 39.22 33.15 22.35
+WATT 7.68 8.80 9.14 23.70 18.36 24.88 22.54 29.56 29.36 35.52 50.82 21.44 14.60 34.90 34.40 24.38
+CLIPArTT 19.10 19.80 19.10 14.10 18.50 19.40 17.00 26.40 29.00 35.20 47.40 2.20 22.20 28.10 34.65 23.47
+BATCLIP 17.64 19.00 16.50 24.46 19.04 27.20 24.20 31.90 29.74 37.46 52.20 21.56 18.82 32.12 35.56 27.16
+Ours 19.80 21.14 19.20 24.78 21.38 26.82 24.40 32.04 32.36 38.54 49.44 24.42 20.16 35.92 34.84 28.35
+Table 13: Unified corruption robustness results (ViT-L/14, severity = 5).
+Top-1 accuracy (%) on CIFAR-10C, CIFAR-100C, Tiny-ImageNet-C, and ImageNet-C.
+Best method in bold (with gray cell), second best underlined.
+Method Gaussian Shot Impulse Defocus Glass Motion Zoom Snow Frost Fog Bright Contrast Elastic Pixelate JPEG Mean $\uparrow$
+CIFAR-10C
+CLIP 64.64 67.82 78.21 80.73 50.29 80.75 82.75 83.01 84.90 78.44 91.67 84.20 65.45 75.10 72.58 76.04
+TENT 68.87 71.95 80.22 83.10 57.12 82.69 84.91 85.99 87.15 81.30 93.07 87.93 69.96 78.88 75.49 79.18
+TPT 64.44 66.81 76.46 79.01 49.64 78.85 82.32 82.69 84.63 77.56 90.94 82.88 64.81 77.89 71.18 75.01
+WATT 72.73 74.60 80.95 83.15 62.35 82.61 85.44 85.61 86.88 81.79 92.59 87.38 71.25 77.67 75.84 80.06
+CLIPArTT 70.04 71.44 79.42 81.75 58.13 80.76 83.39 84.48 85.21 79.27 91.87 86.19 67.43 77.88 74.46 78.06
+BATCLIP 75.04 79.39 84.19 86.93 71.01 86.80 87.56 89.49 89.44 86.88 94.04 92.00 79.42 84.38 80.32 83.79
+Ours 80.47 82.64 87.80 90.33 77.72 88.93 91.69 91.56 91.43 90.82 95.86 94.53 82.16 87.13 82.60 86.35
+CIFAR-100C
+CLIP 30.55 34.58 44.89 48.88 23.46 50.83 56.02 49.03 53.27 48.51 60.53 50.24 35.07 43.86 39.11 44.59
+TENT 36.93 38.23 49.09 55.23 27.02 56.03 61.19 55.60 58.21 49.26 67.34 59.91 38.49 48.37 43.44 50.14
+TPT 36.10 40.96 49.69 50.43 24.35 51.94 56.96 54.89 58.15 53.37 66.60 53.64 35.72 44.32 44.42 47.58
+WATT 44.13 46.63 56.26 57.66 33.54 57.81 62.74 61.04 62.76 54.70 71.60 63.95 41.27 51.22 49.78 54.34
+CLIPArTT 41.46 44.27 51.44 56.55 30.47 56.98 62.56 58.81 60.38 54.38 69.63 63.39 39.57 50.45 47.45 52.52
+BATCLIP 39.87 43.50 50.11 51.55 30.95 51.03 55.98 53.77 54.53 49.08 65.52 59.07 34.07 47.96 45.61 48.84
+Ours 52.26 54.59 63.02 65.20 48.19 65.09 68.92 66.81 67.71 65.54 76.25 72.34 52.72 59.55 54.98 62.21
+Tiny-ImageNet-C
+CLIP 17.26 21.57 17.32 35.63 13.62 50.11 43.65 48.04 49.36 37.58 55.85 5.78 34.15 45.90 50.13 34.98
+TENT 26.60 28.93 22.22 38.66 10.98 52.53 46.02 49.13 50.46 38.63 58.25 8.66 36.81 50.77 51.88 40.28
+TPT 23.22 27.39 19.79 39.44 15.72 53.91 48.82 51.50 53.89 40.43 61.07 7.48 40.65 52.28 52.38 41.07
+WATT 27.98 31.47 25.38 40.55 17.44 54.46 48.28 52.60 53.77 42.57 62.01 9.21 39.85 51.83 54.19 43.28
+CLIPArTT 29.27 33.01 27.93 39.99 17.09 53.63 47.39 51.97 51.49 42.35 60.69 11.25 38.44 52.22 52.46 42.98
+BATCLIP 23.98 27.34 23.06 35.40 14.42 48.27 42.72 41.82 46.54 36.29 54.78 6.90 36.84 47.59 49.89 35.72
+Ours 38.14 41.02 35.48 46.03 27.93 57.78 52.67 56.85 57.56 50.49 65.28 21.57 47.86 58.23 56.82 50.36
+ImageNet-C
+CLIP 19.26 21.62 17.88 27.20 20.26 31.78 29.22 43.16 38.08 43.38 61.52 29.50 20.76 44.22 32.94 32.05
+TENT 20.48 22.70 15.86 28.00 21.74 33.40 30.44 43.96 38.58 43.92 61.88 32.22 22.22 45.84 35.10 33.09
+TPT 19.00 26.00 17.00 28.00 16.00 30.00 30.00 42.00 31.00 41.00 59.00 32.00 15.00 43.00 31.00 30.67
+WATT 23.32 26.18 6.06 32.26 27.66 38.54 34.96 47.14 41.52 46.46 63.32 38.66 26.92 50.10 41.38 36.30
+CLIPArTT 27.10 29.00 29.20 21.80 21.30 31.00 28.40 39.60 38.50 41.50 64.60 18.40 30.70 43.30 48.00 34.13
+BATCLIP 19.26 22.64 16.50 33.02 29.12 40.50 36.32 47.08 40.86 48.10 63.28 37.28 27.88 50.56 43.64 37.07
+Ours 27.68 29.50 28.74 34.52 30.58 39.08 35.50 47.08 42.26 48.48 59.60 39.84 30.20 48.42 40.88 38.82
+D Comparison to OT-VP
+The very recent OT-VP [zhang2024ot] has presented a solution integrating optimal transport for test-time adaptation.
+Nevertheless, it presents several fundamental differences with our work.
+First, OT-VP involves learning a universal visual prompt for the target domain, for which an optimal transport distance is optimized.
+Thus, optimal transport is used for a different task.
+Second, OT-VP is tailored to only visual models, not being capable of leveraging the valuable information found on the text modality.
+And third, the adaptation scenario strongly differs from our setting.
+In particular, OT-VP first fine-tunes the pre-trained model (pre-trained on ImageNet-1k) to a subdomain (e.g., PACS) with labeled data, to later adapt at test-time to the other subdomains (e.g., OfficeHome).
+While this strategy can be done in a single-source setting, OT-VP also evaluates the performance when multiple domains are used as the source (e.g., supervised adaptation on PACS), and only one left out for testing, referred to as Multi-Source in Table 14.
+In contrast, we follow the protocol of recent literature of CLIP test-time adaptation, where CLIP is directly exposed to the unsupervised test data points, without intermediate adaptation steps.
+To expand the extent of our empirical validation, we compare our approach to OT-VP [zhang2024ot] whose results are reported in Table 14.
+To ensure a rigorous comparison, we use the same visual encoder as the backbone used in OT-VP and other prior models, i.e., ViT-B/16.
+Table 14: Performance comparison to [zhang2024ot] across PACS and OfficeHome datasets.
+Best results highlighted in bold, whereas value indicates the difference wrt OT-VP in both single (first value) and multi (second value) source scenarios.
+Method PACS OfficeHome Avg.
+First Setting : Single Source
+[…]
+[…]
+Second Setting : Multi Source
+[…]
+[…]
+Our setting: No specific source
+SAT(Ours) 96.9+23.4/+9.2 83.4+15.3/+8.3 86.8+16.8/+5.6
+E Discussion on the Two Hyperparameters of WATT and Motivation for SAT
+WATT [osowiechi2024watt] introduces two key hyperparameters: $L$ (number of adaptation iterations for each text embedding) and $M$ (number of repetitions of the weight averaging process).
 While these hyperparameters play an essential role in improving WATT’s adaptation capabilities, they introduce significant computational and scalability challenges, particularly as the number of classes grows.
@@
 This dependence on repeated updates for each class and template becomes especially prohibitive for datasets with a large number of classes, as demonstrated in Figure 4, where runtime increases from CIFAR-10 (10 classes) to Tiny-ImageNet (200 classes).
-These limitations motivate our proposed approach, CLIP-OT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$.
-CLIP-OT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging.
+These limitations motivate our proposed approach, SAT, which simplifies and optimizes the adaptation process, achieving robust performance without the drawbacks associated with $L$ and $M$.
+SAT eliminates these limitations by leveraging the Sinkhorn algorithm for single-pass optimization and precomputing averaged class embeddings, removing the need for iterative updates and weight averaging.
 This not only drastically reduces computational overhead but also ensures scalable, robust adaptation across diverse corruption levels and datasets, offering a practical alternative to WATT’s hyperparameter-dependent framework.
-Appendix I Further results on other backbones
-In this section, we present additional experimental results leveraging different backbone architectures to evaluate the robustness and generalizability of our proposed method.
-Specifically, we analyze the performance of CLIP models with ViT-B/16 and ViT-L/14 backbones across various datasets, including CIFAR, Tiny-ImageNet and their corruptions.
-These experiments validate our approach across diverse settings and highlight its adaptability to challenging domain shifts.
-Table 12 details the performance of CLIP-OT with the ViT-B/16 backbone on CIFAR-10, CIFAR-10.1, and CIFAR-10-C datasets.
-Our proposed method demonstrates consistent improvements across both clean and corrupted data, outperforming baseline methods such as CLIP, TENT, and TPT.
-Similarly for ViT-L/14 backbone on these datasets, we obtain similar gains as illustrated in Table 13.
-Table 14 extends the evaluation to CIFAR-100 and CIFAR-100-C, demonstrating the effectiveness of our method on more complex datasets with larger label spaces.
-Using CLIP with the ViT-B/16 backbone, our approach achieves the highest average accuracy, with substantial improvements under corruptions such as Glass Blur ($+5.30\%$) and Contrast ($+6.23\%$).
-We also evaluate these datasets with backbone ViT/L-14 in Table 15.
-For Tiny-ImageNet and its corrupted version, Tiny-ImageNet-C, Table 16 and Table 17 provides a detailed breakdown.
-On datasets with higher complexity and larger label spaces, such as Tiny-ImageNet and Tiny-ImageNet-C (Table 16), CLIP-OT exhibits substantial gains, consistent with the main paper’s observations.
-Specifically, our method improves over WATT-S by over 6% on average across Tiny-ImageNet-C corruptions, showcasing its scalability to more challenging scenarios.
-I.1 Tiny-Imagenet Experiments
-We report the mean and standard deviation over Tiny-Imagenet experiments (Table 2, main paper) in Table 18.
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
-CIFAR-10 89.25 92.75 $\pm$0.17 89.80 $\pm$0.05 92.61 $\pm$0.05 92.31 $\pm$0.10 91.97 $\pm$0.03 94.69 $\pm$0.18
-CIFAR-10.1 84.00 88.52 $\pm$0.33 83.75 $\pm$0.21 88.72 $\pm$0.33 87.90 $\pm$0.11 88.10 $\pm$0.08 89.25 $\pm$0.08
-Gaussian Noise 37.75 31.04 $\pm$0.38 35.35 $\pm$0.15 60.89 $\pm$0.26 63.10 $\pm$0.12 65.57 $\pm$0.22 68.50 $\pm$0.19
-Shot Noise 41.10 40.54 $\pm$0.41 41.03 $\pm$0.19 65.19 $\pm$0.21 66.31 $\pm$0.10 68.67 $\pm$0.37 71.24 $\pm$0.22
-Impulse Noise 51.71 58.03 $\pm$0.16 54.86 $\pm$0.07 67.55 $\pm$0.09 69.62 $\pm$0.12 70.39 $\pm$0.11 76.97 $\pm$0.17
-Defocus Blur 70.07 77.57 $\pm$0.03 70.29 $\pm$0.02 78.92 $\pm$0.12 79.64 $\pm$0.08 79.90 $\pm$0.07 83.33 $\pm$0.09
-Glass Blur 42.24 47.16 $\pm$0.05 37.86 $\pm$0.17 57.18 $\pm$0.20 58.98 $\pm$0.12 61.62 $\pm$0.21 68.33 $\pm$0.20
-Motion Blur 65.81 76.16 $\pm$0.05 67.43 $\pm$0.11 76.59 $\pm$0.06 78.32 $\pm$0.16 79.02 $\pm$0.07 83.13 $\pm$0.10
-Zoom Blur 72.50 79.64 $\pm$0.12 72.91 $\pm$0.02 79.62 $\pm$0.11 80.67 $\pm$0.07 81.10 $\pm$0.08 85.35 $\pm$0.11
-Snow 73.23 81.68 $\pm$0.03 72.98 $\pm$0.32 81.13 $\pm$0.29 81.99 $\pm$0.10 82.54 $\pm$0.18 86.65 $\pm$0.12
-Frost 76.52 83.22 $\pm$0.05 75.87 $\pm$0.16 81.24 $\pm$0.08 83.41 $\pm$0.16 83.46 $\pm$0.15 86.37 $\pm$0.09
-Fog 68.35 80.78 $\pm$0.15 69.13 $\pm$0.27 78.47 $\pm$0.19 81.36 $\pm$0.12 81.88 $\pm$0.12 85.60 $\pm$0.24
-Brightness 83.36 89.85 $\pm$0.11 83.67 $\pm$0.14 88.66 $\pm$0.15 89.06 $\pm$0.05 89.10 $\pm$0.14 92.64 $\pm$0.11
-Contrast 61.90 79.24 $\pm$0.19 62.16 $\pm$0.06 75.15 $\pm$0.07 81.57 $\pm$0.23 83.79 $\pm$0.12 86.92 $\pm$0.18
-Elastic Transform 53.16 62.54 $\pm$0.08 51.26 $\pm$0.23 69.49 $\pm$0.08 69.14 $\pm$0.09 70.93 $\pm$0.20 75.18 $\pm$0.14
-Pixelate 48.48 67.08 $\pm$0.24 44.65 $\pm$0.21 71.80 $\pm$0.16 73.38 $\pm$0.29 75.67 $\pm$0.32 78.73 $\pm$0.30
-JPEG Compression 56.05 65.42 $\pm$0.05 56.73 $\pm$0.07 66.42 $\pm$0.25 69.02 $\pm$0.10 69.65 $\pm$0.23 72.65 $\pm$0.18
-Mean 60.15 68.00 59.75 73.22 75.04 76.22 80.11
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
-CIFAR-10 95.36 96.13 $\pm$0.06 95.18 $\pm$0.02 95.16 $\pm$0.03 95.91 $\pm$0.10 95.71 $\pm$0.03 97.61 $\pm$0.03
-CIFAR-10.1 91.20 92.22 $\pm$0.25 91.32 $\pm$0.12 91.02 $\pm$0.02 92.97 $\pm$0.13 92.10 $\pm$0.33 93.45 $\pm$0.30
-Gaussian Noise 64.64 68.87 $\pm$0.20 64.44 $\pm$0.11 70.04 $\pm$0.31 72.81 $\pm$0.09 72.73 $\pm$0.03 80.47 $\pm$0.26
-Shot Noise 67.82 71.95 $\pm$0.06 66.81 $\pm$0.19 71.44 $\pm$0.16 74.45 $\pm$0.16 74.60 $\pm$0.03 82.64 $\pm$0.29
-Impulse Noise 78.21 80.22 $\pm$0.19 76.46 $\pm$0.17 79.42 $\pm$0.15 81.36 $\pm$0.09 80.95 $\pm$0.15 87.80 $\pm$0.15
-Defocus Blur 80.73 83.10 $\pm$0.03 79.01 $\pm$0.23 81.75 $\pm$0.19 83.20 $\pm$0.10 83.15 $\pm$0.18 90.33 $\pm$0.12
-Glass Blur 50.29 57.12 $\pm$0.07 49.64 $\pm$0.23 58.13 $\pm$0.23 61.51 $\pm$0.07 62.35 $\pm$0.15 77.72 $\pm$0.31
-Motion Blur 80.75 82.69 $\pm$0.11 78.85 $\pm$0.04 80.76 $\pm$0.12 82.60 $\pm$0.13 82.61 $\pm$0.12 88.93 $\pm$0.17
-Zoom Blur 82.75 84.91 $\pm$0.08 82.32 $\pm$0.13 83.39 $\pm$0.05 85.76 $\pm$0.06 85.44 $\pm$0.13 91.69 $\pm$0.13
-Snow 83.01 85.99 $\pm$0.11 82.69 $\pm$0.10 84.48 $\pm$0.07 84.91 $\pm$0.13 85.61 $\pm$0.15 91.56 $\pm$0.11
-Frost 84.90 87.15 $\pm$0.12 84.63 $\pm$0.08 85.21 $\pm$0.06 87.17 $\pm$0.13 86.88 $\pm$0.04 91.43 $\pm$0.12
-Fog 78.44 81.30 $\pm$0.07 77.56 $\pm$0.17 79.27 $\pm$0.07 81.80 $\pm$0.11 81.79 $\pm$0.09 90.82 $\pm$0.11
-Brightness 91.67 93.07 $\pm$0.04 90.94 $\pm$0.04 91.87 $\pm$0.09 92.78 $\pm$0.05 92.59 $\pm$0.16 95.86 $\pm$0.00
-Contrast 84.20 87.93 $\pm$0.04 82.88 $\pm$0.09 86.19 $\pm$0.06 87.54 $\pm$0.12 87.38 $\pm$0.02 94.53 $\pm$0.08
-Elastic Transform 65.45 69.96 $\pm$0.12 64.81 $\pm$0.14 67.43 $\pm$0.24 71.19 $\pm$0.07 71.25 $\pm$0.09 82.16 $\pm$0.23
-Pixelate 75.10 77.89 $\pm$0.05 72.92 $\pm$0.12 77.11 $\pm$0.10 77.88 $\pm$0.13 77.67 $\pm$0.16 87.13 $\pm$0.16
-JPEG Compression 72.58 75.49 $\pm$0.07 71.18 $\pm$0.19 74.46 $\pm$0.11 75.88 $\pm$0.16 75.84 $\pm$0.18 82.60 $\pm$0.18
-Mean 76.04 79.18 75.01 78.06 80.05 80.06 86.35
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
-CIFAR-100 64.76 71.73 $\pm$0.14 67.15 $\pm$0.23 71.34 $\pm$0.07 72.98 $\pm$0.07 72.85 $\pm$0.15 74.07 $\pm$0.20
-Gaussian Noise 15.88 12.28 $\pm$0.20 15.43 $\pm$0.03 19.01 $\pm$0.24 34.23 $\pm$0.03 35.95 $\pm$0.27 38.80 $\pm$0.51
-Shot Noise 17.49 15.07 $\pm$0.21 16.88 $\pm$0.07 20.27 $\pm$0.21 36.68 $\pm$0.10 37.96 $\pm$0.15 40.89 $\pm$0.10
-Impulse Noise 21.43 13.13 $\pm$0.16 22.12 $\pm$0.15 17.66 $\pm$0.10 43.17 $\pm$0.35 44.62 $\pm$0.20 46.39 $\pm$0.17
-Defocus Blur 40.10 50.35 $\pm$0.03 41.08 $\pm$0.22 49.86 $\pm$0.13 53.13 $\pm$0.12 53.80 $\pm$0.12 55.50 $\pm$0.24
-Glass Blur 13.48 04.84 $\pm$0.14 18.43 $\pm$0.15 18.34 $\pm$0.31 32.53 $\pm$0.03 33.39 $\pm$0.11 37.89 $\pm$0.68
-Motion Blur 39.82 49.85 $\pm$0.37 40.85 $\pm$0.26 50.00 $\pm$0.09 51.63 $\pm$0.06 52.72 $\pm$0.30 55.10 $\pm$0.28
-Zoom Blur 45.45 54.76 $\pm$0.04 46.77 $\pm$0.06 54.13 $\pm$0.08 56.81 $\pm$0.11 57.51 $\pm$0.09 58.77 $\pm$0.33
-Snow 42.77 52.38 $\pm$0.18 47.24 $\pm$0.18 52.80 $\pm$0.27 56.04 $\pm$0.06 56.73 $\pm$0.27 57.84 $\pm$0.13
-Frost 45.39 51.66 $\pm$0.04 48.61 $\pm$0.14 49.56 $\pm$0.08 56.00 $\pm$0.11 56.48 $\pm$0.34 58.09 $\pm$0.04
-Fog 38.98 50.74 $\pm$0.14 39.92 $\pm$0.16 49.92 $\pm$0.11 52.88 $\pm$0.33 53.83 $\pm$0.19 55.68 $\pm$0.38
-Brightness 52.55 64.26 $\pm$0.09 55.83 $\pm$0.10 63.76 $\pm$0.13 65.58 $\pm$0.07 66.67 $\pm$0.19 67.76 $\pm$0.21
-Contrast 33.32 48.69 $\pm$0.08 33.13 $\pm$0.16 47.86 $\pm$0.02 52.90 $\pm$0.06 55.06 $\pm$0.15 58.29 $\pm$0.51
-Elastic Transform 24.39 33.56 $\pm$0.28 27.36 $\pm$0.10 32.93 $\pm$0.23 39.82 $\pm$0.21 40.37 $\pm$0.26 44.32 $\pm$0.18
-Pixelate 21.89 36.20 $\pm$0.28 21.26 $\pm$0.10 39.49 $\pm$0.21 45.10 $\pm$0.06 47.02 $\pm$0.04 49.46 $\pm$0.17
-JPEG Compression 27.21 30.80 $\pm$0.05 30.97 $\pm$0.10 35.56 $\pm$0.23 41.43 $\pm$0.18 42.13 $\pm$0.21 43.87 $\pm$0.23
-Mean 32.01 37.90 33.73 40.08 47.86 48.95 51.24
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-P NeurIPS’24 WATT-S NeurIPS’24 CLIP-OT Ours
-CIFAR-100 73.28 78.03 $\pm$0.08 76.85 $\pm$0.06 79.42 $\pm$0.08 79.33 $\pm$0.05 78.85 $\pm$0.19 81.38 $\pm$0.15
-Gaussian Noise 30.55 36.93 $\pm$0.03 36.10 $\pm$0.11 41.46 $\pm$0.15 43.99 $\pm$0.13 44.13 $\pm$0.11 52.26 $\pm$0.18
-Shot Noise 34.58 40.96 $\pm$0.16 38.23 $\pm$0.13 44.27 $\pm$0.09 46.28 $\pm$0.22 46.63 $\pm$0.17 54.59 $\pm$0.26
-Impulse Noise 44.89 49.09 $\pm$0.14 49.69 $\pm$0.21 51.44 $\pm$0.23 56.15 $\pm$0.04 56.26 $\pm$0.22 63.02 $\pm$0.12
-Defocus Blur 48.88 55.23 $\pm$0.07 50.43 $\pm$0.19 56.55 $\pm$0.22 57.46 $\pm$0.01 57.66 $\pm$0.42 65.20 $\pm$0.17
-Glass Blur 23.46 27.02 $\pm$0.23 24.35 $\pm$0.22 30.47 $\pm$0.14 32.54 $\pm$0.12 33.54 $\pm$0.16 48.19 $\pm$0.17
-Motion Blur 50.83 56.03 $\pm$0.20 51.94 $\pm$0.04 56.98 $\pm$0.18 58.22 $\pm$0.10 57.81 $\pm$0.05 65.09 $\pm$0.17
-Zoom Blur 56.02 61.19 $\pm$0.10 56.96 $\pm$0.16 62.56 $\pm$0.04 62.94 $\pm$0.02 62.74 $\pm$0.06 68.92 $\pm$0.15
-Snow 49.03 55.60 $\pm$0.09 54.89 $\pm$0.11 58.81 $\pm$0.11 60.68 $\pm$0.06 61.04 $\pm$0.27 66.81 $\pm$0.15
-Frost 53.27 58.21 $\pm$0.15 58.15 $\pm$0.33 60.38 $\pm$0.23 62.34 $\pm$0.14 62.76 $\pm$0.22 67.71 $\pm$0.29
-Fog 48.51 53.37 $\pm$0.25 49.26 $\pm$0.13 54.38 $\pm$0.04 54.71 $\pm$0.31 54.70 $\pm$0.13 65.54 $\pm$0.23
-Brightness 60.53 67.34 $\pm$0.17 66.60 $\pm$0.10 69.63 $\pm$0.14 71.52 $\pm$0.07 71.60 $\pm$0.09 76.25 $\pm$0.15
-Contrast 50.24 59.91 $\pm$0.13 53.64 $\pm$0.24 63.39 $\pm$0.13 62.77 $\pm$0.22 63.95 $\pm$0.04 72.34 $\pm$0.27
-Elastic Transform 35.07 38.49 $\pm$0.12 35.72 $\pm$0.09 39.57 $\pm$0.39 41.28 $\pm$0.25 41.27 $\pm$0.15 52.72 $\pm$0.24
-Pixelate 43.86 48.37 $\pm$0.17 44.32 $\pm$0.10 50.45 $\pm$0.16 51.15 $\pm$0.15 51.22 $\pm$0.13 59.55 $\pm$0.06
-JPEG Compression 39.11 44.42 $\pm$0.09 43.44 $\pm$0.11 47.45 $\pm$0.14 49.40 $\pm$0.17 49.78 $\pm$0.18 54.98 $\pm$0.20
-Mean 44.59 50.14 47.58 52.52 54.10 54.34 69.48
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
-Tiny-ImageNet 58.12 63.59 63.87 $\pm$0.09 64.73 $\pm$0.11 66.05 $\pm$0.03 66.34 $\pm$0.07
-Gaussian Noise 4.80 13.44 $\pm$0.21 10.53 $\pm$0.25 19.25 $\pm$0.18 16.78 $\pm$0.09 24.88 $\pm$0.28
-Shot Noise 6.44 17.46 $\pm$0.20 12.90 $\pm$0.45 21.80 $\pm$0.11 19.27 $\pm$0.13 27.99 $\pm$0.32
-Impulse Noise 4.36 9.09 $\pm$0.14 6.57 $\pm$0.20 13.46 $\pm$0.01 10.42 $\pm$0.07 20.69 $\pm$0.18
-Defocus Blur 22.23 32.33 $\pm$0.09 29.16 $\pm$0.32 33.49 $\pm$0.29 $3$$\pm$0.03 37.27 $\pm$0.16
-Glass Blur 7.50 12.47 $\pm$0.14 10.88 $\pm$0.18 18.33 $\pm$0.09 16.26 $\pm$0.08 22.72 $\pm$0.14
-Motion Blur 33.46 44.17 $\pm$0.06 39.91 $\pm$0.43 45.17 $\pm$0.11 44.21 $\pm$0.10 48.75 $\pm$0.37
-Zoom Blur 29.86 39.24 $\pm$0.26 36.62 $\pm$0.11 40.53 $\pm$0.19 40.26 $\pm$0.21 45.16 $\pm$0.08
-Snow 31.64 38.91 $\pm$0.29 38.66 $\pm$0.19 42.71 $\pm$0.11 42.31 $\pm$0.18 45.74 $\pm$0.15
-Frost 33.26 42.04 $\pm$0.15 40.95 $\pm$0.20 44.28 $\pm$0.14 44.83 $\pm$0.13 48.09 $\pm$0.05
-Fog 24.26 31.06 $\pm$0.18 29.22 $\pm$0.24 34.78 $\pm$0.11 34.08 $\pm$0.13 41.21 $\pm$0.19
-Brightness 40.48 49.26 $\pm$0.12 47.78 $\pm$0.44 51.61 $\pm$0.39 52.16 $\pm$0.08 55.75 $\pm$0.24
-Contrast 1.72 3.06 $\pm$0.09 3.58 $\pm$0.07 7.24 $\pm$0.09 5.43 $\pm$0.08 13.06 $\pm$0.18
-Elastic Transform 23.85 34.57 $\pm$0.05 32.11 $\pm$0.10 35.16 $\pm$0.05 35.11 $\pm$0.05 41.69 $\pm$0.12
-Pixelate 22.35 37.73 $\pm$0.18 29.02 $\pm$0.22 41.11 $\pm$0.18 37.47 $\pm$0.21 45.83 $\pm$0.17
-JPEG Compression 27.65 41.92 $\pm$0.45 36.51 $\pm$0.12 44.58 $\pm$0.04 42.93 $\pm$0.08 46.55 $\pm$0.26
-Mean 20.92 29.78 26.96 32.90 31.66 37.69
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
-Tiny-ImageNet 71.34 74.09 $\pm$0.12 74.51 $\pm$0.20 74.57 $\pm$0.15 75.04 $\pm$0.21 75.01 $\pm$0.09
-Gaussian Noise 17.26 26.60 $\pm$0.16 23.22 $\pm$0.12 29.27 $\pm$0.03 27.98 $\pm$0.14 38.14 $\pm$0.09
-Shot Noise 21.57 28.93 $\pm$0.27 27.39 $\pm$0.21 33.01 $\pm$0.18 31.47 $\pm$0.04 41.02 $\pm$0.13
-Impulse Noise 17.32 $2$$\pm$0.29 19.79 $\pm$0.18 27.93 $\pm$0.09 25.38 $\pm$0.15 35.48 $\pm$0.10
-Defocus Blur 35.63 38.66 $\pm$0.19 39.44 $\pm$0.27 39.99 $\pm$0.15 40.55 $\pm$0.04 46.03 $\pm$0.17
-Glass Blur 13.62 10.98 $\pm$0.13 15.72 $\pm$0.18 17.09 $\pm$0.15 17.44 $\pm$0.16 27.93 $\pm$0.18
-Motion Blur 50.11 52.53 $\pm$0.01 53.91 $\pm$0.32 53.63 $\pm$0.28 54.46 $\pm$0.16 57.78 $\pm$0.06
-Zoom Blur 43.65 46.02 $\pm$0.15 48.82 $\pm$0.11 47.39 $\pm$0.04 48.28 $\pm$0.09 52.67 $\pm$0.09
-Snow 48.04 49.13 $\pm$0.33 51.50 $\pm$0.27 51.97 $\pm$0.14 52.60 $\pm$0.04 56.85 $\pm$0.13
-Frost 49.36 50.46 $\pm$0.09 53.89 $\pm$0.22 51.49 $\pm$0.11 53.77 $\pm$0.16 57.56 $\pm$0.35
-Fog 37.58 38.63 $\pm$0.15 40.43 $\pm$0.11 42.35 $\pm$0.04 42.57 $\pm$0.09 50.49 $\pm$0.09
-Brightness 55.85 59.25 $\pm$0.04 61.07 $\pm$0.09 60.69 $\pm$0.12 62.01 $\pm$0.04 65.28 $\pm$0.31
-Contrast 5.78 8.66 $\pm$0.14 7.48 $\pm$0.18 11.25 $\pm$0.08 9.21 $\pm$0.15 21.57 $\pm$0.13
-Elastic Transform 34.15 36.81 $\pm$0.19 40.65 $\pm$0.27 38.44 $\pm$0.22 39.85 $\pm$0.08 47.86 $\pm$0.17
-Pixelate 45.90 50.77 $\pm$0.18 52.28 $\pm$0.31 52.22 $\pm$0.06 51.83 $\pm$0.04 58.23 $\pm$0.29
-JPEG Compression 50.13 51.88 $\pm$0.13 52.38 $\pm$0.24 52.46 $\pm$0.19 54.19 $\pm$0.16 56.82 $\pm$0.24
-Mean 34.98 40.28 41.07 42.98 43.28 50.36
-[…]
-Dataset CLIP ICLR’21 TENT ICLR’21 TPT NeurIPS’22 CLIPArTT WACV’25 WATT-S NeurIPS’24 CLIP-OT Ours
-Tiny-Imagenet 58.29 57.72 $\pm$0.14 58.90 $\pm$0.15 59.85 $\pm$0.01 61.35 $\pm$0.17 63.69 $\pm$0.15
-Gaussian Noise 7.08 8.01 $\pm$0.05 9.29 $\pm$0.03 14.44 $\pm$0.25 13.02 $\pm$0.07 21.40 $\pm$0.37
-Shot Noise 9.41 10.04 $\pm$0.07 11.70 $\pm$0.07 17.44 $\pm$0.12 15.94 $\pm$0.07 24.90 $\pm$0.17
-Impulse Noise 3.44 4.18 $\pm$0.02 4.85 $\pm$0.01 10.37 $\pm$0.16 6.90 $\pm$0.01 17.34 $\pm$0.14
-Defocus Blur 21.71 24.53 $\pm$0.05 27.56 $\pm$0.10 31.46 $\pm$0.26 29.91 $\pm$0.13 35.39 $\pm$0.16
-Glass Blur 9.12 10.09 $\pm$0.06 11.03 $\pm$0.10 15.84 $\pm$0.09 14.01 $\pm$0.10 21.16 $\pm$0.19
-Motion Blur 34.52 36.94 $\pm$0.07 38.97 $\pm$0.05 41.34 $\pm$0.20 41.26 $\pm$0.04 46.26 $\pm$0.37
-Zoom Blur 27.44 29.48 $\pm$0.06 34.29 $\pm$0.07 35.06 $\pm$0.15 33.96 $\pm$0.07 40.93 $\pm$0.05
-Snow 32.51 32.20 $\pm$0.03 34.45 $\pm$0.10 36.79 $\pm$0.08 37.76 $\pm$0.09 42.32 $\pm$0.14
-Frost 36.33 35.72 $\pm$0.02 37.13 $\pm$0.05 38.37 $\pm$0.13 39.65 $\pm$0.03 44.97 $\pm$0.17
-Fog 25.94 27.46 $\pm$0.04 28.89 $\pm$0.08 33.51 $\pm$0.19 32.13 $\pm$0.08 38.60 $\pm$0.32
-Brightness 40.15 39.79 $\pm$0.08 43.31 $\pm$0.10 46.52 $\pm$0.06 46.93 $\pm$0.13 53.10 $\pm$0.05
-Contrast 1.81 2.24 $\pm$0.04 3.15 $\pm$0.04 6.07 $\pm$0.12 3.53 $\pm$0.02 11.88 $\pm$0.23
-Elastic Transform 30.40 31.92 $\pm$0.05 33.88 $\pm$0.14 33.74 $\pm$0.11 35.01 $\pm$0.13 40.73 $\pm$0.16
-Pixelate 22.78 24.79 $\pm$0.03 27.70 $\pm$0.12 34.84 $\pm$0.09 31.55 $\pm$0.08 41.84 $\pm$0.09
-JPEG Compression 29.59 30.93 $\pm$0.13 33.60 $\pm$0.15 37.29 $\pm$0.14 36.46 $\pm$0.11 42.86 $\pm$0.04
-Mean 22.14 23.22 25.32 28.87 27.87 34.91
-[…]
+F Differences wrt PLOT [chenplot]
+Note that despite the apparent similarities, our work presents fundamental differences w.r.t.
+Prompt Learning with Optimal Transport (PLOT) [chenplot], which prevents from adding the later in the empirical evaluation.
+a) Target task.
+PLOT tackles few-shot adaptation, requiring labeled data to adapt no novel tasks, whereas our method studies test-time adaptation, where labels or test samples are not available.
+Thus PLOT cannot be compared in the TTA setting.
+b) Adaptation.
+Our approach updates the layer norm parameters of CLIP visual encoder, whereas PLOT optimizes the input text prompts (i.e., it falls into the Prompt Learning category).
+c) Use of OT.
+PLOT leverages OT between learnable prompts and a set of local visual features, whose transport plan is directly used as predicted probabilities.
+In contrast, we compute the transportation cost between the whole image features and fixed text embeddings, which is later used as a pseudo-supervision to distill knowledge during test-time adaptation.